### v1

# Cold Jupiters and small planets: friends, foes, or indifferent?

11institutetext: INAF - Osservatorio Astrofisico di Torino, via Osservatorio 20, 10025 Pino Torinese, Italy 22institutetext: Dipartimento di Fisica, Università degli Studi di Torino, via Pietro Giuria 1, 10125 Torino, Italy 33institutetext: University of Southern Queensland, Centre for Astrophysics, West Street, Toowoomba, QLD 4350, Australia
Cold Jupiters and small planets: friends, foes, or indifferent?
A search for correlations with the largest exoplanet samples
A. S. Bonomo 11    L. Naponiello 11    E. Pezzetta 22    A. Sozzetti 11    D. Gandolfi 22    R. Wittenmyer 33    M. Pinamonti 11 aldo.bonomo@inaf.it
(Received 7 October 2024 / Accepted 20 May 2025)
Finding out whether there is any correlation between the presence of short-period small planets (SPs) with $P\lesssim 100$ d ($a\lesssim 0.4$ AU) and $1<M_{\rm p}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ and that of outer cold Jupiters (CJs) with $a=1-10$ AU and $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ around solar-type stars may provide crucial constraints on the models of formation and/or migration of SPs. However, somehow discrepant results about the occurrence rates of CJs in SP systems have been reported in the literature, with a few recent works claiming a strong SP-CJ correlation but only at super-solar metallicity and/or mass of the host stars. Here we homogeneously recomputed the occurrence rates of CJs at average, sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity as well as at average and sub-intervals of stellar mass, namely 0.6-0.8, 0.8-1.0, and 1.0-1.2 $\mathrm{M}_{\odot}$, with (i) a carefully-selected sample of 217 SP systems detected through transits or radial velocity, and (ii) a large comparison sample of 1167 solar-type stars, regardless of the possible presence of SPs. We determined the integrated occurrence rate of CJs in SP systems to be $f_{\rm CJ|SP}=11.1^{+2.5}_{-1.8}\%$ at average metallicity $\overline{\rm[Fe/H]}=-0.011\pm 0.005$ and mass $\overline{M_{\star}}=0.916\pm 0.012\,\hbox{$\mathrm{M}_{\odot}$}$; this is consistent with the estimated frequencies of CJs in both the comparison sample ($f_{\rm CJ}=9.8^{+0.9}_{-0.8}\%$ at $\overline{\rm[Fe/H]}=-0.072\pm 0.009$ and $\overline{M_{\star}}=0.994\pm 0.004\,\hbox{$\mathrm{M}_{\odot}$}$) and the HARPS-N survey of transiting SP systems. We found a possible correlation ($f_{\rm CJ|SP}>f_{\rm CJ}$) only at super-solar mass and metallicity, namely $1.0\leq M_{\star}<1.2\,\hbox{$\mathrm{M}_{\odot}$}$ and $\rm[Fe/H]>0.1$, though with a statistical confidence less than $3\sigma$. To test some theoretical predictions, we also searched for possible SP-CJ relations as a function of SP and CJ multiplicity as well as SP composition, albeit with the inevitably limited current sample, and found none. We show that the architectures of SP systems are not indifferent to the presence of CJs as the multiplicity of SPs strongly depends on the CJ eccentricity, as expected from planetary dynamics. A more comprehensive understanding of the relation between SPs and CJs requires larger samples of SP systems. The increasing number of well-characterized systems and the anticipated discoveries from both the Gaia and PLATO missions will enable a definitive assessment of the impact of CJs on the formation of SPs.
Key Words.:
Planetary systems – Planets and satellites: formation – Stars: statistics – Stars: solar-type – Methods: statistical – Techniques: radial velocities.
1 Introduction
Whether the presence of short-period small planets (SPs) with $P\lesssim 100$ d (corresponding to semi-major axis of $a\lesssim 0.4$ AU) and $1<M_{\rm p}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ around solar-type stars is somehow correlated with that of cold Jupiters (CJs), that is gaseous giant planets with $a\simeq 1-10$ AU and mass $0.5\leq M_{\rm p}\leq 20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ (or $0.3\leq M_{\rm p}\leq 13\,\hbox{$\mathrm{M}_{\rm Jup}$}$, depending on the definition), is becoming a very intriguing question in exoplanetary science because it could potentially reveal crucial features of planet formation and evolution. However, there are strong discrepancies about this possible correlation in the literature from both theoretical and observational perspectives.
As for the theory, an anti-correlation between SPs and CJs would be expected if CJs could either (i) act as a dynamical barrier to the inward migration of icy SPs from beyond the water snowline at $\sim 1-3$ AU (Izidoro et al. 2015) and/or (ii) reduce considerably the inward flux of pebbles to the inner regions of the protoplanetary disk, which is required to form dry SPs within the water snowline (Lambrechts et al. 2019). From the Generation 3 Bern simulations of planet formation and evolution, Schlecker et al. (2021) instead found no or a weak SP-CJ correlation, and predicted a strong architecture-composition link, with more icy (rocky) SPs in the absence (presence) of one or more CJs, due to the same above-mentioned effect of dynamical barrier by the CJ (Izidoro et al. 2015). Furthermore, Schlecker et al. (2021) predicted a scarcity of SPs at high stellar metallicity due to the disruption of inner planetary systems by more dynamically active CJs (see their Fig. 19). A recent work by Bitsch & Izidoro (2023) shows that it would also be possible to obtain a correlation between the presence of SPs and that of CJs, depending on the gas contraction rate for the formation of CJs: for slow gas contraction rates, the cores that form in the proximity of the water snowline are too small to effectively accrete large envelopes and can migrate inward thus becoming SPs, while cores at greater distances can grow into cold gas giants.
[…]
The works by Zhu (2024) and BL24 were partly motivated by the fact that those authors noticed that several stars in the Kepler and K2 sample by B23 are metal poor, which might have biased the B23 estimate of $f_{\rm CJ|SP}$ towards lower values for the well-known relation between giant planets and stellar metallicity (e.g., Fischer & Valenti 2005; Johnson et al. 2010; Sousa et al. 2011; Mortier et al. 2013). However, the average metallicity of the B23 sample ($\overline{\rm[Fe/H]}=-0.066\pm 0.014$) and that of the Anglo Australian Telescope (AAT) RV survey ($\overline{\rm[Fe/H]}=-0.045\pm 0.018$) (Wittenmyer et al. 2020), which was used as a comparison sample by B23, are consistent within $1\sigma$ and are both lower than the peak of the metallicity at about $-0.1$ dex in a volume-limited sample of the solar neighborhood (e.g., Sousa et al. 2011).
Building upon the work of BL24, in the present article we recompute (i) $f_{\rm CJ|SP}$ with a larger and more carefully selected sample of SP systems, and (ii) $f_{\rm CJ}$ using three different blind RV campaigns, namely the AAT, CLS, and HARPS (Mayor et al. 2011; Trifonov et al. 2020) surveys, and merging the stars of these three surveys into a unique large sample. We then compare $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ to search for possible SP-CJ correlations as a function of stellar metallicity and/or mass as well as planet properties, and draw our conclusions about their possible existence.
2 Samples and radial velocity data
2.1 Small planet transit and radial velocity sample
The sample of SP systems to compute $f_{\rm CJ|SP}$ was built through queries to the NASA Exoplanet Archive [^3] https://exoplanetarchive.ipac.caltech.edu/ . on 2 December 2024 following the same criteria as BL24: the systems must have (1) at least one confirmed SP with $1<M_{\rm p}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ [^4] Unlike BL24, we did not consider the criterion on the radius to be $R_{\rm p}\leq 4\,\hbox{$\mathrm{R}_{\oplus}$}$ as this information is not available for the RV planets. for transiting planets and $1<M_{\rm p}\sin{i}<20\,\hbox{$\mathrm{M}_{\oplus}$}$ for planets discovered by RV surveys; (2) a solar-type (i.e., FGK dwarf or sub-giant) star as host, which thus excludes M dwarfs with $M_{\star}\leq 0.6\,\hbox{$\mathrm{M}_{\odot}$}$, and (3) at least one publicly available RV dataset with more than 20 measurements and baseline $\Delta T\geq 1$ yr. However, unlike BL24, we initially discarded “mixed systems” with both SP and short-period ($P<100$ d) planets with $M_{\rm p}\geq 20\,\hbox{$\mathrm{M}_{\oplus}$}$, such as the WASP-47 system which also hosts a hot Jupiter (WASP-47b) in between the two SPs WASP-47e and d (e.g., Vanderburg et al. 2017; Nascimbeni et al. 2023); indeed, it is unclear whether the theoretical predictions mentioned in Sect. 1 apply to such systems in which giant planets, especially the hot Jupiters, may have undergone a substantial migration process (e.g., Dawson & Johnson 2018). Nonetheless, we computed the occurrence rates by also taking the mixed systems into account, and noted no significant differences (see Sect. 3). We cleaned the BL24 sample from a couple of outliers like TOI-1853b, which has $M_{\rm p}=73\pm 3>20\,\hbox{$\mathrm{M}_{\oplus}$}$ (Naponiello et al. 2023), or Kepler-22b, whose orbital period is $P=290>100$ d (Borucki et al. 2012).
Our SP sample results in 217 systems, of which 134 were detected by transit surveys and 83 with RVs. It is therefore considerably larger than that of BL24 (184 mixed systems), and contains 148 systems in common (see Table LABEL:table_SPsample). The inclusion of the mixed systems leads to 242 systems, 151 and 91 of which were found with transits and RVs, respectively (Tables LABEL:table_SPsample and LABEL:table_Mixed_sample). As in BL24, for each system we selected the available RV dataset with the best sensitivity to the detection of CJs (see Sect. 3).
2.2 Radial velocity comparison samples
[…]
The mass, metallicity and uncertainty on the metallicity of the selected solar-type stars from the three surveys were all taken from the Hypatia catalog [^5] https://www.hypatiacatalog.com/ . (Hinkel et al. 2014) and are given in Table LABEL:table_RVsurveys.
We also merged the three aforementioned RV surveys to create a large comparison sample with 1167 stars, 213 of which are in common between two or even all three surveys (Table LABEL:table_RVsurveys). In this way, we could compute a more precise and accurate $f_{\rm CJ}$ (see Tables 19 and 23).
2.3 Comparison of the small planet and the RV merged samples
The distributions of the stars of both the SP and RV merged samples with stellar metallicity and mass are compared in Figs. 4 and 5, respectively.
The Kolmogorov-Smirnov (K-S) test between the two empirical cumulative distribution functions (eCDF) of metallicity gives a relatively low p-value of 0.021 as the SP sample has a lower fraction of metal-poor stars at $\rm[Fe/H]\lesssim-0.2$ than the comparison sample (see Fig. 4).
[…]
In Fig. 6 we show the stellar metallicity distributions and eCDFs of both the SP and RV merged samples at the three intervals of stellar mass considered in Sect. 3, namely $0.6-0.8$, $0.8-1.0$, and $1.0-1.2\,\hbox{$\mathrm{M}_{\odot}$}$. While the eCDFs of metallicity are compatible in the $0.6-0.8\,\hbox{$\mathrm{M}_{\odot}$}$ and $0.8-1.0\,\hbox{$\mathrm{M}_{\odot}$}$ intervals with p-values of 99% and 27%, respectively, they considerably differ at super-solar mass (p-value $<1\%$) because of the considerably larger number of metal-poor stars in the RV merged sample (right panels of Fig. 6).
3 Statistical analyses
We determined $f_{\rm CJ|SP}$ from the SP sample (Sect. 2.1), and $f_{\rm CJ}$ from both the individual RV surveys and the comparison sample (Sect. 2.2), first at the average metallicity of the different samples, and then by considering three intervals in stellar metallicity, i.e. sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity. Similarly to BL24, we also split the samples in two broad intervals, namely $\rm[Fe/H]\leq 0$ and $\rm[Fe/H]>0$, even though, in our view, given the typical uncertainties of $\sim 0.04-0.07$ dex on $\rm[Fe/H]$, a star with $0.10.0<\rm[Fe/H]<$ 0.1 ($-0.1<\rm[Fe/H]<0.0$) should not be classified as metal rich (poor), but rather as solar in metallicity.
Given the different distributions of stellar mass in the SP and comparison samples (Sect. 2.3), we further considered three different intervals in stellar mass, namely $0.6\leq M_{\star}<0.8\,\hbox{$\mathrm{M}_{\odot}$}$, $0.8\leq M_{\star}<1.0\,\hbox{$\mathrm{M}_{\odot}$}$, and $1.0\leq M_{\star}<1.2\,\hbox{$\mathrm{M}_{\odot}$}$ [^6] BL25 also considered similar intervals of stellar mass to investigate the SP-CJ relation as a function of stellar metallicity and mass.. This is of interest because the frequency of giant planets also depends on stellar mass, though to a lesser extent than metallicity (e.g., Johnson et al. 2010; Fulton et al. 2021).
For each sample and its relative sub-samples in metallicity we used the binomial statistics
$b(d|N_{\star,\rm eff},f_{\rm CJ|SP})=\frac{N_{\star,\rm eff}!}{d!(N_{\star,\rm eff}-d)!}f_{\rm CJ|SP}^{d}(1-f_{\rm CJ|SP})^{N_{\star,\rm eff}-d},$ (1)
where $d$ is the number of systems with at least one confirmed CJ, and $N_{\star,\rm eff}$ is the “effective” number of stars given by the product of the number of stars of the survey, $N_{\star}$, and the average survey sensitivity (or completeness) to CJs, $C$, i.e. $N_{\star,\rm eff}=N_{\star}\cdot C$.
The completeness of each individual star, $C_{\rm i}$, was computed in the same way as B23 and BL24 (see their Sect. 3), that is through experiments of injection and recovery of CJ RV signals in a logarithmic grid of 30x30 $\Delta M_{\rm p}$-$\Delta a$ cells covering the ranges $0.3-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ and $1-10$ AU, respectively. Specifically, for each cell, we simulated 200 RV signals of CJs at the epochs of the RV dataset by randomly varying (i) $M_{\rm p}$ and $a$ within the cell bounds, (ii) $T_{\rm c}$ within the range of $P$ corresponding to $a$ and the stellar mass $M_{\star}$ (Tables LABEL:table_SPsample, LABEL:table_Mixed_sample and LABEL:table_RVsurveys) from Kepler’s third law; (iii) $\cos{i}$ from 0 to 1, where $i$ is the orbital inclination; and iv) the argument of periastron $\omega$ from 0 to $2\pi$, while drawing the orbital eccentricity $e$ from a $\beta$ distribution (Kipping 2013). We then shifted every RV point at time $t$ following a Gaussian distribution with mean and standard deviations equal to the RV value and its uncertainty, respectively. $C_{\rm i}$ was estimated as the recovery rate of the simulated signals (Keplerian orbits or linear and quadratic trends if $\Delta T$ was considerably shorter than the simulated CJ orbital period) across the whole grid using the $\Delta BIC$ criterion (see B23 and BL24). The average completeness $C$ of each sample (or sub-sample) was then simply obtained by the mean of the completenesses of the stars in that sample: $C=(1/N_{\star})\sum_{i=1}^{N_{\star}}{C_{\rm i}}$. The average completenesses of all the samples described in Sect. 2 are displayed in Fig. 7. We note a very good agreement when comparing our average completenesses of the SP RV and transit samples (upper panels in Fig. 7) with those shown in Fig. A.1 of BL24 (for $a=1-10$ AU).
We considered two ranges of CJ mass [^7] Note that only minimum masses, $M_{\rm p}\sin{i}$, are known for the majority of CJs as their orbital inclination $i$ cannot be determined from RVs alone., namely $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ and $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ (for comparison with B23), but we mainly rely on the former because the latter suffers from incompleteness issues between 0.3 and 0.5 $\mathrm{M}_{\rm Jup}$ (Fig. 7). For each CJ mass range and the aforementioned bins of stellar metallicity and/or mass we give the number of solar-type stars ($N_{\star}$), the number of stars with confirmed CJs ($d$), the average completenesses ($C$), and CJ occurrence rates from (i) the SP and the large comparison samples in Tables 19 and 20, (ii) the SP transit, RV, and transit+RV samples as well as the SP and mixed transit, RV, and transit+RV samples in Table 21, and (iii) the AAT, CLS, and HARPS surveys in Table 23. Tables LABEL:table_CJs_SPsample, LABEL:table_CJs_SPsample_mixed, and LABEL:table_CJs_RVsurveys list the known CJs and their properties in the SP, SP and mixed, and comparison samples. Fig. 11 shows on the left (right) the architectures of the 27 (23) systems with CJs with $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ ($0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$) in the SP sample, respectively.
4 Results
4.1 Frequency of CJs at average stellar metallicity and mass
At the average stellar metallicity ($\overline{\rm[Fe/H]}=-0.011\pm 0.005$) and mass ($\overline{M_{\star}}=0.916\pm 0.012\,\hbox{$\mathrm{M}_{\odot}$}$) of the SP (transit+RV) stars we determined $f_{\rm CJ|SP}=11.1^{+2.5}_{-1.8}\%$ for the CJ mass range $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$. Adding the 25 mixed systems yielded $f_{\rm CJ|SP+Mixed}=12.5^{+2.5}_{-1.8}\%$. Going back to the original sample of BL24 with $\overline{\rm[Fe/H]}=-0.001\pm 0.005$ and $\overline{M_{\star}}=0.923\pm 0.014\,\hbox{$\mathrm{M}_{\odot}$}$, which includes the “mixed systems”, and accounting for its average completeness [^8] This was extracted from Fig. A.1 in BL24., we found a fully compatible result: $15.2^{+3.1}_{-2.3}\%$. It is interesting to note that both occurrence rates well agree within $1\sigma$ with the estimate of $f_{\rm CJ|SP}=8.7^{+7.3}_{-2.7}\%$ by B23, whose smaller sample is on average slightly poorer in metals by $\sim 0.06$ dex, and are lower than previous claims of $f_{\rm CJ|SP}\sim 40\%$ (Sect. 1).
As for the comparison sample, at its average metallicity ($\overline{\rm[Fe/H]}=-0.072\pm 0.009$) and mass ($\overline{M_{\star}}=0.994\pm 0.004\,\hbox{$\mathrm{M}_{\odot}$}$), we found $f_{\rm CJ}=9.8^{+0.9}_{-0.8}\%$, which is fully compatible with $f_{\rm CJ|SP}$ [^9] The possible, though not significant, hint for a SP-CJ anti-correlation reported by B23 can be ruled out as our $f_{\rm CJ}$ recomputed from the AAT survey is approximately half that derived from Wittenmyer et al. (2020)..
From our occurrence rates $f_{\rm CJ|SP}$ and $f_{\rm CJ}$, and the frequency of SPs $f_{\rm SP}=28.1^{+6.6}_{-5.1}\%$ from Rosenthal et al. (2022) (see also Zhu et al. 2018), we can derive the conditional probability that a solar-type star with a CJ also hosts a SP (Zhu & Wu 2018), namely $f_{\rm SP|CJ}=f_{\rm SP}\times f_{\rm CJ|SP}\leavevmode\nobreak\ /\leavevmode\nobreak\ f_{\rm CJ}=32\pm 11\%$. This agrees within $1\sigma$ with the $f_{\rm SP|CJ}$ computed from a sample of 115 solar-type stars known to host a CJ (Ruggieri, A., et al. in prep.), and is substantially lower than the estimate of $f_{\rm SP|CJ}\sim 100\%$ in Bryan et al. (2019).
4.2 Frequency of CJs as a function of stellar metallicity
We report $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ for the three sub-intervals of stellar metallicity (Sect. 3) in Table 19, and display them in Fig. 1 (see also Fig. 9 for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$). For the same metallicity bins, we also give the estimates of $f_{\rm CJ}$ from the AAT, CLS, and HARPS surveys separately in Table 23, which are fully consistent with each other (see Fig. 8).
By comparing $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ in Table 19, we found that (i) at sub-solar and solar metallicities, they are fully compatible within $1\sigma$, and (ii) at super-solar metallicity, the value of $f_{\rm CJ|SP}$ is higher than $f_{\rm CJ}$ by approximately $3/2$ ($\sim 30\%$ vs $\sim 20\%$; Fig. 1), but the significance of this difference is low, that is $1.7\sigma$ (compared to $2.7\sigma$ in BL24 for $\rm[Fe/H]>0$).
Figure 1: Occurrence rates of cold Jupiters with $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ in the small planet sample (red solid line) and the merged RV comparison sample (black dashed line) at sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity. See also Table 19.
Figure 2: Same as Fig. 1 for three different bins of stellar mass, namely $0.6-0.8\,\hbox{$\mathrm{M}_{\odot}$}$ (left), $0.8-1.0\,\hbox{$\mathrm{M}_{\odot}$}$ (center), and $1.0-1.2\,\hbox{$\mathrm{M}_{\odot}$}$ (right); see Table 20.
4.3 Frequency of CJs as a function of stellar metallicity and mass
We report $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ for the same bins of stellar metallicity as above, and for the three ranges of stellar mass discussed in Sect. 3, in Table 20. We show them in Fig. 2 (see also Fig. 10 for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$).
We found an excess of CJs in SP systems at $2.5\sigma$ significance level only at super-solar mass and metallicity, namely $M_{\star}\geq 1.0\,\hbox{$\mathrm{M}_{\odot}$}$ and $\rm[Fe/H]>0$ (see the right plot in Fig. 2, and Table 20), in agreement with the recent findings of BL25. This is the origin of the higher $f_{\rm CJ|SP}$ compared to $f_{\rm CJ}$ at super-solar metallicity and integrated stellar masses (Sect- 4.2 and Fig. 1).
4.4 SP-CJ relations as a function of SP and CJ multiplicity
We explored possible SP-CJ trends as a function of SP multiplicity, for which purpose we chose the $0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ mass range (instead of $0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$), given the higher number of systems with both CJs and SPs (27 compared to 23). Indeed, if the presence of CJs really hindered the formation/migration of SPs, a lower degree of SP multiplicity would be expected in the presence of CJs, for instance because low-mass “jumpers” overtaking the CJ towards inner orbits most likely form a single SP system (see Fig. 3 in Izidoro et al. 2015). By comparing the frequency of multiple systems with a number of SPs $n_{\rm p}\geq 2$ in the SP sub-samples with and without CJs, we found $f_{n_{\rm p}\geq 2,\rm w/\,CJ}=37^{+10}_{-8}\%$ (10 out of 27 systems) and $f_{n_{\rm p}\geq 2,\rm w/o\,CJ}=43^{+4}_{-3}\%$ (81 out of 190 systems), respectively; these rates are statistically consistent within the uncertainties.
We also searched for possible (anti)correlations as a function of CJ multiplicity to investigate whether two (or more) CJs may hinder the formation of SPs more than a single CJ system (Sect. 1). To this end, we compared the frequency of single SPs in systems with one CJ to that in systems with two (or more) CJs; indeed, the latter should be higher if two (or more) CJs had a more deleterious effect than a single CJ on the formation of multiple SP systems. We did not find any evidence for that: the rates of single SP systems with one CJ and two CJs are compatible, namely $63^{+9}_{-12}\%$ (12 out of 19 systems) and $62^{+13}_{-18}\%$ (5 out of 8 systems), respectively (see Fig. 11).
4.5 SP-CJ relation as a function of SP composition
We investigated the possible presence of an architecture-composition link by determining the frequency of rocky SPs [^10] We define a rocky planet as a planet that is located between the 100% iron and 100% silicate compositions in the mass-radius diagram of small planets (e.g., Zeng & Sasselov 2013). in CJ systems. This is intended to test whether the formation of rocky planets may be favored in the presence of CJs in the event CJs block the migration of icy sub-Neptunes from beyond the water snowline (Schlecker et al. 2021). We derived $f_{\rm rocky,w/\,CJ}=33^{+13}_{-9}\%$ from 5 transiting systems with rocky planets (HD219134, Kepler-139, Kepler-407, K2-312, K2-364) out of 15 in total (including 3 RV systems that were later found transiting). This shows that systems with both rocky planets and CJs do not represent the majority of the SP-CJ transiting systems.
4.6 SP multiplicity as a function of CJ eccentricity
When ordering the SP systems with CJs as a function of the CJ eccentricity (Fig. 3), only single or, more rarely, two SPs are found with eccentric ($e\gtrsim 0.4$) CJs, the most extreme case being K2-312/HD 80653 (B23). On the contrary, higher multiplicity (up to $n_{\rm p}=4$ for HD 219134) SP systems are seen in the company of CJs in quasi-circular orbits. This is in line with the expectation of stronger dynamical perturbations at inner orbital separations by eccentric CJs (e.g., Huang et al. 2017).
Figure 3: The architectures of the 27 systems with both small planets and cold Jupiters in the mass range $0.3\leq{\rm M_{p}}\leq 13\,\hbox{$\mathrm{M}_{\rm Jup}$}$, ordered according to increasing orbital eccentricity of the cold Jupiters (from top to bottom; in case of multiple cold Jupiters, the highest eccentricity was considered). The planet circles have sizes proportional to their mass (see legend in the top right corner), and colors depending on the orbital eccentricity (more eccentric planets are shown in yellow). The host stars are color coded as a function of stellar metallicity (metallicity increases from blue to red). All the systems with three or more inner small planets are accompanied by cold Jupiters with low eccentricities.
5 Discussion and conclusions
5.1 Cold Jupiters and small planets are friends, foes, or indifferent?
In the present work we first investigated the possible SP-CJ relation as a function of stellar metallicity and mass, and did not reveal any correlation or anticorrelation (i.e., SPs and CJs are neither friends nor foes) in eight of the nine mass and metallicity bins, that is at sub-solar mass for any metallicity as well as at super-solar mass for solar and sub-solar metallicity. We instead found an enhancement of CJs in SP systems (i.e., SPs and CJs could be friends) at super-solar mass and metallicity, though at less than $3\sigma$, in general agreement with BL25.
We could not unveil any SP-CJ trends with planet multiplicity and composition, which would further disfavor the anticorrelation models predicting a dearth of CJs in SP systems (Sect. 1). Nonetheless, it should be noted that (i) the limited sub-samples of SP systems with CJs (27 systems) inevitably result in large error bars on the CJ occurrence rates, making it difficult to reveal possible weak correlations or anticorrelations with planet properties; and (ii) the SP multiplicity considered in this work relies on the known (that is, detected) SPs and may be different from the true one, especially if non-transiting SPs with relatively long period ($P\sim 25-100$ d) and/or low mass went undetected by the RV follow up.
Our investigation of the architecture-composition relation in Sect. 4.5 must be considered with caution because it only concerned rocky planets. Yet, in principle it may also apply to rocky planets with tenuous H/He atmospheres, if some available gas was accreted by the super-Earths at the time of their formation, and was not later on stripped away by post-formation processes such as atmospheric photoevaporation (e.g., Lopez & Fortney 2014) or core-powered mass loss (e.g., Ginzburg et al. 2018). However, the well-known degeneracy in SP composition from the measurement of the bulk density alone does not allow us to distinguish whether or which non-rocky planets in the 10 transiting systems with CJs of our SP sample are actually dry (rocky worlds shrouded by thin H/He atmospheres, which formed within the water snowline) or icy (water worlds with possible H/He envelops, which formed beyond the water snowline and then migrated inwards). This makes the investigation of the architecture-composition link challenging.
We showed that higher multiplicity SP systems are found in systems with CJs in low-eccentricity orbits, being immune to dynamical instabilities generated by highly eccentric CJs. In this sense, SPs are not indifferent to CJs, even in the great majority of cases where $f_{\rm CJ|SP}\sim f_{\rm CJ}$.
5.2 Caveats of our, BL24, and BL25 analyses
The occurrence rates derived in this work, BL24, and BL25 may be slightly underestimated mainly because of two reasons, which however should affect both the SP and comparison samples approximately in the same way, making the search for correlations in principle reliable. Firstly, the completenesses computed in our and BL24 works are optimistic because they only account for the formal RV uncertainties, and not for the uncorrelated jitter (as properly done by B23), which in some cases may be considerably higher than the former [^11] Nevertheless, proper Markov chain Monte Carlo analyses of the RV datasets of $\sim 1350$ systems considered here, similarly to B23, would require unrealistic computation times.. Secondly, a few linear or quadratic RV trends, which might later result in CJ signals, were not taken into account in the computation of the occurrence rates, which relies on confirmed CJs only. Nevertheless, among the targets in the SP sample with relatively short temporal baselines, that is $1\leq\Delta T\leq 4$ yr, we noticed only two targets with RV long-term trends still compatible with CJs (namely TOI-1279 and TOI-1778). Moreover, such trends are very often induced by brown dwarfs or low-mass stellar companions, or they correlate with activity indicators (see B23).
Conversely, occurrence rates might decrease if the minimum masses ($M_{\rm p}\sin{i}$) of some of the considered CJs turned into true mass ($M_{\rm p}$) measurements above the 20 (or 13) $\mathrm{M}_{\rm Jup}$ upper limit, corresponding to brown dwarfs or low-mass stellar objects, due to low orbital inclinations (e.g, Xiao et al. 2023).
5.3 Perspectives
As more and more SP systems are characterized with high-precision RV follow up and monitored over the years, more precise and accurate determinations of $f_{\rm CJ|SP}$ can be derived, by allowing to investigate more deeply the SP-CJ relation, even at the lower CJ mass range $0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$. We plan to update our results over time by adding the increasing number of well-characterized SP systems.
[…]
Acknowledgements.
We are are grateful to the anonymous referee for her/his valuable comments, which allowed us to considerably improve this work. We are thankful to Profs. A. Morbidelli and B. Bitsch for useful discussions about theoretical predictions on SP-CJ relations. We also thank Dr. K. Biazzo for her suggestion to use the Hypatia catalog. We acknowledge financial contribution from the INAF Large Grant 2023 “EXODEMO” and from the European Union - Next Generation EU RRF M4C2 1.1 PRIN MUR 2022 project 20229R43BH. D.G. gratefully acknowledges the financial support from the grant for internationalization (GAND_GFI_23_01) granted by the University of Turin (Italy). This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. The research shown here acknowledges use of the Hypatia Catalog Database, an online compilation of stellar abundance data, which was supported by NASA’s Nexus for Exoplanet System Science (NExSS) research coordination network and the Vanderbilt Initiative in Data-Intensive Astrophysics (VIDA).
Appendix A Sample of solar-type stars in the small planet sample
Table 1: Stars of the small planet sample and their properties.
Target M [Fe/H] $N_{pl}$ Method $N_{obs}$ Baseline CJ CJ Ref. for RVs and parameters
[M] [dex] [yrs] [$0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$] [$0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$] (if different from RVs)
61-Vir $0.91\pm 0.04$ $0.03\pm 0.06$ 3 RV 602 15.20 No No 96
BD-08-2823 $0.74\pm 0.07$ $0.00\pm 0.02^{1}$ 2 RV 83 5.04 No No 49
BD-11-4672 $0.65\pm 0.03$ $-0.35\pm 0.15$ 2 RV 111 14.16 Yes Yes 10
CoRoT-7 $0.90\pm 0.12$ $0.06\pm 0.10$ 3 Transit 180 3.90 No No 48
CoRoT-22 $1.10\pm 0.05$ $0.17\pm 0.12$ 1 Transit 23 4.90 No No 77
CoRoT-24 $0.91\pm 0.09$ $0.30\pm 0.15$ 2 Transit 37 3.10 No No 4
DMPP-4 $1.25\pm 0.02$ $-0.01\pm 0.06$ 2 RV 71 4.27 No No 11
GJ-96 $0.60\pm 0.07$ $0.14\pm 0.08$ 1 RV 72 5.53 No No 52
GJ-160.2 $0.69\pm 0.10$ $-0.26\pm 0.10^{1}$ 1 RV 100 6.37 No No 120,111
GJ-338-B $0.64\pm 0.07$ $-0.03\pm 0.16$ 1 RV 114 9.25 No No 33
GJ-414-A $0.69\pm 0.01$ $0.24\pm 0.07$ 2 RV 420 6.48 No No 96
GJ-676 $0.73\pm 0.21$ $0.08\pm 0.10$ 4 RV 127 8.85 Yes Yes 21,105
GJ-2030 $0.96\pm 0.13$ $-0.44\pm 0.14^{1}$ 2 RV 48 8.93 No No 110,39
GJ-2056 $0.62\pm 0.08$ $-0.08\pm 0.10^{1}$ 1 RV 51 9.07 No No 38
GJ-3138 $0.68\pm 0.10$ $-0.30\pm 0.12$ 3 RV 199 8.03 No No 7
GJ-3222 $0.89\pm 0.11$ $0.08\pm 0.04^{1}$ 1 RV 60 4.02 No No 110,39
GJ-3942 $0.63\pm 0.07$ $-0.04\pm 0.09$ 1 RV 145 3.30 No No 87
GJ-9404 $0.62\pm 0.07$ $-0.08\pm 0.04^{1}$ 1 RV 54 3.92 No No 92,14
HD-1461 $1.03\pm 0.05$ $0.18\pm 0.06$ 2 RV 846 23.30 No No 96
HD-4308 $0.93\pm 0.10$ $-0.35\pm 0.10$ 1 RV 348 13.92 No No 110,105
HD-7924 $0.80\pm 0.03$ $-0.12\pm 0.06$ 3 RV 967 18.10 No No 96
HD-12235 $1.10\pm 0.13$ $0.32\pm 0.05^{1}$ 2 RV 44 24.10 No No 39
HD-13808 $0.77\pm 0.02$ $-0.21\pm 0.02$ 2 RV 246 11.10 No No 1
HD-18143 $0.85\pm 0.10$ $0.32\pm 0.06^{1}$ 2 RV 80 20.99 No No 39
HD-20003 $0.88\pm 0.10$ $0.04\pm 0.02$ 2 RV 203 13.73 No No 110,113
HD-20781 $0.70\pm 0.10$ $-0.11\pm 0.02$ 4 RV 237 13.74 No No 110,113
HD-21693 $0.80\pm 0.10$ $0.00\pm 0.10$ 2 RV 213 11.25 No No 110,113
HD-22496 $0.68\pm 0.01$ $-0.08\pm 0.02$ 1 RV 43 6.90 No No 63
HD-24085 $1.22\pm 0.07$ $0.17\pm 0.06^{1}$ 1 RV 25 8.66 No No 37
HD-31527 $0.96\pm 0.10$ $-0.17\pm 0.01$ 3 RV 268 13.84 No No 110,113
HD-34445 $1.11\pm 0.06$ $0.14\pm 0.06$ 6 RV 222 21.80 Yes Yes 96
HD-39091 $1.07\pm 0.04$ $0.09\pm 0.04$ 3 RV 402 2.20 Yes Yes 46
HD-39194 $0.67\pm 0.04$ $-0.61\pm 0.02$ 3 RV 273 13.69 No No 114
HD-39855 $0.87\pm 0.05$ $-0.43\pm 0.04^{1}$ 1 RV 25 6.22 No No 37
HD-40307 $0.77\pm 0.10$ $-0.31\pm 0.03$ 5 RV 226 10.44 No No 32
HD-45184 $1.02\pm 0.05$ $0.04\pm 0.06$ 2 RV 339 23.40 No No 96
HD-48948 $0.69\pm 0.02$ $-0.21\pm 0.03$ 3 RV 189 9.50 No No 27
HD-51608 $0.80\pm 0.10$ $-0.07\pm 0.01$ 2 RV 201 9.95 No No 39,113
HD-69830 $0.89\pm 0.04$ $0.01\pm 0.06$ 3 RV 529 13.24 No No 96
HD-74698 $1.04\pm 0.03$ $0.07\pm 0.02$ 2 RV 119 14.10 Yes No 41
HD-77338 $1.37\pm 0.32$ $0.32\pm 0.10$ 1 RV 32 7.22 No No 56,105
HD-86226 $1.02\pm 0.06$ $0.02\pm 0.05$ 2 RV 27 5.10 Yes No 109
HD-90156 $0.86\pm 0.04$ $-0.18\pm 0.06$ 1 RV 167 23.18 No No 96
HD-93351 $0.94\pm 0.12$ $-0.20\pm 0.02$ 2 RV 37 13.97 No No 110,39
HD-93385 $1.04\pm 0.01$ $0.02\pm 0.01$ 3 RV 241 13.23 No No 114
HD-96700 $0.89\pm 0.01$ $-0.18\pm 0.01$ 3 RV 254 14.40 No No 114
HD-97658 $0.77\pm 0.03$ $-0.23\pm 0.06$ 1 RV 577 22.83 No No 96
HD-125612 $1.11\pm 0.20$ $0.25\pm 0.04$ 3 RV 72 12.13 Yes Yes 75
HD-134060 $1.09\pm 0.10$ $0.14\pm 0.01$ 2 RV 339 13.21 No No 110,113
HD-134606 $1.05\pm 0.10$ $0.34\pm 0.08$ 5 RV 66 17.30 No No 61
HD-136352 $0.87\pm 0.04$ $-0.24\pm 0.05$ 3 RV 246 13.20 No No 30,57
HD-141004 $1.05\pm 0.05$ $0.03\pm 0.06$ 1 RV 873 6.64 No No 96
HD-144899 $1.16\pm 0.16$ $0.37\pm 0.12^{1}$ 1 RV 22 6.34 No No 110,39
HD-153557 $0.79\pm 0.09$ $-0.26\pm 0.09$ 3 RV 36 22.20 No No 96
HD-154088 $0.91\pm 0.02$ $0.28\pm 0.03$ 1 RV 187 11.35 No No 114
HD-155918 $1.03\pm 0.13$ $-0.65\pm 0.06^{1}$ 2 RV 73 13.43 No No 39
HD-156098 $1.29\pm 0.20$ $0.19\pm 0.10^{1}$ 2 RV 61 12.94 Yes Yes 110,39
HD-156668 $0.77\pm 0.02$ $0.05\pm 0.06$ 1 RV 619 16.67 No No 96
HD-158259 $1.08\pm 0.10$ $0.00\pm 0.12^{1}$ 5 RV 580 7.01 No No 45
HD-160691 $1.13\pm 0.02$ $0.28\pm 0.03$ 4 RV 180 17.31 Yes Yes 16
HD-164595 $0.99\pm 0.03$ $-0.04\pm 0.08$ 1 RV 75 2.33 No No 25
HD-164922 $0.92\pm 0.03$ $0.20\pm 0.06$ 4 RV 413 23.68 Yes No 96
HD-168009 $0.99\pm 0.04$ $0.02\pm 0.06$ 1 RV 450 5.74 No No 96
HD-175607 $0.71\pm 0.01$ $-0.62\pm 0.01$ 1 RV 110 9.29 No No 76
HD-176986 $0.79\pm 0.02$ $0.03\pm 0.05$ 2 RV 259 13.16 No No 106
HD-177565 $1.00\pm 0.10$ $0.08\pm 0.01^{1}$ 1 RV 68 4.61 No No 36
HD-179079 $1.14\pm 0.19$ $0.23\pm 0.10$ 1 RV 86 11.20 No No 75
HD-181433 $0.78\pm 0.10$ $0.33\pm 0.13$ 3 RV 200 13.78 Yes Yes 54
HD-189567 $0.83\pm 0.01$ $-0.24\pm 0.01$ 2 RV 256 13.20 No No 114
HD-190007 $0.77\pm 0.02$ $0.16\pm 0.05$ 1 RV 157 6.23 No No 22,104
HD-190360 $0.99\pm 0.04$ $0.23\pm 0.06$ 2 RV 754 6.72 Yes Yes 96
HD-192310 $0.84\pm 0.03$ $0.08\pm 0.06$ 2 RV 429 15.23 No No 96
HD-204313 $1.05\pm 0.03$ $0.18\pm 0.02$ 3 RV 113 11.39 Yes Yes 110,119
HD-211970 $0.61\pm 0.04$ $-0.50\pm 0.10^{1}$ 1 RV 52 8.50 No No 37
HD-213519 $1.05\pm 0.13$ $0.03\pm 0.02^{1}$ 1 RV 28 19.96 No No 39
HD-215497 $0.92\pm 0.05$ $0.23\pm 0.07$ 2 RV 99 5.05 Yes No 65
HD-216520 $0.82\pm 0.04$ $-0.16\pm 0.10$ 2 RV 503 15.74 No No 22
HD-217786 $1.02\pm 0.03$ $-0.14\pm 0.04$ 2 RV 27 8.17 No No 110,119
HD-219134 $0.79\pm 0.03$ $0.08\pm 0.06$ 5 RV 584 6.07 Yes No 96
HD-219828 $1.18\pm 0.20$ $0.18\pm 0.10$ 2 RV 91 8.51 No Yes 75,98
HIP-4845 $0.62\pm 0.04$ $-0.60\pm 0.04^{1}$ 1 RV 57 9.13 No No 38
HIP-19976 $0.67\pm 0.08$ $-0.29\pm 0.12^{1}$ 1 RV 43 11.01 No No 39
HIP-35173 $0.79\pm 0.05$ $-0.05\pm 0.02^{1}$ 1 RV 39 8.96 No No 37
HIP-38594 $0.61\pm 0.02$ $0.30\pm 0.08^{1}$ 2 RV 37 8.91 No No 38
HIP-71135 $0.66\pm 0.04$ $0.12\pm 0.14$ 1 RV 44 8.24 No No 37
HIP-107772 $0.63\pm 0.08$ $-0.66\pm 0.10$ 1 RV 49 9.20 No No 38
K2-2/HIP-116454 $0.76\pm 0.03$ $-0.16\pm 0.08$ 1 Transit 115 10.30 No No 18,79
K2-12 $0.97\pm 0.06$ $0.00\pm 0.07$ 1 Transit 75 9.20 No No 18,79
K2-24 $1.07\pm 0.06$ $0.34\pm 0.04$ 2 Transit 64 2.30 No No 88
K2-32 $0.83\pm 0.02$ $-0.06\pm 0.03$ 4 Transit 131 3.20 No No 62
K2-36 $0.79\pm 0.01$ $-0.09\pm 0.06$ 2 Transit 85 5.10 No No 18
K2-38 $1.05\pm 0.07$ $0.23\pm 0.07$ 2 Transit 45 8.20 No No 18,79
K2-66 $1.11\pm 0.04$ $-0.05\pm 0.02$ 1 Transit 38 1.30 No No 101
K2-79 $1.07\pm 0.07$ $0.04\pm 0.06$ 1 Transit 81 9.00 No No 18,79
K2-96/HD-3167 $0.84\pm 0.05$ $0.04\pm 0.05$ 4 Transit 218 7.30 No No 18,79
K2-106 $0.95\pm 0.06$ $0.11\pm 0.07$ 2 Transit 44 3.40 No No 18
K2-110 $0.74\pm 0.02$ $-0.34\pm 0.03$ 1 Transit 20 8.10 No No 18
K2-111 $0.84\pm 0.02$ $-0.46\pm 0.05$ 2 Transit 114 3.40 No No 18
K2-131 $0.80\pm 0.03$ $-0.04\pm 0.07$ 1 Transit 88 6.30 No No 18,79
K2-135/GJ-9827 $0.61\pm 0.02$ $-0.26\pm 0.09$ 3 Transit 51 7.20 No No 18
K2-136 $0.74\pm 0.04$ $0.05\pm 0.10$ 3 Transit 93 2.20 No No 74
K2-141 $0.71\pm 0.03$ $0.00\pm 0.06$ 2 Transit 49 4.10 No No 18
K2-167 $1.01\pm 0.08$ $-0.40\pm 0.06$ 1 Transit 84 9.00 No No 18,79
K2-199 $0.71\pm 0.02$ $-0.01\pm 0.09$ 2 Transit 45 2.90 No No 2
K2-222 $0.99\pm 0.07$ $-0.32\pm 0.06$ 1 Transit 70 4.40 No No 18
K2-262/Wolf-503 $0.69\pm 0.02$ $-0.47\pm 0.08$ 1 Transit 25 3.00 No No 18
K2-263 $0.88\pm 0.03$ $-0.08\pm 0.05$ 1 Transit 95 4.20 No No 18
K2-265 $0.92\pm 0.02$ $0.08\pm 0.02$ 1 Transit 144 1.10 No No 60
K2-312/HD-80653 $1.18\pm 0.04$ $0.26\pm 0.07$ 2 Transit 208 2.90 Yes Yes 18
K2-314 $1.05\pm 0.05$ $0.20\pm 0.05$ 3 Transit 72 1.40 No No 51
K2-364/HD-137496 $1.04\pm 0.02$ $-0.03\pm 0.04$ 2 Transit 142 2.00 Yes Yes 9
K2-370 $0.96\pm 0.04$ $0.08\pm 0.05$ 1 Transit 85 6.10 No No 103
K2-418 $0.97\pm 0.04$ $-0.12\pm 0.05$ 1 Transit 129 7.00 No No 18,79
Kepler-10 $0.91\pm 0.02$ $-0.15\pm 0.04$ 3 Transit 88 12.00 No No 118,18
Kepler-11 $0.99\pm 0.03$ $0.05\pm 0.04$ 6 Transit 31 11.60 No No 118,42
Kepler-18 $0.97\pm 0.04$ $0.19\pm 0.06$ 3 Transit 25 8.80 No No 118,24
Kepler-19 $0.94\pm 0.04$ $-0.13\pm 0.06$ 3 Transit 104 8.00 No No 18
Kepler-20 $0.93\pm 0.05$ $0.07\pm 0.08$ 6 Transit 131 5.40 No No 18
Kepler-21 $1.41\pm 0.03$ $-0.03\pm 0.10$ 1 Transit 102 10.50 No No 18,79
Kepler-25 $1.19\pm 0.06$ $-0.04\pm 0.10$ 3 Transit 101 13.30 No No 118,71
Kepler-36 $1.03\pm 0.02$ $-0.18\pm 0.04$ 2 Transit 25 9.20 No No 118,117
Kepler-37 $0.79\pm 0.03$ $-0.36\pm 0.05$ 3 Transit 114 5.40 No No 18
Kepler-48 $0.88\pm 0.06$ $0.17\pm 0.07$ 5 Transit 59 12.80 Yes Yes 118,71
Kepler-50 $1.15\pm 0.05$ $-0.09\pm 0.04$ 2 Transit 40 11.20 No No 118,42
Kepler-65 $1.25\pm 0.03$ $0.13\pm 0.04$ 4 Transit 79 11.10 No No 118,42
Kepler-68 $1.08\pm 0.05$ $0.12\pm 0.07$ 4 Transit 98 13.40 Yes Yes 118,71
Kepler-78 $0.78\pm 0.05$ $-0.18\pm 0.08$ 1 Transit 121 11.30 No No 18,79
Kepler-90/KOI-351 $1.11\pm 0.04$ $0.10\pm 0.04$ 8 Transit 34 11.20 No No 118,42
Kepler-92 $1.30\pm 0.09$ $0.20\pm 0.06$ 3 Transit 23 9.90 No No 118,90
Kepler-93 $0.91\pm 0.03$ $-0.18\pm 0.10$ 1 Transit 100 14.10 No No 118,18,79
Kepler-94 $0.81\pm 0.06$ $0.34\pm 0.07$ 2 Transit 39 12.00 Yes Yes 118,71
Kepler-95 $1.08\pm 0.08$ $0.30\pm 0.10$ 1 Transit 36 7.90 No No 118,71
Kepler-96 $1.00\pm 0.06$ $0.04\pm 0.07$ 1 Transit 56 13.10 No No 118,71
Kepler-97 $0.94\pm 0.06$ $-0.20\pm 0.07$ 1 Transit 31 8.00 No No 118,71
Kepler-98 $0.99\pm 0.06$ $0.18\pm 0.07$ 1 Transit 42 7.90 No No 118,71
Kepler-99 $0.79\pm 0.06$ $0.18\pm 0.07$ 2 Transit 45 7.00 No No 118,71
Kepler-102 $0.80\pm 0.02$ $0.11\pm 0.04$ 5 Transit 74 8.10 No No 18
Kepler-103 $1.21\pm 0.03$ $0.16\pm 0.04$ 2 Transit 28 12.20 No No 118,18
Kepler-104 $0.82\pm 0.01$ $-0.42\pm 0.04$ 3 Transit 44 11.30 No No 118,42
Kepler-106 $1.00\pm 0.06$ $-0.12\pm 0.11$ 4 Transit 48 9.80 No No 118,71
Kepler-107 $1.24\pm 0.03$ $0.32\pm 0.07$ 4 Transit 121 6.00 No No 18
Kepler-109 $1.04\pm 0.06$ $-0.08\pm 0.10$ 2 Transit 35 10.80 No No 118,71
Kepler-113 $0.75\pm 0.06$ $0.05\pm 0.07$ 2 Transit 42 12.10 No No 118,71
Kepler-129 $1.18\pm 0.03$ $0.29\pm 0.10$ 3 Transit 55 9.00 Yes Yes 118,121
Kepler-131 $1.02\pm 0.06$ $0.12\pm 0.07$ 2 Transit 46 7.00 No No 118,71
Kepler-139 $1.08\pm 0.03$ $0.35\pm 0.04$ 4 Transit 38 11.80 Yes Yes 118,42
Kepler-323 $1.02\pm 0.07$ $-0.14\pm 0.07$ 2 Transit 51 10.30 No No 18,79
Kepler-406 $1.07\pm 0.06$ $0.18\pm 0.07$ 2 Transit 58 11.90 No No 118,71
Kepler-407 $1.00\pm 0.06$ $0.33\pm 0.07$ 2 Transit 98 11.00 Yes Yes 118,71
Kepler-409 $0.91\pm 0.05$ $0.05\pm 0.07$ 1 Transit 97 10.70 No No 118,18
Kepler-454 $1.03\pm 0.04$ $0.32\pm 0.08$ 3 Transit 125 10.20 Yes Yes 18,79
Kepler-507 $1.14\pm 0.03$ $0.14\pm 0.04$ 1 Transit 49 10.80 No No 118,42
Kepler-538 $0.89\pm 0.05$ $-0.09\pm 0.07$ 1 Transit 46 13.00 No No 118,18
Kepler-1655 $1.03\pm 0.04$ $-0.24\pm 0.05$ 1 Transit 99 9.20 No No 18,79
TIC-365102760 $1.17\pm 0.12$ $0.09\pm 0.14$ 1 Transit 20 1.10 No No 44
TOI-130/HD-5278 $1.13\pm 0.04$ $-0.12\pm 0.04$ 2 Transit 41 1.10 No No 102
TOI-134/L-168-9 $0.62\pm 0.03$ $0.04\pm 0.17$ 1 Transit 49 10.40 No No 8
TOI-174/HD-23472 $0.67\pm 0.03$ $-0.20\pm 0.05$ 5 Transit 104 1.70 No No 12
TOI-238 $0.79\pm 0.02$ $-0.11\pm 0.05$ 2 Transit 68 2.40 No No 107
TOI-260 $0.62\pm 0.03$ $-0.47\pm 0.03$ 1 Transit 45 1.70 No No 53
TOI-266/HIP-8152 $0.91\pm 0.06$ $-0.09\pm 0.09$ 2 Transit 94 3.00 No No 3
TOI-286 $0.83\pm 0.05$ $-0.05\pm 0.02$ 2 Transit 47 1.20 No No 53
TOI-402/HD-15337 $0.85\pm 0.04$ $0.03\pm 0.04$ 2 Transit 85 13.70 No No 35
TOI-431 $0.78\pm 0.07$ $0.20\pm 0.05$ 3 Transit 162 15.80 No No 83
TOI-469/HIP-29442 $0.89\pm 0.04$ $0.24\pm 0.05$ 3 Transit 83 2.50 No No 28
TOI-480 $1.28\pm 0.03$ $0.16\pm 0.08$ 1 Transit 181 2.00 No No 93
TOI-509/HD-63935 $0.93\pm 0.05$ $0.07\pm 0.06$ 2 Transit 138 1.50 No No 99
TOI-544 $0.63\pm 0.02$ $-0.17\pm 0.09$ 2 Transit 108 2.90 No No 85
TOI-554/HD-25463 $1.19\pm 0.07$ $-0.20\pm 0.09$ 2 Transit 97 2.50 No No 3
TOI-561 $0.81\pm 0.03$ $-0.41\pm 0.05$ 4 Transit 102 2.90 No No 20
TOI1611/HD-207897 $0.80\pm 0.03$ $-0.04\pm 0.04$ 1 Transit 36 17.40 No No 50
TOI-669 $0.90\pm 0.06$ $-0.06\pm 0.09$ 1 Transit 61 2.10 No No 3
TOI-757 $0.80\pm 0.05$ $-0.27\pm 0.04$ 1 Transit 24 1.10 No No 5
TOI-784/HD-307842 $0.91\pm 0.10$ $-0.13\pm 0.08$ 1 Transit 35 2.10 No No 55
TOI-836 $0.68\pm 0.05$ $-0.28\pm 0.07$ 2 Transit 52 1.00 No No 47
TOI-1055/HD-183579 $1.03\pm 0.05$ $-0.07\pm 0.09$ 1 Transit 56 7.90 No No 43
TOI-1099/HD-207496 $0.80\pm 0.04$ $0.10\pm 0.04$ 1 Transit 88 1.20 No No 13
TOI-1136 $1.02\pm 0.03$ $0.07\pm 0.06$ 6 Transit 155 2.60 No No 15
TOI-1184 $0.73\pm 0.03$ $0.02\pm 0.09$ 1 Transit 59 2.10 No No 93
TOI-1235 $0.64\pm 0.02$ $0.05\pm 0.09$ 1 Transit 37 4.40 No No 23
TOI-1246 $0.87\pm 0.03$ $0.17\pm 0.06$ 2 Transit 100 2.00 No No 112
TOI-1247/HD-135694 $0.90\pm 0.06$ $-0.21\pm 0.09$ 1 Transit 72 2.30 No No 3
TOI-1249 $1.01\pm 0.03$ $0.30\pm 0.06$ 1 Transit 22 2.50 No No 93
TOI-1279 $0.88\pm 0.03$ $-0.08\pm 0.06$ 1 Transit 38 2.60 No No 93
TOI-1339/HD-191939 $0.81\pm 0.04$ $-0.15\pm 0.06$ 6 Transit 138 1.80 Yes Yes 81
TOI-1347 $0.91\pm 0.03$ $0.04\pm 0.06$ 2 Transit 120 2.10 No No 97
TOI-1411 $0.60\pm 0.02$ $0.01\pm 0.07$ 1 Transit 45 2.50 No No 93
TOI-1416 $0.80\pm 0.04$ $0.08\pm 0.05$ 1 Transit 96 2.00 No No 29
TOI-1422 $0.98\pm 0.07$ $-0.09\pm 0.07$ 1 Transit 112 1.60 No No 78
TOI-1430/HD-235088 $0.85\pm 0.02$ $0.10\pm 0.06$ 1 Transit 88 2.60 No No 93
TOI-1437 $1.01\pm 0.04$ $-0.16\pm 0.05$ 1 Transit 87 3.10 No No 91,69
TOI-1444 $0.93\pm 0.04$ $0.13\pm 0.06$ 2 Transit 56 1.00 No No 26
TOI-1451 $1.00\pm 0.04$ $0.02\pm 0.06$ 1 Transit 64 2.60 No No 93
TOI-1471/HIP-9618 $1.02\pm 0.08$ $-0.07\pm 0.06$ 2 Transit 27 10.20 No No 84
TOI-1472 $0.92\pm 0.02$ $0.29\pm 0.06$ 1 Transit 23 1.90 No No 93
TOI-1473/HD-6061 $1.00\pm 0.06$ $-0.08\pm 0.09$ 1 Transit 64 1.70 No No 3
TOI-1669 $1.00\pm 0.05$ $0.26\pm 0.06$ 2 Transit 20 2.20 Yes Yes 116
TOI-1691 $0.96\pm 0.04$ $0.05\pm 0.06$ 1 Transit 27 2.10 No No 93
TOI-1710 $0.99\pm 0.07$ $0.12\pm 0.06$ 1 Transit 70 1.20 No No 82
TOI-1723 $1.04\pm 0.04$ $0.16\pm 0.06$ 1 Transit 67 1.70 No No 93
TOI-1736 $1.08\pm 0.04$ $0.14\pm 0.01$ 2 Transit 152 2.60 Yes Yes 73
TOI-1742 $1.09\pm 0.03$ $0.19\pm 0.06$ 1 Transit 273 2.70 No No 93
TOI-1751 $0.90\pm 0.03$ $-0.40\pm 0.06$ 1 Transit 71 3.00 No No 31
TOI-1759 $0.61\pm 0.02$ $0.10\pm 0.20$ 1 Transit 204 1.20 No No 72
TOI-1778/HD-77946 $1.17\pm 0.06$ $0.18\pm 0.05$ 1 Transit 111 3.90 No No 86
TOI-1794 $0.97\pm 0.03$ $0.04\pm 0.06$ 1 Transit 24 2.10 No No 93
TOI-1798 $0.85\pm 0.02$ $0.07\pm 0.06$ 2 Transit 112 2.00 No No 93
TOI-1799 $0.94\pm 0.03$ $-0.06\pm 0.06$ 1 Transit 66 2.10 No No 93
TOI-1807 $0.76\pm 0.03$ $-0.04\pm 0.02$ 1 Transit 161 2.30 No No 80
TOI-1824 $0.88\pm 0.02$ $0.12\pm 0.06$ 1 Transit 275 2.80 No No 93
TOI-1835/HD-110067 $0.80\pm 0.04$ $-0.20\pm 0.04$ 6 Transit 72 1.90 No No 67
TOI-1836 $1.25\pm 0.04$ $-0.14\pm 0.07$ 2 Transit 53 1.80 No No 93
TOI-2076 $0.86\pm 0.02$ $0.02\pm 0.06$ 3 Transit 63 1.80 No No 93
TOI-2128 $1.03\pm 0.04$ $-0.07\pm 0.06$ 1 Transit 81 1.90 No No 93
TOI-2141 $0.94\pm 0.02$ $-0.12\pm 0.01$ 1 Transit 90 1.60 No No 73
TOI-2221/AU-Mic $0.60\pm 0.04$ $0.12\pm 0.10$ 3 Transit 225 3.10 No No 34
TOI-5140/HD-73344 $1.20\pm 0.02$ $0.14\pm 0.03$ 1 Transit 312 1.30 No No 108
 [^13] From left to right, the columns report the name of the system, the mass and metallicity of the host star, the planet multiplicity, the method by which the system was detected (either RV or transit), the number and baseline of observations of the best RV dataset used, the presence or absence of confirmed CJs in the $0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ and $0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ mass ranges (“Yes” appears in both columns if at least one CJ in the overlapping range $0.5-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ is present), the reference for the RV dataset and stellar parameters. References are given after Table LABEL:table_Mixed_sample.
Table 2: Stars of the mixed systems and their properties.
Target M [Fe/H] $N_{pl}$ Method $N_{obs}$ Baseline CJ CJ Ref. for RVs and parameters
[M] [dex] [yrs] [$0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$] [$0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$] (if different from RVs)
55-Cnc $0.98\pm 0.05$ $0.38\pm 0.06$ 5 RV 679 15.30 Yes Yes 96
HD-10180 $1.06\pm 0.05$ $0.08\pm 0.01$ 6 RV 190 6.70 No No 66
HD-99492 $0.85\pm 0.02$ $0.30\pm 0.03$ 2 RV 197 23.10 No No 96,104
HD-104067 $0.82\pm 0.03$ $0.09\pm 0.06$ 2 RV 53 14.90 No No 96
HD-107148 $1.11\pm 0.04$ $0.30\pm 0.06$ 2 RV 110 15.60 No No 96
HD-109271 $1.05\pm 0.02$ $0.10\pm 0.05$ 2 RV 106 9.50 No No 64
HIP-57274 $0.73\pm 0.05$ $0.09\pm 0.05$ 3 RV 110 5.50 Yes Yes 40
K2-19 $0.88\pm 0.03$ $0.06\pm 0.05$ 3 Transit 51 2.90 No No 89
Kepler-56 $1.67\pm 0.15$ $0.40\pm 0.04$ 3 Transit 28 10.10 Yes Yes 118,90
Kepler-88/KOI-142 $0.99\pm 0.02$ $0.27\pm 0.04$ 2 Transit 47 9.00 No No 118,90
Kepler-89/KOI-94 $1.21\pm 0.04$ $0.07\pm 0.06$ 4 Transit 71 12.70 No No 118,90
Kepler-100 $1.10\pm 0.03$ $0.10\pm 0.04$ 4 Transit 112 13.50 No No 118,90
Kepler-101 $1.17\pm 0.07$ $0.33\pm 0.07$ 2 Transit 50 6.10 No No 17
rho-CrB $0.95\pm 0.01$ $-0.20\pm 0.01$ 4 RV 587 13.70 No No 96,19
TOI-500 $0.74\pm 0.02$ $0.12\pm 0.08$ 4 Transit 197 1.00 No No 100
TOI-815 $0.78\pm 0.04$ $-0.09\pm 0.05$ 2 Transit 28 1.70 No No 94
TOI-1272 $0.85\pm 0.05$ $0.17\pm 0.06$ 2 Transit 62 1.80 No No 68
TOI-1408 $1.31\pm 0.09$ $0.25\pm 0.06$ 2 Transit 60 3.10 No No 58
TOI-1410 $0.79\pm 0.02$ $0.33\pm 0.09$ 2 Transit 56 2.70 No No 93
TOI-1726/HD-63433 $0.99\pm 0.03$ $0.08\pm 0.05$ 3 Transit 23 1.70 No No 93
TOI-2134 $0.74\pm 0.03$ $0.12\pm 0.02$ 2 Transit 108 2.20 No No 95
TOI-4010 $0.89\pm 0.03$ $0.37\pm 0.07$ 4 Transit 128 2.60 Yes Yes 59
TOI-5398 $1.15\pm 0.01$ $0.09\pm 0.06$ 2 Transit 86 1.20 No No 70
WASP-47 $1.04\pm 0.03$ $0.38\pm 0.05$ 4 Transit 142 9.20 Yes Yes 115
WASP-84 $0.84\pm 0.04$ $0.00\pm 0.10$ 2 Transit 64 5.20 No No 6
 [^14] Same as Table LABEL:table_SPsample
References: (1) Ahrer et al. (2021), (2) Akana Murphy et al. (2021), (3) Akana Murphy et al. (2023), (4) Alonso et al. (2014), (5) Alqasim et al. (2024), (6) Anderson et al. (2014), (7) Astudillo-Defru et al. (2017), (8) Astudillo-Defru et al. (2020), (9) Azevedo Silva et al. (2022), (10) Barbato et al. (2020), (11) Barnes et al. (2023), (12) Barros et al. (2022), (13) Barros et al. (2023), (14) Baycroft et al. (2023), (15) Beard et al. (2024), (16) Benedict et al. (2022), (17) Bonomo et al. (2014), (18) Bonomo et al. (2023), (19) Brewer et al. (2023), (20) Brinkman et al. (2023), (21) Bryan et al. (2019), (22) Burt et al. (2021), (23) Cloutier et al. (2020), (24) Cochran et al. (2011), (25) Courcol et al. (2015), (26) Dai et al. (2021), (27) Dalal et al. (2024), (28) Damasso et al. (2023), (29) Deeg et al. (2023), (30) Delrez et al. (2021), (31) Desai et al. (2024), (32) Díaz et al. (2016), (33) DiTomasso et al. (2023), (34) Donati et al. (2023), (35) Dumusque et al. (2019), (36) Feng et al. (2017), (37) Feng et al. (2019), (38) Feng et al. (2020), (39) Feng et al. (2022), (40) Fischer et al. (2012), (41) Frensch et al. (2023), (42) Fulton & Petigura (2018), (43) Gan et al. (2021), (44) Grunblatt et al. (2024), (45) Hara et al. (2020), (46) Hatzes et al. (2022), (47) Hawthorn et al. (2023), (48) Haywood et al. (2014), (49) Hébrard et al. (2010), (50) Heidari et al. (2022), (51) Hidalgo et al. (2020), (52) Hobson et al. (2018), (53) Hobson et al. (2024), (54) Horner et al. (2019), (55) Hua et al. (2023), (56) Jenkins et al. (2013), (57) Kane et al. (2020), (58) Korth et al. (2024), (59) Kunimoto et al. (2023), (60) Lam et al. (2018), (61) Li et al. (2024), (62) Lillo-Box et al. (2020), (63) Lillo-Box et al. (2021), (64) Lo Curto et al. (2013), (65) Lo Curto et al. (2010), (66) Lovis et al. (2011), (67) Luque et al. (2023), (68) MacDougall et al. (2022), (69) MacDougall et al. (2023), (70) Mantovan et al. (2024), (71) Marcy et al. (2014), (72) Martioli et al. (2022), (73) Martioli et al. (2023), (74) Mayo et al. (2023), (75) Ment et al. (2018), (76) Mortier et al. (2016), (77) Moutou et al. (2014), (78) Naponiello et al. (2022), (79) Naponiello, L., et al. (in prep.), (80) Nardiello et al. (2022), (81) Orell-Miquel et al. (2023), (82) Orell-Miquel et al. (2024), (83) Osborn et al. (2021), (84) Osborn et al. (2023), (85) Osborne et al. (2024), (86) Palethorpe et al. (2024), (87) Perger et al. (2017), (88) Petigura et al. (2018), (89) Petigura et al. (2020), (90) Petigura et al. (2022), (91) Pidhorodetska et al. (2024), (92) Pinamonti et al. (2022), (93) Polanski et al. (2024), (94) Psaridi et al. (2024), (95) Rescigno et al. (2024), (96) Rosenthal et al. (2021), (97) Rubenzahl et al. (2024), (98) Santos et al. (2016), (99) Scarsdale et al. (2021), (100) Serrano et al. (2022), (101) Sinukoff et al. (2017), (102) Sozzetti et al. (2021), (103) Sozzetti et al. (2024), (104) Stalport et al. (2023), (105) Stassun et al. (2017), (106) Suárez Mascareño et al. (2018), (107) Suárez Mascareño et al. (2024), (108) Sulis et al. (2024), (109) Teske et al. (2020), (110) Trifonov et al. (2020), (111) Tuomi et al. (2014), (112) Turtelboom et al. (2022), (113) Udry et al. (2019), (114) Unger et al. (2021), (115) Vanderburg et al. (2017), (116) Van Zandt et al. (2023), (117) Vissapragada et al. (2020), (118) Weiss et al. (2024), (119) Xiao et al. (2023), (120) Zechmeister et al. (2009), (121) Zhang et al. (2021)
1 Metallicity taken from the Hypatia catalog (Hinkel et al. 2014) as not provided in the reference paper.
Appendix B Sample of solar-type stars in the AAT, CLS, and HARPS radial velocity surveys
Table 3: Stars of the AAT, CLS, HARPS radial velocity surveys and their properties.
Target M [Fe/H] CJ CJ Ref.
[M] [dex] [$0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$] [$0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$]
BD-103166 $1.00$ $0.370\pm 0.060$ No No CLS
GJ 10 $1.19$ $-0.340\pm 0.100$ No No HARPS
GJ 1021 $1.08$ $-0.06\pm 0.10$ No No HARPS
GJ 1043 $1.30$ $-0.25\pm 0.10$ No No HARPS
GJ 1079 $0.80$ $-0.140\pm 0.080$ No No HARPS
GJ 1085 $1.02$ $-0.23\pm 0.80$ No No HARPS
GJ 1126 $0.71$ $-0.57\pm 0.10$ No No HARPS
GJ 1137 $0.84$ $0.080\pm 0.040$ No No HARPS
GJ 115 $1.07$ $-0.600\pm 0.040$ No No HARPS
GJ 117 $0.89$ $0.07\pm 0.20$ No No HARPS
GJ 1239 $0.72$ $-0.230\pm 0.070$ No No HARPS
GJ 1247 $0.76$ $-0.110\pm 0.060$ No No HARPS
GJ 128 $1.14$ $0.21\pm 0.20$ Yes Yes HARPS
GJ 1282 $1.24$ $-0.040\pm 0.040$ No No HARPS
GJ 132 $1.06$ $-0.470\pm 0.040$ No No HARPS
GJ 138 $1.06$ $-0.25\pm 0.30$ No No HARPS
GJ 143 $0.73$ $-0.050\pm 0.040$ No No HARPS
GJ 149 $0.72$ $-0.140\pm 0.040$ No No HARPS
GJ 159 $1.20$ $0.06\pm 0.10$ No No HARPS
GJ 17 $1.09$ $-0.21\pm 0.30$ No No HARPS
GJ 177 $1.05$ $0.01\pm 0.20$ No No HARPS
GJ 189.2 $1.30$ $-0.29\pm 0.20$ No No HARPS
GJ 2030 $0.96$ $-0.48\pm 0.10$ No No HARPS
GJ 2046 $0.79$ $-0.33\pm 0.10$ No No HARPS
GJ 214 $0.96$ $0.26\pm 0.10$ No No HARPS
GJ 225.1 $1.02$ $-0.380\pm 0.040$ No No HARPS
GJ 245.1 $1.24$ $-0.260\pm 0.040$ No No HARPS
GJ 253 $0.93$ $-0.100\pm 0.040$ No No HARPS
GJ 260 $0.95$ $0.150\pm 0.040$ No No HARPS
GJ 275 $0.99$ $0.030\pm 0.040$ No No HARPS
GJ 279 $1.20$ $-0.17\pm 0.20$ Yes Yes HARPS
GJ 288 $1.08$ $-0.81\pm 0.20$ No No HARPS
GJ 29 $1.12$ $-0.280\pm 0.100$ No No HARPS
GJ 294 A $1.10$ $-0.32\pm 0.10$ No No HARPS
GJ 3 $0.73$ $-0.690\pm 0.080$ No No HARPS
GJ 302 $0.93$ $-0.04\pm 0.10$ No No HARPS
GJ 309 $0.89$ $-0.41\pm 0.10$ No No HARPS
GJ 312 $1.11$ $0.130\pm 0.090$ No No HARPS
GJ 3187 $0.85$ $-0.460\pm 0.060$ No No HARPS
GJ 3200 $1.03$ $-0.153\pm 0.070$ No No HARPS
GJ 3222 $0.89$ $0.040\pm 0.040$ No No HARPS
GJ 3436 $1.02$ $-0.560\pm 0.080$ No No HARPS
GJ 3446 $1.05$ $-0.070\pm 0.040$ No No HARPS
GJ 3449 $1.08$ $-0.70\pm 0.20$ No No HARPS
GJ 3519 $1.30$ $0.090\pm 0.040$ No No HARPS
GJ 3535 $1.05$ $-0.260\pm 0.060$ No No HARPS
GJ 355 $0.84$ $0.330\pm 0.040$ No No HARPS
GJ 3615 $0.86$ $0.040\pm 0.080$ No No HARPS
GJ 3677 $1.21$ $-0.270\pm 0.040$ No No HARPS
GJ 3752 $0.94$ $-0.620\pm 0.060$ No No HARPS
GJ 379.1 $1.06$ $-0.17\pm 0.20$ No No HARPS
GJ 3830 $0.95$ $-0.130\pm 0.060$ No No HARPS
GJ 3944 $1.04$ $-0.270\pm 0.060$ No No HARPS
GJ 3952 $1.54$ $0.0600\pm 0.0100$ No No HARPS
GJ 398.1 $1.14$ $-0.23\pm 0.10$ No No HARPS
GJ 423.1 $0.93$ $-0.260\pm 0.080$ No No HARPS
GJ 4291 $0.79$ $-0.08\pm 0.10$ No No HARPS
GJ 4340 $1.06$ $0.30\pm 0.20$ No No HARPS
GJ 4377 $0.79$ $-0.020\pm 0.020$ No No HARPS
GJ 442 $1.01$ $-0.31\pm 0.40$ No No HARPS
GJ 446 $0.98$ $-0.290\pm 0.080$ No No HARPS
GJ 449 $1.13$ $0.13\pm 0.20$ No No HARPS
GJ 472 $0.86$ $-0.310\pm 0.060$ No No HARPS
GJ 506 $0.98$ $-0.02\pm 0.50$ No No HARPS
GJ 530 $1.01$ $-0.050\pm 0.080$ No No HARPS
GJ 541.1 $1.01$ $0.27\pm 0.20$ No No HARPS
GJ 542 $0.80$ $0.24\pm 0.20$ No No HARPS
GJ 550 $1.00$ $-0.100\pm 0.050$ No No HARPS
GJ 558.1 $1.24$ $-0.170\pm 0.040$ No No HARPS
GJ 570.1 $1.01$ $0.082\pm 0.080$ No No HARPS
GJ 570 A $0.75$ $-0.07\pm 0.20$ No No HARPS
GJ 579.2 $0.82$ $-1.40\pm 0.20$ No No HARPS
GJ 579.3 $0.89$ $-0.620\pm 0.060$ No No HARPS
GJ 58 $0.97$ $-0.410\pm 0.080$ No No HARPS
GJ 587 $0.90$ $-0.560\pm 0.020$ No No HARPS
GJ 604 $0.72$ $-0.370\pm 0.100$ No No HARPS
GJ 613.0 $0.85$ $-0.430\pm 0.060$ No No HARPS
GJ 615 $0.83$ $-1.26\pm 0.10$ No No HARPS
GJ 616 $1.04$ $0.04\pm 0.10$ No No HARPS
GJ 624 $1.14$ $-0.080\pm 0.080$ No No HARPS
GJ 634.1 $1.01$ $-0.36\pm 0.10$ No No HARPS
GJ 641 $0.94$ $-0.05\pm 0.10$ No No HARPS
GJ 656 $0.80$ $-0.710\pm 0.050$ No No HARPS
GJ 666 $0.93$ $-0.340\pm 0.030$ No No HARPS
GJ 691 $1.03$ $0.28\pm 0.20$ Yes Yes HARPS
GJ 695.1 $0.99$ $0.140\pm 0.080$ No No HARPS
GJ 699.2 $1.39$ $-0.110\pm 0.040$ No No HARPS
GJ 724.2 $1.10$ $0.030\pm 0.080$ No No HARPS
GJ 744 $1.00$ $0.08\pm 0.10$ No No HARPS
GJ 755 $1.04$ $-0.010\pm 0.080$ No No HARPS
GJ 768.1 $1.15$ $0.110\pm 0.100$ No No HARPS
GJ 776 $1.02$ $-0.26\pm 0.30$ No No HARPS
GJ 780 $0.99$ $0.33\pm 0.10$ No No HARPS
GJ 781.2 $0.74$ $-0.260\pm 0.030$ No No HARPS
GJ 785 $0.85$ $-0.01\pm 0.20$ No No HARPS
GJ 787 $0.70$ $-0.310\pm 0.050$ No No HARPS
GJ 805 $1.35$ $0.040\pm 0.040$ No No HARPS
GJ 811 $1.22$ $-0.090\pm 0.070$ No No HARPS
GJ 812.1 $1.04$ $-0.650\pm 0.060$ No No HARPS
GJ 825.2 $1.05$ $-0.002\pm 0.060$ No No HARPS
GJ 827.1 $0.79$ $-0.710\pm 0.070$ No No HARPS
GJ 836.7 $1.08$ $-0.07\pm 0.10$ No No HARPS
GJ 838 $1.08$ $-0.03\pm 0.10$ No No HARPS
GJ 845 $0.73$ $-0.13\pm 0.40$ No No HARPS
GJ 848 $1.36$ $-0.16\pm 0.10$ No No HARPS
GJ 849.1 $1.22$ $0.090\pm 0.070$ No No HARPS
GJ 853 $1.06$ $-0.24\pm 0.20$ No No HARPS
GJ 9003 $0.87$ $-0.550\pm 0.040$ No No HARPS
GJ 9007 $1.01$ $-0.300\pm 0.040$ No No HARPS
GJ 9008 $1.08$ $-0.030\pm 0.050$ No No HARPS
GJ 902 $0.76$ $-0.37\pm 0.20$ No No HARPS
GJ 9029 $1.12$ $-0.390\pm 0.060$ No No HARPS
GJ 904 $1.21$ $-0.12\pm 0.20$ No No HARPS
GJ 9075 $1.16$ $0.040\pm 0.040$ No No HARPS
GJ 9121 $1.07$ $-0.84\pm 0.20$ No No HARPS
GJ 9156 $1.02$ $-0.054\pm 0.040$ No No HARPS
GJ 9178 $0.90$ $-0.170\pm 0.040$ No No HARPS
GJ 9191 $0.88$ $-0.640\pm 0.080$ No No HARPS
GJ 9222 $1.25$ $0.090\pm 0.020$ No No HARPS
GJ 9252 $1.07$ $-0.190\pm 0.070$ No No HARPS
GJ 9255 $1.22$ $-0.060\pm 0.080$ No No HARPS
GJ 9263 $1.02$ $-0.081\pm 0.050$ No No HARPS
GJ 9269 $0.99$ $0.29\pm 0.10$ No No HARPS
GJ 9273 $1.05$ $0.13\pm 0.10$ No No HARPS
GJ 9322 $1.00$ $-0.650\pm 0.100$ No No HARPS
GJ 9324 $1.31$ $0.120\pm 0.040$ No No HARPS
GJ 9332 $1.16$ $-0.46\pm 0.30$ No No HARPS
GJ 9338 $1.33$ $0.070\pm 0.060$ No No HARPS
GJ 9356 $0.76$ $-0.090\pm 0.040$ No No HARPS
GJ 9367 $1.32$ $0.060\pm 0.080$ No No HARPS
GJ 9396 $1.01$ $0.110\pm 0.080$ No No HARPS
GJ 9467 $1.10$ $-0.150\pm 0.060$ No No HARPS
GJ 9473 $1.18$ $-0.06\pm 0.20$ No No HARPS
GJ 9476 $1.17$ $-0.710\pm 0.030$ No No HARPS
GJ 9479 $0.75$ $0.090\pm 0.020$ No No HARPS
GJ 9527 $1.01$ $0.036\pm 0.100$ No No HARPS
GJ 9630 $1.00$ $-0.05\pm 0.10$ No No HARPS
GJ 9659 $1.04$ $0.029\pm 0.090$ No No HARPS
GJ 9663 $1.27$ $0.050\pm 0.050$ No No HARPS
GJ 97 $1.07$ $-0.070\pm 0.100$ No No HARPS
GJ 9765 $0.86$ $-0.170\pm 0.040$ No No HARPS
GJ 9787 $0.98$ $-0.020\pm 0.040$ No No HARPS
GJ 9790 $0.98$ $-0.480\pm 0.060$ No No HARPS
GJ 9797 $0.95$ $-0.200\pm 0.040$ No No HARPS
GJ 9801 $1.22$ $-0.24\pm 0.20$ No No HARPS
GJ 9803 $1.00$ $-0.38\pm 0.30$ No No HARPS
GJ 9829 $1.10$ $-0.230\pm 0.080$ No No HARPS
GJ 9833 $0.90$ $-0.19\pm 0.20$ No No HARPS
GL397 $0.70$ $0.225\pm 0.070$ No No CLS
HD 10002 $0.93$ $0.251\pm 0.060$ No No CLS
HD 10008 $0.90$ $0.024\pm 0.060$ No No CLS
HD 100180 $1.06$ $0.009\pm 0.060$ No No CLS
HD 100508 $0.94$ $0.360\pm 0.030$ No No HARPS
HD 100623 $0.75$ $-0.38\pm 0.10$ No No AAT+CLS*
HD 100777 $0.97$ $0.250\pm 0.060$ Yes Yes HARPS
HD 101339 $1.02$ $-0.130\pm 0.020$ No No HARPS
HD 10145 $0.92$ $0.004\pm 0.060$ No No CLS
HD 101501 $0.92$ $-0.009\pm 0.060$ No No CLS
HD 101581 $0.74$ $-0.55\pm 0.10$ No No AAT
HD 101612 $1.24$ $-0.390\pm 0.070$ No No HARPS
HD 101644 $1.01$ $-0.590\pm 0.020$ No No HARPS
HD 101650 $0.69$ $-0.480\pm 0.020$ No No HARPS
HD 10180 $1.07$ $0.050\pm 0.080$ No No HARPS*+AAT
HD 101959 $1.12$ $-0.1100\pm 0.0100$ No No AAT
HD 102117 $1.01$ $0.274\pm 0.090$ No No HARPS*+AAT
HD 102195 $0.90$ $0.050\pm 0.040$ No No HARPS
HD 103095 $0.63$ $-1.146\pm 0.060$ No No CLS
HD 103197 $0.90$ $0.190\pm 0.020$ No No HARPS
HD 10360 $0.83$ $-0.24\pm 0.10$ No No AAT
HD 10361 $0.86$ $-0.200\pm 0.020$ No No AAT
HD 103720 $0.83$ $-0.010\pm 0.060$ No No HARPS
HD 103743 $1.00$ $-0.010\pm 0.020$ No No HARPS
HD 103774 $1.37$ $0.140\pm 0.060$ No No HARPS
HD 103891 $1.10$ $-0.220\pm 0.080$ Yes Yes HARPS
HD 103932 $0.72$ $0.117\pm 0.070$ No No CLS
HD 104006 $0.85$ $-0.78\pm 0.30$ No No HARPS
HD 104067 $0.81$ $-0.010\pm 0.100$ No No HARPS+CLS*
HD 104263 $0.96$ $-0.010\pm 0.060$ No No HARPS
HD 104304 $1.03$ $0.294\pm 0.060$ No No CLS
HD 10476 $0.84$ $0.004\pm 0.060$ No No CLS
HD 104800 $1.06$ $-0.820\pm 0.070$ No No HARPS
HD 104982 $1.01$ $-0.220\pm 0.090$ No No HARPS
HD 105 $1.12$ $0.05\pm 0.10$ No No HARPS
HD 105113 $1.05$ $-0.120\pm 0.060$ No No CLS
HD 105328 $1.09$ $0.090\pm 0.060$ No No AAT
HD 105631 $0.98$ $0.203\pm 0.060$ No No CLS
HD 105690 $0.99$ $-0.030\pm 0.020$ No No HARPS
HD 105779 $1.04$ $-0.250\pm 0.020$ Yes Yes HARPS
HD 105837 $1.08$ $-0.540\pm 0.080$ No No HARPS
HD 106275 $0.85$ $-0.110\pm 0.060$ No No HARPS
HD 106453 $1.00$ $0.050\pm 0.060$ No No AAT
HD 10647 $1.17$ $-0.03\pm 0.10$ Yes Yes HARPS*+AAT
HD 10697 $1.16$ $0.160\pm 0.060$ Yes Yes CLS
HD 10700 $0.92$ $-0.52\pm 0.20$ No No HARPS*+CLS+AAT
HD 107148 $1.03$ $0.280\pm 0.070$ No No HARPS+CLS*
HD 107692 $1.06$ $0.14\pm 0.10$ No No AAT
HD 10780 $0.90$ $0.055\pm 0.060$ No No CLS
HD 10790 $1.02$ $0.375\pm 0.060$ No No CLS
HD 108063 $1.17$ $0.520\pm 0.020$ No No HARPS
HD 108147 $1.22$ $0.17\pm 0.20$ No No HARPS*+AAT
HD 108309 $1.04$ $0.09\pm 0.10$ No No HARPS+AAT*
HD 108341 $0.89$ $0.020\pm 0.020$ Yes Yes HARPS
HD 10853 $0.72$ $-0.045\pm 0.080$ No No CLS
HD 108564 $0.74$ $-1.000\pm 0.040$ No No HARPS
HD 108874 $1.00$ $0.221\pm 0.060$ Yes Yes CLS
HD 109271 $1.05$ $0.070\pm 0.040$ No No HARPS
HD 109358 $0.90$ $-0.199\pm 0.060$ No No CLS
HD 109409 $1.03$ $0.300\pm 0.060$ No No HARPS
HD 109684 $1.10$ $-0.370\pm 0.040$ No No HARPS
HD 109988 $0.88$ $0.130\pm 0.060$ No No HARPS
HD 110315 $0.72$ $0.068\pm 0.050$ No No CLS
HD 110537 $1.02$ $0.104\pm 0.060$ No No HARPS+CLS*
HD 110557 $0.90$ $-0.090\pm 0.020$ No No HARPS
HD 110619 $0.99$ $-0.410\pm 0.020$ No No HARPS
HD 110810 $0.83$ $0.070\pm 0.040$ No No AAT
HD 110897 $0.82$ $-0.460\pm 0.060$ No No CLS
HD 111031 $1.10$ $0.272\pm 0.060$ No No CLS
HD 11112 $1.07$ $0.180\pm 0.070$ No No AAT
HD 111232 $0.96$ $-0.450\pm 0.090$ Yes Yes HARPS
HD 111395 $1.01$ $0.119\pm 0.060$ No No CLS
HD 111484 B $1.20$ $0.118\pm 0.060$ No No CLS
HD 111564 $1.11$ $0.040\pm 0.040$ No No HARPS
HD 111631 $0.68$ $0.324\pm 0.080$ No No CLS
HD 111777 $1.01$ $-0.71\pm 0.10$ No No HARPS
HD 11195 $1.02$ $-0.099\pm 0.040$ No No HARPS
HD 112257 $0.93$ $-0.011\pm 0.060$ No No CLS
HD 11226 $1.14$ $0.010\pm 0.060$ No No HARPS
HD 112914 $0.71$ $-0.254\pm 0.060$ No No CLS
HD 11397 $0.98$ $-0.570\pm 0.090$ No No HARPS
HD 114076 $0.88$ $-0.470\pm 0.020$ No No HARPS
HD 114174 $0.97$ $0.061\pm 0.060$ No No CLS
HD 114613 $1.00$ $0.16\pm 0.20$ No No HARPS+CLS*+AAT
HD 114642 $1.27$ $-0.120\pm 0.100$ No No HARPS
HD 114729 $1.05$ $-0.28\pm 0.20$ Yes Yes HARPS*+CLS
HD 114747 $0.87$ $0.210\pm 0.060$ No No HARPS
HD 114762 $0.83$ $-0.603\pm 0.060$ No No CLS
HD 114783 $0.87$ $0.151\pm 0.060$ Yes Yes CLS
HD 114853 $1.02$ $-0.240\pm 0.070$ No No HARPS*+AAT
HD 115031 $1.08$ $-0.030\pm 0.030$ No No HARPS
HD 11505 $1.03$ $-0.23\pm 0.10$ No No HARPS
HD 115169 $1.03$ $-0.043\pm 0.040$ No No HARPS
HD 115404 A $0.78$ $-0.113\pm 0.060$ No No CLS
HD 115585 $1.01$ $0.320\pm 0.060$ No No HARPS+AAT*
HD 115589 $0.92$ $0.247\pm 0.060$ No No CLS
HD 115674 $1.00$ $-0.200\pm 0.050$ No No HARPS
HD 115773 $1.21$ $-0.110\pm 0.080$ No No HARPS
HD 11608 $0.82$ $0.240\pm 0.100$ No No HARPS
HD 116321 $1.64$ $0.290\pm 0.060$ No No CLS
HD 116442 $0.76$ $-0.310\pm 0.060$ No No CLS
HD 116443 $0.74$ $-0.289\pm 0.060$ No No CLS
HD 116956 $0.96$ $0.172\pm 0.060$ No No CLS
HD 117105 $1.07$ $-0.320\pm 0.070$ No No HARPS*+AAT
HD 117176 $1.06$ $-0.040\pm 0.060$ No No CLS
HD 117207 $0.99$ $0.220\pm 0.080$ Yes Yes HARPS*+CLS
HD 117359 $0.90$ $-0.160\pm 0.020$ No No HARPS
HD 117618 $1.10$ $0.000\pm 0.080$ No No HARPS*+AAT
HD 117936 $0.79$ $0.088\pm 0.060$ No No CLS
HD 117939 $1.01$ $-0.220\pm 0.030$ No No AAT
HD 118466 $0.84$ $0.200\pm 0.020$ No No HARPS
HD 118972 $0.89$ $-0.020\pm 0.060$ No No AAT
HD 119173 $1.08$ $-0.630\pm 0.090$ No No HARPS
HD 11938 $0.71$ $0.030\pm 0.050$ No No HARPS
HD 119638 $1.14$ $-0.180\pm 0.040$ No No HARPS
HD 11964 $0.91$ $0.080\pm 0.100$ Yes Yes HARPS
HD 119949 $1.22$ $-0.440\pm 0.020$ No No HARPS
HD 120066 $1.10$ $0.101\pm 0.060$ No No CLS
HD 120136 $1.40$ $0.272\pm 0.060$ No No CLS
HD 120237 A $1.15$ $-0.0200\pm 0.0100$ No No AAT
HD 120344 $1.00$ $-0.190\pm 0.020$ No No HARPS
HD 12039 $1.01$ $-0.020\pm 0.020$ No No HARPS
HD 120467 $0.73$ $0.242\pm 0.070$ No No CLS
HD 12051 $0.97$ $0.230\pm 0.060$ No No CLS
HD 121004 $1.01$ $-0.71\pm 0.10$ No No HARPS
HD 121504 $1.12$ $0.13\pm 0.20$ No No HARPS
HD 122064 $0.80$ $0.147\pm 0.080$ No No CLS
HD 122120 $0.74$ $0.185\pm 0.070$ No No CLS
HD 122194 $1.06$ $0.060\pm 0.030$ No No HARPS
HD 12264 $1.04$ $-0.01\pm 0.10$ No No HARPS
HD 122652 $1.14$ $0.019\pm 0.060$ No No CLS
HD 123265 $0.92$ $0.160\pm 0.080$ No No HARPS
HD 123319 $1.03$ $-0.550\pm 0.020$ No No HARPS
HD 12345 $0.94$ $-0.210\pm 0.100$ No No HARPS
HD 123651 $1.10$ $-0.510\pm 0.040$ No No HARPS
HD 12387 $1.02$ $-0.270\pm 0.060$ No No HARPS
HD 124523 $1.03$ $-0.026\pm 0.030$ No No HARPS
HD 124584 $1.08$ $-0.060\pm 0.040$ No No AAT
HD 124642 $0.76$ $0.104\pm 0.080$ No No CLS
HD 125455 $0.81$ $-0.088\pm 0.060$ No No CLS
HD 125881 $1.12$ $0.030\pm 0.080$ No No HARPS*+AAT
HD 126053 $0.84$ $-0.282\pm 0.060$ No No CLS
HD 12661 $1.09$ $0.338\pm 0.060$ Yes Yes CLS
HD 126614 $1.02$ $0.408\pm 0.050$ Yes No CLS
HD 126793 $1.05$ $-0.740\pm 0.100$ No No HARPS
HD 126803 $0.96$ $-0.640\pm 0.060$ No No HARPS
HD 127124 $0.83$ $-0.030\pm 0.020$ No No HARPS
HD 127334 $1.05$ $0.229\pm 0.060$ No No CLS
HD 128165 $0.77$ $0.015\pm 0.060$ No No CLS
HD 128311 $0.84$ $0.134\pm 0.060$ Yes Yes CLS
HD 128356 $0.79$ $0.06\pm 0.40$ No No HARPS
HD 12846 $0.87$ $-0.195\pm 0.060$ No No CLS
HD 128571 $1.18$ $-0.400\pm 0.020$ No No HARPS
HD 128620 $1.11$ $0.2000\pm 0.0100$ No No AAT
HD 128621 $0.90$ $0.2400\pm 0.0100$ No No AAT
HD 128674 $0.98$ $-0.380\pm 0.060$ No No HARPS*+AAT
HD 129060 $1.16$ $0.070\pm 0.040$ No No AAT
HD 129191 $1.03$ $0.210\pm 0.050$ No No HARPS
HD 129814 $1.06$ $-0.026\pm 0.040$ No No HARPS
HD 130004 $0.75$ $-0.208\pm 0.050$ No No CLS
HD 130307 $0.78$ $-0.131\pm 0.060$ No No CLS
HD 130322 $0.93$ $0.01\pm 0.10$ No No HARPS
HD 13043 $1.09$ $0.092\pm 0.060$ No No CLS
HD 13060 $0.90$ $0.020\pm 0.060$ No No HARPS
HD 130930 $0.83$ $-0.010\pm 0.020$ No No HARPS
HD 130992 $0.76$ $-0.030\pm 0.060$ No No CLS
HD 131653 $0.92$ $-0.660\pm 0.020$ No No HARPS
HD 131664 $1.06$ $0.280\pm 0.040$ No No HARPS
HD 132142 $0.76$ $-0.294\pm 0.060$ No No CLS
HD 132569 $0.84$ $-0.260\pm 0.020$ No No HARPS
HD 132648 $0.95$ $-0.370\pm 0.060$ No No HARPS
HD 13350 $0.97$ $0.42\pm 0.20$ No No HARPS
HD 133600 $1.04$ $0.020\pm 0.030$ No No HARPS
HD 133633 $1.00$ $-0.480\pm 0.020$ No No HARPS
HD 134060 $1.09$ $0.110\pm 0.060$ No No HARPS*+AAT
HD 134088 $1.01$ $-0.780\pm 0.060$ No No HARPS
HD 134113 $1.03$ $-0.77\pm 0.10$ No No HARPS
HD 134330 $0.99$ $0.120\pm 0.080$ No No AAT
HD 134331 $1.09$ $0.000\pm 0.080$ No No AAT
HD 13445 $0.88$ $-0.23\pm 0.10$ No No AAT
HD 134606 $1.00$ $0.260\pm 0.100$ No No HARPS*+AAT
HD 134664 $1.05$ $0.070\pm 0.060$ No No HARPS
HD 134987 $1.02$ $0.26\pm 0.20$ Yes Yes HARPS+CLS*+AAT
HD 135101 A $1.01$ $0.078\pm 0.060$ No No CLS
HD 135468 $1.25$ $-0.050\pm 0.020$ No No HARPS
HD 135625 $1.12$ $0.090\pm 0.020$ Yes Yes HARPS
HD 13612B $0.99$ $0.107\pm 0.060$ No No CLS
HD 136352 $1.03$ $-0.34\pm 0.20$ No No HARPS*+CLS+AAT
HD 136713 $0.85$ $0.222\pm 0.060$ No No CLS
HD 136894 $0.94$ $-0.110\pm 0.020$ No No HARPS
HD 136925 $0.88$ $-0.217\pm 0.060$ Yes Yes CLS
HD 13724 $1.05$ $0.200\pm 0.020$ No No HARPS
HD 137388 $0.88$ $0.180\pm 0.070$ No No HARPS
HD 137778 $0.90$ $0.236\pm 0.060$ No No CLS
HD 13808 $0.85$ $-0.210\pm 0.060$ No No HARPS
HD 138549 $0.98$ $-0.030\pm 0.030$ No No HARPS*+CLS
HD 138573 $1.02$ $-0.02\pm 0.10$ No No HARPS
HD 138763 $1.14$ $0.040\pm 0.040$ No No HARPS
HD 138799 $0.89$ $-0.010\pm 0.020$ No No HARPS
HD 139189 $0.84$ $0.010\pm 0.020$ No No HARPS
HD 139211 $1.24$ $-0.24\pm 0.10$ No No HARPS
HD 13931 $1.02$ $0.041\pm 0.060$ Yes Yes CLS
HD 139323 $0.91$ $0.349\pm 0.060$ No No CLS
HD 139477 $0.74$ $0.030\pm 0.080$ No No CLS
HD 139590 $1.17$ $0.100\pm 0.040$ No No HARPS
HD 139879 $1.21$ $0.270\pm 0.020$ No No HARPS
HD 140785 $1.03$ $-0.05\pm 0.10$ No No AAT
HD 140901 $0.99$ $0.090\pm 0.080$ No No AAT
HD 141004 $1.05$ $0.025\pm 0.060$ No No CLS
HD 141272 $0.90$ $0.033\pm 0.060$ No No CLS
HD 141399 $1.09$ $0.366\pm 0.060$ Yes Yes CLS
HD 141624 $1.06$ $-0.410\pm 0.030$ No No HARPS
HD 142267 $0.83$ $-0.381\pm 0.060$ No No CLS
HD 142331 $1.01$ $0.017\pm 0.040$ No No HARPS
HD 142 A $1.21$ $0.07\pm 0.20$ Yes Yes HARPS*+AAT
HD 143114 $1.05$ $-0.440\pm 0.070$ No No HARPS*+AAT
HD 143291 $0.76$ $-0.344\pm 0.060$ No No CLS
HD 143361 $0.97$ $0.250\pm 0.080$ Yes Yes HARPS
HD 143638 $1.09$ $-0.300\pm 0.020$ No No HARPS
HD 143761 $0.91$ $-0.194\pm 0.060$ No No CLS
HD 143790 $1.34$ $-0.090\pm 0.020$ No No HARPS
HD 144009 $0.99$ $0.060\pm 0.040$ No No AAT
HD 14412 $0.77$ $-0.405\pm 0.060$ No No CLS
HD 144287 $0.87$ $-0.007\pm 0.060$ No No CLS
HD 144411 $0.79$ $-0.320\pm 0.050$ No No HARPS
HD 144550 $1.02$ $0.25\pm 0.10$ No No HARPS
HD 144579 $0.73$ $-0.529\pm 0.050$ No No CLS
HD 144585 $1.04$ $0.300\pm 0.100$ No No HARPS*+CLS
HD 144589 $1.24$ $-0.080\pm 0.020$ No No HARPS
HD 144899 $1.16$ $0.32\pm 0.10$ No No HARPS
HD 145377 $1.10$ $0.090\pm 0.040$ No No HARPS
HD 145598 $0.95$ $-0.790\pm 0.040$ No No HARPS
HD 145675 $0.97$ $0.405\pm 0.050$ Yes Yes CLS
HD 145927 $1.04$ $-0.041\pm 0.100$ No No HARPS
HD 145958 A $0.91$ $0.052\pm 0.060$ No No CLS
HD 145958 B $0.90$ $0.058\pm 0.060$ No No CLS
HD 1461 $1.03$ $0.180\pm 0.070$ No No HARPS+CLS*
HD 146481 $1.02$ $-0.460\pm 0.020$ No No AAT
HD 146775 $0.96$ $-0.055\pm 0.060$ No No CLS
HD 147231 $0.92$ $0.018\pm 0.060$ No No CLS
HD 14745 $1.21$ $-0.170\pm 0.040$ No No HARPS
HD 14747 $0.98$ $-0.390\pm 0.040$ No No HARPS
HD 147512 $0.97$ $-0.090\pm 0.030$ No No HARPS
HD 147518 $1.00$ $-0.660\pm 0.070$ No No HARPS
HD 147722 $1.10$ $0.090\pm 0.080$ No No AAT
HD 147723 $1.12$ $0.130\pm 0.040$ No No AAT
HD 147873 $1.18$ $0.08\pm 0.30$ Yes Yes HARPS
HD 148156 $1.18$ $0.220\pm 0.030$ Yes Yes HARPS
HD 148211 $1.08$ $-0.650\pm 0.080$ No No HARPS
HD 148467 $0.68$ $0.012\pm 0.070$ No No CLS
HD 148577 $1.01$ $-0.120\pm 0.020$ No No HARPS
HD 148816 $1.06$ $-0.74\pm 0.20$ No No HARPS
HD 149143 $1.06$ $0.25\pm 0.10$ No No HARPS
HD 149189 $1.08$ $0.25\pm 0.10$ No No HARPS
HD 149396 $1.01$ $0.190\pm 0.040$ No No HARPS
HD 149612 $1.02$ $-0.440\pm 0.040$ No No AAT
HD 149661 $0.89$ $0.074\pm 0.060$ No No CLS
HD 149806 $0.95$ $0.238\pm 0.060$ No No CLS
HD 150177 $1.16$ $-0.58\pm 0.10$ No No HARPS
HD 150437 $1.05$ $0.270\pm 0.050$ No No HARPS
HD 150474 $0.95$ $0.000\pm 0.060$ No No AAT
HD 151288 $0.65$ $0.056\pm 0.080$ No No CLS
HD 151337 $1.37$ $0.020\pm 0.040$ No No AAT
HD 151504 $0.95$ $0.060\pm 0.060$ No No HARPS
HD 151541 $0.83$ $-0.114\pm 0.060$ No No CLS
HD 152555 $1.10$ $0.090\pm 0.060$ No No HARPS
HD 153075 $1.05$ $-0.540\pm 0.060$ No No AAT
HD 15337 $0.88$ $0.030\pm 0.100$ No No HARPS
HD 153525 $0.77$ $-0.030\pm 0.060$ No No CLS
HD 153950 $1.11$ $-0.040\pm 0.030$ Yes Yes HARPS
HD 154088 $0.94$ $0.28\pm 0.10$ No No HARPS+CLS*
HD 154221 $1.09$ $0.120\pm 0.080$ No No HARPS
HD 154345 $0.89$ $-0.070\pm 0.060$ Yes Yes CLS
HD 154363 $0.64$ $-0.234\pm 0.070$ No No CLS
HD 154672 $1.03$ $0.250\pm 0.090$ No No HARPS
HD 154857 $0.99$ $-0.220\pm 0.040$ Yes Yes AAT
HD 154962 $1.04$ $0.290\pm 0.050$ No No HARPS
HD 155712 $0.78$ $-0.048\pm 0.050$ No No CLS
HD 155918 $1.03$ $-0.690\pm 0.060$ No No AAT
HD 155974 $1.23$ $-0.210\pm 0.060$ No No AAT
HD 156079 $1.07$ $0.250\pm 0.060$ No No HARPS
HD 156098 $1.29$ $0.150\pm 0.100$ Yes Yes HARPS
HD 15612 $0.90$ $-0.110\pm 0.020$ No No HARPS
HD 156274 B $0.60$ $-0.460\pm 0.060$ No No AAT
HD 156279 $0.96$ $0.192\pm 0.060$ Yes Yes CLS
HD 156365 $1.35$ $0.289\pm 0.060$ No No CLS
HD 156411 $1.07$ $-0.110\pm 0.100$ Yes Yes HARPS
HD 156423 $0.88$ $-0.450\pm 0.020$ No No HARPS
HD 156668 $0.79$ $0.049\pm 0.060$ No No CLS
HD 156985 $0.75$ $-0.082\pm 0.060$ No No CLS
HD 157172 $0.95$ $0.110\pm 0.040$ No No HARPS
HD 157214 $0.85$ $-0.296\pm 0.060$ No No CLS
HD 157338 $1.09$ $-0.090\pm 0.060$ No No HARPS+CLS*
HD 157347 $1.02$ $0.030\pm 0.080$ No No HARPS+CLS*
HD 157830 $0.98$ $-0.250\pm 0.050$ No No HARPS
HD 158633 $0.76$ $-0.356\pm 0.060$ No No CLS
HD 159062 $0.79$ $-0.272\pm 0.060$ No No CLS
HD 159063 $1.36$ $0.254\pm 0.060$ No No CLS
HD 159222 $1.10$ $0.166\pm 0.060$ No No CLS
HD 159868 $0.97$ $-0.07\pm 0.10$ Yes Yes AAT
HD 16008 $1.03$ $-0.090\pm 0.040$ No No HARPS
HD 160693 $0.83$ $-0.435\pm 0.060$ No No CLS
HD 160836 $0.77$ $-0.150\pm 0.020$ No No HARPS
HD 161050 $1.10$ $-0.100\pm 0.050$ No No AAT
HD 161098 $0.99$ $-0.270\pm 0.080$ No No HARPS
HD 161256 $1.01$ $0.130\pm 0.040$ No No HARPS
HD 16141 $1.03$ $0.13\pm 0.20$ No No HARPS+CLS*
HD 161566 $1.17$ $-0.310\pm 0.020$ No No HARPS
HD 16160 $0.75$ $-0.036\pm 0.070$ No No CLS
HD 161797 $1.16$ $0.318\pm 0.060$ No No CLS
HD 162396 $1.15$ $-0.35\pm 0.10$ No No HARPS
HD 16280 $0.77$ $-0.180\pm 0.020$ No No HARPS
HD 163153 $1.21$ $0.376\pm 0.060$ No No CLS
HD 163272 $1.13$ $0.2200\pm 0.0100$ No No AAT
HD 163441 $1.04$ $0.020\pm 0.050$ No No HARPS
HD 16417 $1.05$ $0.120\pm 0.100$ No No HARPS*+AAT
HD 164427 $1.07$ $0.060\pm 0.080$ No No AAT
HD 164509 $1.06$ $0.210\pm 0.040$ No No HARPS
HD 164595 $0.91$ $-0.061\pm 0.060$ No No CLS
HD 164922 $0.92$ $0.197\pm 0.060$ Yes No CLS
HD 165131 $1.06$ $0.030\pm 0.020$ No Yes HARPS
HD 165155 $0.94$ $0.210\pm 0.080$ Yes Yes HARPS
HD 165269 $1.14$ $0.270\pm 0.030$ No No AAT
HD 165401 $0.87$ $-0.322\pm 0.060$ No No CLS
HD 16548 $1.02$ $0.12\pm 0.10$ No No HARPS
HD 165920 $0.92$ $0.290\pm 0.050$ No No HARPS
HD 166 $0.98$ $0.139\pm 0.060$ No No CLS
HD 166553 $1.09$ $-0.020\pm 0.060$ No No AAT
HD 166620 $0.78$ $-0.112\pm 0.060$ No No CLS
HD 166724 $0.86$ $-0.090\pm 0.050$ Yes Yes HARPS
HD 167060 $1.05$ $-0.015\pm 0.040$ No No HARPS
HD 16714 $0.97$ $-0.200\pm 0.040$ No No HARPS
HD 167215 $1.05$ $-0.306\pm 0.060$ No No CLS
HD 167300 $1.05$ $-0.480\pm 0.080$ No No HARPS
HD 167665 $1.06$ $-0.141\pm 0.060$ No No CLS
HD 167677 $0.96$ $-0.320\pm 0.100$ Yes Yes HARPS
HD 168009 $0.99$ $0.018\pm 0.060$ No No CLS
HD 168060 $0.98$ $0.280\pm 0.040$ No No AAT
HD 168443 $1.04$ $0.076\pm 0.060$ Yes Yes CLS
HD 168746 $0.98$ $-0.10\pm 0.20$ No No HARPS
HD 168769 $0.91$ $-0.010\pm 0.040$ No No HARPS
HD 168863 $0.80$ $0.180\pm 0.020$ No No HARPS
HD 168871 $1.10$ $-0.120\pm 0.080$ No No HARPS*+AAT
HD 16905 $0.79$ $0.170\pm 0.080$ Yes Yes HARPS
HD 169830 $1.17$ $0.17\pm 0.20$ Yes Yes HARPS+CLS*
HD 170469 $1.14$ $0.311\pm 0.060$ Yes Yes CLS
HD 170493 $0.79$ $0.206\pm 0.080$ No No CLS
HD 17051 $1.17$ $0.16\pm 0.20$ No No AAT
HD 170657 $0.79$ $-0.126\pm 0.060$ No No CLS
HD 171028 $1.01$ $-0.480\pm 0.100$ Yes Yes HARPS
HD 171067 $0.95$ $-0.012\pm 0.060$ No No CLS
HD 171587 $0.95$ $-0.640\pm 0.020$ No No HARPS
HD 171834 $1.46$ $-0.100\pm 0.040$ No No HARPS
HD 17190 $0.84$ $-0.026\pm 0.060$ No No CLS
HD 172051 $1.00$ $-0.207\pm 0.060$ No No CLS*+AAT
HD 17230 $0.71$ $0.166\pm 0.070$ No No CLS
HD 172513 $0.97$ $-0.050\pm 0.080$ No No HARPS+CLS*
HD 172568 $1.02$ $-0.390\pm 0.020$ No No HARPS
HD 173701 $1.00$ $0.353\pm 0.060$ No No CLS
HD 17382 $0.91$ $0.103\pm 0.050$ No No CLS
HD 174080 $0.76$ $0.112\pm 0.080$ No No CLS
HD 17439 $1.04$ $0.050\pm 0.020$ No No HARPS
HD 174545 $0.88$ $0.190\pm 0.080$ No No HARPS
HD 17548 $1.11$ $-0.560\pm 0.070$ No No HARPS
HD 175607 $0.94$ $-0.620\pm 0.060$ No No HARPS
HD 176377 $0.91$ $-0.229\pm 0.060$ No No CLS
HD 176986 $0.82$ $0.050\pm 0.060$ No No HARPS
HD 177153 $1.04$ $-0.049\pm 0.060$ No No CLS
HD 177409 $1.07$ $-0.070\pm 0.040$ No No HARPS
HD 177758 $1.07$ $-0.610\pm 0.080$ No No HARPS
HD 178911 B $1.03$ $0.260\pm 0.060$ No No CLS
HD 179346 $1.16$ $-0.060\pm 0.050$ No No HARPS
HD 179949 $1.17$ $0.19\pm 0.20$ No No HARPS+CLS+AAT*
HD 179957 $0.94$ $0.009\pm 0.060$ No No CLS
HD 179958 $0.97$ $0.039\pm 0.060$ No No CLS
HD 180161 $0.97$ $0.184\pm 0.060$ No No CLS
HD 180409 $1.11$ $-0.200\pm 0.040$ No No HARPS
HD 181327 $1.32$ $-0.17\pm 0.10$ No No HARPS
HD 181428 $1.13$ $0.03\pm 0.10$ No No AAT
HD 18143 $0.83$ $0.168\pm 0.060$ No No CLS
HD 181433 $0.81$ $0.38\pm 0.10$ Yes Yes HARPS
HD 181720 $1.04$ $-0.560\pm 0.030$ Yes No HARPS
HD 182488 $0.95$ $0.199\pm 0.060$ No No CLS
HD 182572 $1.10$ $0.370\pm 0.060$ No No CLS
HD 183263 $1.17$ $0.278\pm 0.060$ Yes Yes CLS
HD 183414 $1.02$ $-0.020\pm 0.040$ No No HARPS
HD 183579 $1.04$ $-0.036\pm 0.030$ No No HARPS
HD 183650 $1.14$ $0.309\pm 0.060$ No No CLS
HD 183870 $0.82$ $0.019\pm 0.060$ No No CLS
HD 183877 $0.99$ $-0.200\pm 0.060$ No No AAT
HD 185144 $0.81$ $-0.172\pm 0.060$ No No CLS
HD 185283 $0.79$ $-0.070\pm 0.060$ Yes Yes HARPS
HD 185414 $0.95$ $-0.095\pm 0.060$ No No CLS
HD 185615 $0.98$ $0.080\pm 0.030$ No No HARPS
HD 186194 $1.03$ $0.190\pm 0.080$ No No HARPS
HD 18632 $0.89$ $0.230\pm 0.060$ No No CLS
HD 186408 $1.02$ $0.089\pm 0.060$ No No CLS
HD 186427 $0.98$ $0.074\pm 0.060$ Yes Yes CLS
HD 187085 $1.15$ $0.090\pm 0.100$ Yes Yes AAT
HD 187123 $1.02$ $0.106\pm 0.060$ Yes Yes CLS
HD 187237 $1.04$ $0.065\pm 0.060$ No No CLS
HD 187532 $1.48$ $-0.140\pm 0.040$ No No HARPS
HD 187897 $1.11$ $0.129\pm 0.060$ No No CLS
HD 187923 $0.98$ $-0.087\pm 0.060$ No No CLS
HD 188015 $1.04$ $0.253\pm 0.060$ Yes Yes CLS
HD 18803 $0.98$ $0.128\pm 0.060$ No No CLS
HD 188748 $1.00$ $-0.120\pm 0.040$ No No HARPS
HD 188815 $1.21$ $-0.56\pm 0.10$ No No HARPS
HD 189625 $1.06$ $0.178\pm 0.030$ No No HARPS+CLS*
HD 189733 $0.83$ $0.041\pm 0.060$ No No CLS
HD 190007 $0.76$ $0.209\pm 0.070$ No No CLS
HD 190067 $0.79$ $-0.288\pm 0.060$ No No CLS
HD 190125 $1.03$ $0.170\pm 0.080$ No No HARPS
HD 19034 $0.80$ $-0.354\pm 0.060$ No No CLS
HD 190360 $0.99$ $0.230\pm 0.060$ Yes Yes CLS
HD 190404 $0.70$ $-0.476\pm 0.050$ No No CLS
HD 190406 $1.07$ $0.051\pm 0.060$ No No CLS
HD 190647 $1.00$ $0.230\pm 0.100$ Yes Yes HARPS
HD 190954 $0.95$ $-0.410\pm 0.020$ No No HARPS
HD 190984 $1.10$ $-0.520\pm 0.020$ Yes Yes HARPS
HD 191408 $0.84$ $-0.522\pm 0.060$ No No CLS+AAT*
HD 191785 $0.81$ $-0.044\pm 0.050$ No No CLS
HD 192020 $0.85$ $0.016\pm 0.060$ No No CLS
HD 192031 $0.89$ $-0.870\pm 0.050$ No No HARPS
HD 192263 $0.82$ $0.03\pm 0.10$ No No HARPS+CLS*
HD 19308 $1.06$ $0.150\pm 0.060$ No No CLS
HD 193202 $0.70$ $0.235\pm 0.070$ No No CLS
HD 193307 $1.12$ $-0.290\pm 0.080$ No No AAT
HD 19373 $1.14$ $0.116\pm 0.060$ No No CLS
HD 194490 $1.06$ $0.070\pm 0.080$ No No HARPS
HD 194640 $0.98$ $-0.040\pm 0.020$ No No AAT
HD 195019 $1.05$ $0.050\pm 0.060$ No No CLS
HD 195145 $1.05$ $0.170\pm 0.060$ No No HARPS
HD 19518 $1.04$ $-0.11\pm 0.10$ No No HARPS
HD 195564 $1.12$ $0.089\pm 0.060$ No No CLS
HD 196050 $1.07$ $0.22\pm 0.20$ Yes Yes HARPS+AAT*
HD 196068 $1.08$ $0.200\pm 0.080$ No No AAT
HD 19632 $1.03$ $0.12\pm 0.20$ No No AAT
HD 196378 $1.13$ $-0.43\pm 0.20$ No No HARPS+AAT*
HD 196390 $1.07$ $0.057\pm 0.060$ No No HARPS
HD 19641 $1.06$ $-0.040\pm 0.020$ No No HARPS
HD 19668 $0.95$ $-0.040\pm 0.040$ No No HARPS
HD 196761 $0.96$ $-0.197\pm 0.060$ No No CLS*+AAT
HD 196800 $1.10$ $0.160\pm 0.060$ No No HARPS*+AAT
HD 196850 $0.92$ $-0.094\pm 0.060$ No No CLS
HD 196892 $1.08$ $-1.05\pm 0.30$ No No HARPS
HD 197027 $1.02$ $-0.016\pm 0.030$ No No HARPS
HD 197076 $0.98$ $-0.065\pm 0.060$ No No CLS
HD 197083 $1.03$ $-0.480\pm 0.050$ No No HARPS
HD 197197 $1.04$ $-0.490\pm 0.020$ No No HARPS
HD 197536 $1.13$ $-0.440\pm 0.060$ No No HARPS
HD 197823 $0.93$ $0.110\pm 0.080$ No No HARPS
HD 1979 $1.00$ $-0.090\pm 0.020$ No No HARPS
HD 198075 $1.05$ $-0.270\pm 0.060$ No No HARPS
HD 198390 $1.28$ $-0.32\pm 0.10$ No No HARPS
HD 198425 $0.79$ $-0.039\pm 0.060$ No No CLS
HD 198802 $1.24$ $0.009\pm 0.060$ No No CLS
HD 199190 $1.08$ $0.140\pm 0.070$ No No HARPS+AAT*
HD 199289 $1.06$ $-1.04\pm 0.10$ No No HARPS
HD 199476 $0.80$ $-0.311\pm 0.060$ No No CLS
HD 199509 $1.04$ $-0.310\pm 0.050$ No No AAT
HD 199604 $1.05$ $-0.650\pm 0.060$ No No HARPS
HD 199847 $1.03$ $-0.570\pm 0.020$ No No HARPS
HD 199960 $1.07$ $0.25\pm 0.20$ No No HARPS+CLS*
HD 20003 $0.96$ $0.040\pm 0.060$ No No HARPS
HD 200083 $0.78$ $-0.120\pm 0.020$ No No HARPS
HD 200143 $0.86$ $0.010\pm 0.020$ No No HARPS
HD 20029 $1.17$ $0.120\pm 0.060$ No No AAT
HD 200349 $0.78$ $-0.290\pm 0.020$ No No HARPS
HD 200538 $1.11$ $0.098\pm 0.020$ No No HARPS
HD 200565 $1.06$ $0.049\pm 0.060$ No No CLS
HD 200633 $1.05$ $0.040\pm 0.040$ No No HARPS
HD 200968 $0.87$ $0.096\pm 0.060$ No No CLS
HD 201092 $0.62$ $-0.395\pm 0.050$ No No CLS
HD 20165 $0.84$ $0.023\pm 0.060$ No No CLS
HD 20201 $1.10$ $0.080\pm 0.060$ No No AAT
HD 202206 $1.03$ $0.29\pm 0.20$ No Yes HARPS
HD 202575 $0.74$ $-0.074\pm 0.060$ No No CLS
HD 202751 $0.75$ $-0.024\pm 0.070$ No No CLS
HD 202871 $1.11$ $-0.120\pm 0.040$ No No HARPS
HD 203384 $0.98$ $0.230\pm 0.050$ No No HARPS
HD 203432 $0.99$ $0.278\pm 0.060$ No No HARPS
HD 2039 $1.09$ $0.31\pm 0.20$ Yes Yes AAT
HD 204287 $1.05$ $-0.070\pm 0.020$ No No AAT
HD 204313 $1.03$ $0.160\pm 0.040$ Yes Yes HARPS
HD 204385 $1.12$ $0.040\pm 0.090$ No No HARPS*+AAT
HD 204587 $0.67$ $0.071\pm 0.070$ No No CLS
HD 204941 $0.83$ $-0.200\pm 0.050$ No No HARPS
HD 205294 $1.23$ $-0.280\pm 0.060$ No No HARPS
HD 205390 $0.84$ $-0.170\pm 0.040$ No No AAT
HD 205536 $0.95$ $-0.050\pm 0.030$ No No HARPS
HD 205591 $1.32$ $-0.110\pm 0.020$ No No HARPS
HD 205739 $1.24$ $0.180\pm 0.040$ No No HARPS
HD 206172 $0.99$ $-0.240\pm 0.040$ No No HARPS
HD 20619 $0.91$ $-0.172\pm 0.060$ No No CLS
HD 206387 $1.15$ $0.226\pm 0.060$ No No CLS
HD 206395 $1.20$ $0.200\pm 0.040$ No No AAT
HD 206683 $1.14$ $0.290\pm 0.080$ No No HARPS
HD 20675 $1.54$ $0.162\pm 0.060$ No No CLS
HD 206998 $1.05$ $-0.720\pm 0.080$ No No HARPS
HD 2071 $1.02$ $-0.090\pm 0.040$ No No HARPS
HD 207190 $1.18$ $-0.450\pm 0.050$ No No HARPS
HD 20766 $1.03$ $-0.23\pm 0.20$ No No AAT
HD 207700 $1.01$ $0.030\pm 0.040$ No No HARPS*+AAT
HD 20781 $0.90$ $-0.110\pm 0.040$ No No HARPS
HD 20782 $1.04$ $-0.060\pm 0.080$ Yes Yes HARPS*+AAT
HD 207832 $1.02$ $0.130\pm 0.040$ Yes Yes HARPS
HD 207869 $0.98$ $-0.450\pm 0.020$ No No HARPS
HD 20794 $0.94$ $-0.40\pm 0.20$ No No AAT
HD 208 $1.08$ $-0.340\pm 0.030$ No No HARPS
HD 208038 $0.81$ $-0.028\pm 0.060$ No No CLS
HD 208272 $0.89$ $-0.070\pm 0.060$ No No HARPS
HD 208313 $0.82$ $0.010\pm 0.060$ No No CLS
HD 208487 $1.17$ $0.050\pm 0.090$ No No HARPS*+AAT
HD 20852 $1.45$ $-0.380\pm 0.020$ No No HARPS
HD 20868 $0.78$ $0.070\pm 0.020$ No No HARPS
HD 208704 $1.05$ $-0.096\pm 0.050$ No No HARPS
HD 208998 $1.08$ $-0.330\pm 0.100$ No No AAT
HD 209268 $1.13$ $-0.080\pm 0.040$ No No AAT
HD 209458 $1.07$ $-0.006\pm 0.060$ No No CLS
HD 209653 $1.10$ $-0.140\pm 0.040$ No No AAT
HD 209875 $1.13$ $-0.087\pm 0.060$ No No CLS
HD 210277 $0.96$ $0.20\pm 0.10$ Yes Yes HARPS+CLS*
HD 210312 $1.07$ $0.253\pm 0.060$ No No CLS
HD 210918 $1.16$ $-0.110\pm 0.070$ No No AAT
HD 210975 $0.76$ $-0.460\pm 0.060$ No No HARPS
HD 211080 $1.26$ $0.330\pm 0.060$ No No CLS
HD 211188 $0.84$ $-0.150\pm 0.030$ No No HARPS
HD 211317 $1.09$ $0.240\pm 0.100$ No No AAT
HD 21132 $1.25$ $-0.400\pm 0.040$ No No HARPS
HD 211976 $1.34$ $-0.110\pm 0.040$ No No HARPS
HD 211998 $1.45$ $-1.51\pm 0.20$ No No AAT
HD 212036 $1.04$ $-0.010\pm 0.020$ No No HARPS
HD 21209 A $0.74$ $-0.440\pm 0.040$ No No HARPS
HD 212168 $1.08$ $-0.010\pm 0.040$ No No AAT
HD 212291 $0.90$ $-0.119\pm 0.060$ No No CLS
HD 212330 $1.01$ $-0.06\pm 0.20$ No No AAT
HD 212708 $1.01$ $0.255\pm 0.090$ No No HARPS+AAT*
HD 212918 $0.84$ $-0.240\pm 0.020$ No No HARPS
HD 213042 $0.79$ $0.199\pm 0.080$ No No CLS
HD 213240 $1.10$ $0.14\pm 0.20$ Yes Yes AAT
HD 213472 $1.05$ $0.125\pm 0.060$ No No CLS
HD 213575 $1.01$ $-0.17\pm 0.20$ No No HARPS
HD 214094 $1.19$ $-0.040\pm 0.060$ No No HARPS
HD 214385 $1.01$ $-0.370\pm 0.060$ No No HARPS
HD 214683 $0.71$ $-0.262\pm 0.060$ No No CLS
HD 214749 $0.72$ $0.088\pm 0.070$ No No CLS
HD 214953 A $1.15$ $-0.050\pm 0.050$ No No AAT
HD 2151 $1.33$ $-0.12\pm 0.20$ No No AAT
HD 215257 $1.06$ $-0.64\pm 0.10$ No No HARPS
HD 215497 $0.86$ $0.170\pm 0.020$ Yes No HARPS
HD 215641 $0.97$ $0.020\pm 0.040$ No No HARPS
HD 215906 $1.21$ $-0.310\pm 0.020$ No No HARPS
HD 216259 $0.68$ $-0.556\pm 0.040$ No No CLS
HD 216435 $1.10$ $0.21\pm 0.20$ Yes Yes HARPS+AAT*
HD 216437 $1.06$ $0.22\pm 0.20$ Yes Yes HARPS+AAT*
HD 216520 $0.79$ $-0.124\pm 0.060$ No No CLS
HD 216770 $0.94$ $0.240\pm 0.100$ No No HARPS
HD 21693 $0.95$ $-0.030\pm 0.050$ No No HARPS
HD 217004 $1.04$ $-0.030\pm 0.060$ No No CLS
HD 217014 $1.07$ $0.206\pm 0.060$ No No CLS
HD 217107 $1.06$ $0.340\pm 0.060$ Yes Yes CLS
HD 217165 $1.07$ $0.044\pm 0.060$ No No CLS
HD 217343 $1.03$ $0.00\pm 0.10$ No No HARPS
HD 217357 $0.63$ $0.025\pm 0.070$ No No CLS
HD 217786 $1.14$ $-0.170\pm 0.030$ No Yes HARPS
HD 217877 $1.00$ $-0.089\pm 0.060$ No No CLS
HD 217958 $1.08$ $0.25\pm 0.10$ No No AAT
HD 218209 $0.81$ $-0.376\pm 0.060$ No No CLS
HD 218249 $0.82$ $-0.430\pm 0.060$ No No HARPS
HD 218379 $1.10$ $0.120\pm 0.020$ No No HARPS
HD 218504 $1.09$ $-0.580\pm 0.080$ No No HARPS
HD 218544 $1.04$ $-0.017\pm 0.040$ No No HARPS
HD 218566 $0.80$ $0.28\pm 0.10$ No No HARPS+CLS*
HD 218860 $0.98$ $0.110\pm 0.020$ No No HARPS
HD 218868 $1.00$ $0.231\pm 0.060$ No No CLS
HD 219057 $1.05$ $-0.063\pm 0.040$ No No HARPS
HD 219077 $0.92$ $-0.16\pm 0.10$ Yes Yes HARPS+AAT*
HD 219134 $0.79$ $0.083\pm 0.060$ Yes No CLS
HD 219249 $0.96$ $-0.400\pm 0.060$ No No HARPS
HD 21938 $1.04$ $-0.480\pm 0.040$ No No HARPS
HD 219538 $0.83$ $0.008\pm 0.060$ No No CLS
HD 219542 $1.04$ $0.165\pm 0.060$ No No CLS
HD 219828 $1.04$ $0.15\pm 0.10$ No Yes HARPS
HD 219834 B $0.90$ $0.224\pm 0.060$ No No CLS
HD 220221 $0.81$ $0.161\pm 0.060$ No No CLS
HD 220339 $0.75$ $-0.213\pm 0.060$ No No CLS
HD 220367 $1.16$ $-0.240\pm 0.020$ No No HARPS
HD 22049 $0.85$ $-0.09\pm 0.50$ Yes Yes HARPS+CLS*
HD 220507 $1.01$ $0.000\pm 0.060$ No No HARPS
HD 22104 $1.04$ $0.320\pm 0.100$ No No AAT
HD 221146 $1.07$ $0.080\pm 0.040$ No No HARPS
HD 221287 $1.27$ $0.01\pm 0.10$ Yes Yes HARPS
HD 221354 $0.86$ $0.099\pm 0.060$ No No CLS
HD 221575 $0.85$ $-0.060\pm 0.080$ No No HARPS
HD 221851 $0.84$ $-0.025\pm 0.060$ No No CLS
HD 222480 $1.05$ $0.150\pm 0.020$ No No HARPS+AAT*
HD 222582 $1.04$ $-0.01\pm 0.20$ Yes Yes HARPS+CLS*
HD 222595 $1.00$ $0.010\pm 0.040$ No No HARPS
HD 222669 $1.07$ $0.030\pm 0.050$ No No HARPS
HD 223121 $0.85$ $0.060\pm 0.040$ No No HARPS
HD 223171 $1.05$ $0.110\pm 0.090$ No No HARPS*+AAT
HD 223238 $1.06$ $0.024\pm 0.060$ No No HARPS+CLS*
HD 223282 $0.92$ $-0.430\pm 0.040$ No No HARPS
HD 223498 $0.97$ $0.184\pm 0.060$ No No CLS
HD 223691 $1.06$ $-0.169\pm 0.060$ No No CLS
HD 223854 $1.14$ $-0.570\pm 0.080$ No No HARPS
HD 224383 $1.04$ $-0.036\pm 0.080$ No No HARPS
HD 224393 $1.04$ $-0.410\pm 0.030$ No No HARPS
HD 224538 $1.23$ $0.28\pm 0.10$ Yes Yes HARPS
HD 224619 $0.96$ $-0.220\pm 0.080$ No No HARPS*+CLS
HD 224685 $0.97$ $-0.400\pm 0.030$ No No HARPS
HD 224693 $1.07$ $0.250\pm 0.070$ No No HARPS
HD 224789 $0.88$ $-0.040\pm 0.040$ No No HARPS
HD 224817 $1.07$ $-0.560\pm 0.060$ No No HARPS
HD 224983 $0.88$ $0.147\pm 0.060$ No No CLS
HD 225261 $0.77$ $-0.323\pm 0.060$ No No CLS
HD 23079 $1.10$ $-0.14\pm 0.10$ Yes Yes HARPS+AAT*
HD 23127 $1.05$ $0.34\pm 0.20$ Yes Yes AAT
HD 231701 $1.15$ $-0.010\pm 0.080$ No No HARPS
HD 23249 $1.22$ $0.12\pm 0.60$ No No AAT
HD 23356 $0.82$ $-0.10\pm 0.20$ No No HARPS+CLS*
HD 23439 $0.68$ $-0.733\pm 0.040$ No No CLS
HD 23456 $1.18$ $-0.350\pm 0.080$ No No HARPS
HD 23901 $0.91$ $-0.400\pm 0.020$ No No HARPS
HD 24040 $1.10$ $0.197\pm 0.060$ Yes Yes CLS
HD 24062 $1.12$ $0.250\pm 0.070$ No No HARPS
HD 24112 $1.16$ $0.130\pm 0.040$ No No HARPS
HD 24238 $0.72$ $-0.357\pm 0.040$ No No CLS
HD 24496 $0.93$ $0.030\pm 0.060$ No No CLS
HD 245409 $0.62$ $0.066\pm 0.080$ No No CLS
HD 24558 $0.92$ $-0.490\pm 0.040$ No No HARPS
HD 24712 $1.63$ $-5.100\pm 0.040$ No No HARPS
HD 24892 $0.95$ $-0.271\pm 0.060$ No No CLS
HD 24916 $0.69$ $-0.140\pm 0.100$ No No HARPS
HD 25061 $0.91$ $0.070\pm 0.070$ No No HARPS
HD 25171 $1.20$ $-0.140\pm 0.040$ Yes Yes HARPS
HD 25665 $0.81$ $0.021\pm 0.060$ No No CLS
HD 2567 $1.14$ $0.190\pm 0.020$ No No HARPS
HD 25673 $0.87$ $-0.510\pm 0.080$ No No HARPS
HD 25912 $1.08$ $0.090\pm 0.040$ No No HARPS
HD 26151 $0.98$ $0.308\pm 0.060$ No No CLS
HD 26161 $1.13$ $-0.009\pm 0.060$ No No CLS
HD 2638 $0.88$ $0.120\pm 0.060$ No No HARPS
HD 26736 $1.10$ $0.194\pm 0.060$ No No CLS
HD 26887 $1.11$ $-0.380\pm 0.030$ No No HARPS
HD 26923 $1.08$ $-0.030\pm 0.040$ No No HARPS
HD 26965 $0.85$ $-0.34\pm 0.20$ No No HARPS*+CLS+AAT
HD 27063 $1.04$ $0.05\pm 0.10$ No No HARPS
HD 27274 $0.71$ $-0.15$ No No AAT
HD 27442 A $1.21$ $0.39\pm 0.20$ Yes Yes AAT
HD 2768 $0.98$ $-0.060\pm 0.020$ No No HARPS
HD 27894 $0.81$ $0.26\pm 0.20$ Yes Yes HARPS
HD 28005 $1.13$ $0.313\pm 0.060$ No No CLS
HD 28185 $1.02$ $0.227\pm 0.060$ Yes Yes CLS
HD 28254 $1.00$ $0.330\pm 0.020$ Yes Yes HARPS
HD 28255 A $1.07$ $0.050\pm 0.040$ No No AAT
HD 28255b $1.00$ $0.090\pm 0.040$ No No AAT
HD 28343 $0.70$ $0.375\pm 0.070$ No No CLS
HD 28635 $1.15$ $0.100\pm 0.020$ No No HARPS
HD 28701 $1.02$ $-0.320\pm 0.060$ No No HARPS
HD 28821 $1.01$ $-0.120\pm 0.050$ No No HARPS
HD 28946 $0.85$ $-0.101\pm 0.060$ No No CLS
HD 290038 $0.83$ $0.030\pm 0.020$ No No HARPS
HD 290327 $0.97$ $-0.170\pm 0.020$ Yes Yes HARPS
HD 29137 $1.03$ $0.270\pm 0.020$ No No HARPS
HD 297396 $0.76$ $0.080\pm 0.040$ No No HARPS
HD 29883 $0.78$ $-0.094\pm 0.050$ No No CLS
HD 30053 $1.19$ $-0.250\pm 0.060$ No No HARPS
HD 30177 $0.98$ $0.37\pm 0.20$ Yes Yes AAT
HD 30295 $0.94$ $0.300\pm 0.090$ No No AAT
HD 30306 $0.98$ $0.173\pm 0.060$ No No HARPS
HD 30669 $0.94$ $0.130\pm 0.040$ Yes No HARPS
HD 3074 A $1.14$ $0.056\pm 0.060$ No No CLS
HD 30876 $0.84$ $-0.060\pm 0.040$ No No AAT
HD 31103 $1.14$ $0.060\pm 0.020$ No No HARPS
HD 31253 $1.33$ $0.140\pm 0.060$ Yes No CLS
HD 31527 $1.07$ $-0.200\pm 0.070$ No No HARPS*+AAT
HD 31532 $1.07$ $-0.110\pm 0.020$ No No HARPS
HD 31560 $0.77$ $0.130\pm 0.080$ No No CLS
HD 31822 $1.13$ $-0.220\pm 0.040$ No No HARPS
HD 31827 $0.98$ $0.44\pm 0.10$ No No AAT
HD 32147 $0.80$ $0.186\pm 0.080$ No No CLS
HD 32564 $0.98$ $-0.010\pm 0.030$ No No HARPS
HD 32724 $1.05$ $-0.200\pm 0.060$ No No HARPS
HD 32923 $1.00$ $-0.140\pm 0.060$ No No CLS
HD 32963 $0.98$ $0.075\pm 0.060$ Yes Yes CLS
HD 33473 $1.03$ $-0.13\pm 0.10$ No No HARPS
HD 33636 $1.02$ $-0.071\pm 0.060$ No No CLS
HD 34411 $1.03$ $0.074\pm 0.060$ No No CLS
HD 34445 $1.11$ $0.138\pm 0.060$ Yes Yes CLS
HD 34721 $1.05$ $-0.046\pm 0.060$ No No CLS
HD 36003 $0.74$ $0.155\pm 0.070$ No No CLS
HD 36108 $1.08$ $-0.240\pm 0.060$ No No AAT
HD 36152 $1.03$ $0.057\pm 0.060$ No No HARPS
HD 36379 $1.12$ $-0.170\pm 0.040$ No No HARPS
HD 36395 $0.62$ $0.397\pm 0.060$ No No CLS
HD 3651 $0.89$ $0.168\pm 0.060$ No No CLS
HD 37008 $0.73$ $-0.322\pm 0.060$ No No CLS
HD 37124 $0.81$ $-0.326\pm 0.060$ Yes Yes CLS
HD 37213 $0.93$ $-0.419\pm 0.060$ No No CLS
HD 37226 $1.18$ $-0.150\pm 0.020$ No No HARPS
HD 37394 $0.92$ $0.159\pm 0.060$ No No CLS
HD 37572 $0.86$ $-0.110\pm 0.040$ No No HARPS
HD 3765 $0.85$ $0.172\pm 0.060$ No No CLS
HD 377 $1.11$ $0.138\pm 0.060$ No No CLS
HD 3795 $0.86$ $-0.513\pm 0.060$ No No CLS
HD 37990 $1.18$ $-0.030\pm 0.050$ No No HARPS
HD 3821 $1.06$ $-0.060\pm 0.040$ No No HARPS
HD 38230 $0.83$ $-0.006\pm 0.050$ No No CLS
HD 38283 $1.10$ $-0.250\pm 0.100$ Yes No HARPS+AAT*
HD 38382 $1.12$ $0.000\pm 0.100$ No No AAT
HD 38459 $0.90$ $0.210\pm 0.080$ No No HARPS
HD 38677 $1.17$ $0.180\pm 0.040$ No No HARPS
HD 38973 $1.11$ $0.030\pm 0.090$ No No AAT
HD 38B $0.62$ $0.054\pm 0.080$ No No CLS
HD 39091 $1.10$ $0.07\pm 0.10$ Yes Yes HARPS*+AAT
HD 39427 $1.01$ $-0.180\pm 0.020$ No No HARPS
HD 3964 $1.04$ $0.050\pm 0.020$ No No HARPS
HD 39881 $0.92$ $-0.091\pm 0.060$ No No CLS
HD 40397 $0.88$ $-0.042\pm 0.060$ No No CLS
HD 40979 $1.24$ $0.203\pm 0.060$ No No CLS
HD 41248 $1.04$ $-0.400\pm 0.020$ No No HARPS
HD 41593 $0.91$ $0.087\pm 0.060$ No No CLS
HD 4203 $1.14$ $0.380\pm 0.060$ Yes Yes CLS
HD 4208 $0.85$ $-0.220\pm 0.060$ Yes Yes CLS
HD 42250 $0.88$ $0.047\pm 0.060$ No No CLS
HD 4256 $0.85$ $0.256\pm 0.060$ No No CLS
HD 42618 $1.03$ $-0.09\pm 0.10$ No No HARPS+CLS*
HD 42936 $0.86$ $0.19\pm 0.20$ No No HARPS
HD 4307 $1.02$ $-0.199\pm 0.060$ No No CLS
HD 4308 $1.01$ $-0.340\pm 0.090$ No No HARPS*+AAT
HD 43197 $0.96$ $0.360\pm 0.040$ Yes Yes HARPS
HD 43834 $0.99$ $0.10\pm 0.10$ No No AAT
HD 43947 $0.92$ $-0.255\pm 0.060$ No No CLS
HD 44219 $1.03$ $0.020\pm 0.020$ Yes Yes HARPS
HD 44420 $1.03$ $0.265\pm 0.070$ No No HARPS
HD 44447 $1.11$ $-0.25\pm 0.10$ No No HARPS*+AAT
HD 4457 $0.80$ $-0.380\pm 0.020$ No No HARPS
HD 44573 $0.84$ $-0.060\pm 0.040$ No No HARPS
HD 44594 $1.05$ $0.130\pm 0.070$ No No HARPS*+AAT
HD 44627 $0.84$ $-0.120\pm 0.090$ No No HARPS
HD 44665 $1.02$ $-0.004\pm 0.050$ No No HARPS
HD 44804 $0.93$ $0.030\pm 0.020$ No No HARPS
HD 45021 $1.01$ $0.007\pm 0.050$ No No HARPS
HD 45067 $1.17$ $-0.049\pm 0.060$ No No CLS
HD 45184 $1.06$ $0.040\pm 0.060$ No No HARPS*+CLS
HD 45270 $1.08$ $-0.030\pm 0.040$ No No HARPS
HD 45289 $1.02$ $-0.050\pm 0.060$ No No HARPS*+AAT
HD 45346 $1.02$ $-0.057\pm 0.050$ No No HARPS
HD 45350 $1.05$ $0.292\pm 0.060$ Yes Yes CLS
HD 45364 $0.95$ $-0.200\pm 0.040$ No No HARPS
HD 457 $1.15$ $0.310\pm 0.060$ No No HARPS
HD 45701 $1.05$ $0.20\pm 0.10$ No No AAT
HD 4614 $0.91$ $-0.241\pm 0.060$ No No CLS
HD 4628 $0.74$ $-0.235\pm 0.060$ No No CLS
HD 46375 $0.93$ $0.271\pm 0.060$ No No CLS
HD 47157 $1.09$ $0.351\pm 0.060$ No No CLS
HD 47186 $1.00$ $0.216\pm 0.080$ Yes No HARPS
HD 4747 $0.83$ $-0.166\pm 0.060$ No No CLS
HD 47752 $0.73$ $-0.049\pm 0.080$ No No CLS
HD 48115 $1.07$ $-0.220\pm 0.020$ No No HARPS
HD 48265 $1.04$ $0.35\pm 0.20$ Yes Yes HARPS
HD 48682 $1.12$ $0.083\pm 0.060$ No No CLS
HD 48938 $0.86$ $-0.404\pm 0.060$ No No CLS
HD 49035 $1.00$ $0.240\pm 0.040$ No No HARPS
HD 4915 $1.01$ $-0.240\pm 0.100$ No No HARPS+CLS*
HD 49674 $1.06$ $0.300\pm 0.060$ No No CLS
HD 49933 A $1.38$ $-0.43\pm 0.10$ No No HARPS
HD 50281 $0.74$ $0.017\pm 0.080$ No No CLS
HD 50499 $1.25$ $0.305\pm 0.060$ Yes Yes CLS
HD 50554 $1.04$ $-0.021\pm 0.060$ Yes Yes CLS
HD 50590 $0.79$ $-0.250\pm 0.060$ No No HARPS
HD 50652 $1.02$ $0.290\pm 0.100$ No No HARPS
HD 50692 $0.91$ $-0.179\pm 0.060$ No No CLS
HD 50806 $1.04$ $0.063\pm 0.060$ No No CLS
HD 51419 $0.84$ $-0.308\pm 0.060$ No No CLS
HD 51754 $1.03$ $-0.580\pm 0.100$ No No HARPS
HD 51866 $0.81$ $0.088\pm 0.060$ No No CLS
HD 52265 $1.11$ $0.20\pm 0.20$ No No HARPS+CLS*
HD 52711 $0.93$ $-0.119\pm 0.060$ No No CLS
HD 52919 $0.74$ $0.073\pm 0.080$ No No CLS
HD 53705 $1.04$ $-0.24\pm 0.30$ No No AAT
HD 53706 $0.91$ $-0.29\pm 0.20$ No No AAT
HD 5372 $1.14$ $0.238\pm 0.060$ No No CLS
HD 5388 $1.21$ $-0.310\pm 0.060$ No No HARPS
HD 54351 $1.05$ $-0.110\pm 0.100$ No No HARPS
HD 55575 $0.87$ $-0.309\pm 0.060$ No No CLS
HD 55693 $1.07$ $0.26\pm 0.20$ No No HARPS*+AAT
HD 55720 $0.96$ $-0.350\pm 0.060$ No No AAT
HD 56380 $0.92$ $-0.440\pm 0.040$ No No HARPS
HD 564 $1.11$ $-0.230\pm 0.020$ Yes No HARPS
HD 58489 $0.77$ $0.120\pm 0.030$ No No HARPS
HD 59711 A $1.02$ $-0.13\pm 0.10$ No No HARPS
HD 60491 $0.79$ $-0.152\pm 0.050$ No No CLS
HD 61005 $0.97$ $0.030\pm 0.040$ No No HARPS
HD 6101 $0.79$ $-0.018\pm 0.050$ No No CLS
HD 61051 $0.94$ $-0.100\pm 0.080$ No No HARPS
HD 61383 $1.02$ $-0.520\pm 0.050$ No No HARPS
HD 61606 $0.81$ $0.033\pm 0.060$ No No CLS
HD 61902 $1.15$ $-0.650\pm 0.080$ No No HARPS
HD 61986 $1.03$ $-0.340\pm 0.030$ No No HARPS
HD 6204 $1.05$ $0.02\pm 0.10$ No No HARPS
HD 62128 $1.04$ $0.330\pm 0.030$ No No HARPS
HD 62364 $1.19$ $-0.140\pm 0.030$ No Yes HARPS
HD 62613 $0.89$ $-0.063\pm 0.060$ No No CLS
HD 63454 $0.78$ $0.130\pm 0.080$ No No HARPS
HD 6348 $0.85$ $-0.590\pm 0.060$ No No HARPS
HD 63487 $1.05$ $0.059\pm 0.040$ No No HARPS
HD 63608 $0.92$ $0.100\pm 0.020$ No No HARPS
HD 63754 $1.45$ $0.175\pm 0.060$ No No CLS
HD 63765 $0.95$ $-0.190\pm 0.040$ No No HARPS
HD 6434 $1.04$ $-0.54\pm 0.20$ No No HARPS
HD 65277 $0.74$ $-0.059\pm 0.070$ No No CLS
HD 65430 $0.82$ $-0.039\pm 0.050$ No No CLS
HD 65562 $0.84$ $-0.130\pm 0.030$ No No HARPS
HD 6558 $1.29$ $0.240\pm 0.060$ No No CLS
HD 65583 $0.73$ $-0.541\pm 0.050$ No No CLS
HD 66040 $0.89$ $0.320\pm 0.020$ No No HARPS
HD 66221 $0.99$ $0.152\pm 0.020$ No No HARPS
HD 66340 $0.91$ $0.030\pm 0.020$ No No HARPS
HD 66428 $1.01$ $0.250\pm 0.090$ Yes Yes HARPS+CLS*
HD 66653 $1.06$ $0.100\pm 0.100$ No No HARPS
HD 67 $1.03$ $0.00\pm 0.80$ No No HARPS
HD 6718 $1.02$ $-0.07\pm 0.10$ Yes Yes HARPS
HD 67199 $0.87$ $0.030\pm 0.020$ No No AAT
HD 67200 $1.18$ $0.290\pm 0.020$ No No HARPS
HD 6735 $1.14$ $-0.090\pm 0.040$ No No HARPS
HD 67556 $1.21$ $0.190\pm 0.040$ No No AAT
HD 6790 $1.17$ $0.010\pm 0.080$ No No HARPS
HD 68017 $0.82$ $-0.324\pm 0.060$ No No CLS
HD 68168 $1.02$ $0.12\pm 0.10$ No No HARPS
HD 68284 $1.08$ $-0.530\pm 0.100$ No No HARPS
HD 68607 $0.88$ $0.040\pm 0.040$ No No HARPS
HD 68978 $1.09$ $0.010\pm 0.040$ No No HARPS
HD 68988 $1.17$ $0.305\pm 0.060$ No No CLS
HD 69611 $1.04$ $-0.580\pm 0.100$ No No HARPS
HD 69655 $1.09$ $-0.210\pm 0.070$ No No HARPS*+AAT
HD 70642 $1.01$ $0.170\pm 0.080$ Yes Yes HARPS+AAT*
HD 70889 $1.11$ $0.086\pm 0.040$ No No HARPS+AAT*
HD 7134 $1.09$ $-0.320\pm 0.070$ No No HARPS
HD 71479 $1.08$ $0.21\pm 0.10$ No No HARPS
HD 71685 $1.12$ $-0.400\pm 0.020$ No No HARPS
HD 71835 $0.95$ $-0.040\pm 0.040$ No No HARPS
HD 7199 $0.93$ $0.270\pm 0.070$ No No HARPS
HD 72579 $0.95$ $0.170\pm 0.050$ No No HARPS
HD 72659 $1.09$ $-0.020\pm 0.080$ Yes Yes HARPS+CLS*
HD 73121 $1.12$ $0.060\pm 0.080$ No No HARPS*+AAT
HD 73256 $0.97$ $0.230\pm 0.060$ No Yes HARPS
HD 73267 $0.93$ $0.050\pm 0.030$ Yes Yes HARPS
HD 73344 $1.18$ $0.131\pm 0.060$ No No CLS
HD 73526 $0.98$ $0.25\pm 0.20$ Yes Yes AAT
HD 73667 $0.71$ $-0.409\pm 0.040$ No No CLS
HD 74000 $1.19$ $-2.01\pm 0.10$ No No HARPS
HD 74014 $0.98$ $0.200\pm 0.060$ No No HARPS
HD 74156 $1.15$ $0.079\pm 0.060$ Yes Yes CLS
HD 7449 $1.12$ $-0.140\pm 0.040$ Yes Yes HARPS
HD 74698 $1.05$ $0.060\pm 0.040$ Yes No HARPS
HD 74868 $1.19$ $0.240\pm 0.040$ No No AAT
HD 74957 $1.07$ $-0.210\pm 0.040$ No No HARPS
HD 75289 $1.12$ $0.28\pm 0.20$ No No HARPS+AAT*
HD 75302 $1.01$ $0.08\pm 0.10$ Yes Yes HARPS
HD 75530 $0.92$ $-0.570\pm 0.100$ No No HARPS
HD 7570 $1.15$ $0.15\pm 0.10$ No No AAT
HD 75732 $0.98$ $0.383\pm 0.060$ Yes Yes CLS
HD 75881 $1.17$ $0.040\pm 0.040$ No No HARPS
HD 76188 $1.11$ $-0.470\pm 0.020$ No No HARPS
HD 7661 $0.95$ $0.020\pm 0.030$ No No HARPS
HD 76700 $1.00$ $0.350\pm 0.060$ No No HARPS+AAT*
HD 76849 $0.90$ $-0.26\pm 0.20$ No No HARPS
HD 76909 $1.09$ $0.341\pm 0.060$ No No CLS
HD 77110 $1.02$ $-0.520\pm 0.020$ No No HARPS
HD 77338 $0.93$ $0.33\pm 0.10$ No No HARPS
HD 78286 $1.05$ $0.230\pm 0.080$ No No HARPS
HD 78429 $1.03$ $0.078\pm 0.070$ No No HARPS*+AAT
HD 78534 $1.04$ $0.038\pm 0.040$ No No HARPS
HD 78558 $1.06$ $-0.440\pm 0.090$ No No HARPS
HD 78660 $1.02$ $-0.012\pm 0.050$ No No HARPS
HD 78747 $1.03$ $-0.70\pm 0.20$ No No HARPS
HD 7924 $0.80$ $-0.123\pm 0.060$ No No CLS
HD 79601 $1.04$ $-0.620\pm 0.080$ No No HARPS
HD 80606 $1.05$ $0.348\pm 0.060$ No No CLS
HD 80913 $1.11$ $-0.6900\pm 0.0100$ No No AAT
HD 81639 $0.97$ $-0.200\pm 0.050$ No No HARPS
HD 81700 $1.08$ $0.120\pm 0.020$ No No HARPS
HD 81767 $0.82$ $0.070\pm 0.100$ No No HARPS
HD 82516 $0.85$ $0.020\pm 0.040$ No No HARPS
HD 82783 $0.92$ $0.180\pm 0.020$ No No HARPS
HD 82943 $1.08$ $0.25\pm 0.20$ Yes Yes HARPS*+CLS
HD 83443 $0.95$ $0.34\pm 0.20$ Yes Yes HARPS*+CLS
HD 83529 $1.07$ $-0.250\pm 0.060$ No No HARPS
HD 8389 $0.95$ $0.406\pm 0.050$ No No CLS
HD 84035 $0.77$ $0.269\pm 0.070$ No No CLS
HD 8406 $1.02$ $-0.10\pm 0.10$ No No HARPS
HD 84117 $1.16$ $-0.035\pm 0.060$ No No CLS*+AAT
HD 84737 $1.19$ $0.136\pm 0.060$ No No CLS
HD 85119 $0.94$ $-0.230\pm 0.070$ No No HARPS
HD 8535 $1.18$ $0.010\pm 0.020$ Yes Yes HARPS
HD 85390 $0.87$ $-0.090\pm 0.030$ No No HARPS
HD 85512 $0.69$ $-0.280\pm 0.080$ No No HARPS*+AAT
HD 85683 $1.17$ $0.040\pm 0.020$ No No AAT
HD 85725 $1.09$ $0.120\pm 0.080$ No No HARPS
HD 8574 $1.09$ $-0.010\pm 0.060$ No No CLS
HD 86065 $0.81$ $-0.030\pm 0.090$ No No HARPS
HD 86140 $0.79$ $-0.280\pm 0.050$ No No HARPS
HD 8648 $1.08$ $0.188\pm 0.060$ No No CLS
HD 86728 $1.07$ $0.226\pm 0.060$ No No CLS
HD 86819 $1.10$ $-0.080\pm 0.020$ No No AAT
HD 87359 $0.98$ $0.070\pm 0.060$ No No CLS
HD 8765 $1.03$ $0.223\pm 0.060$ No No CLS
HD 87836 $1.18$ $0.329\pm 0.060$ No No CLS
HD 87838 $1.15$ $-0.430\pm 0.060$ No No HARPS
HD 87883 $0.83$ $0.104\pm 0.060$ Yes Yes CLS
HD 88072 $1.04$ $0.02\pm 0.10$ No No HARPS
HD 88084 $1.03$ $-0.100\pm 0.050$ No No HARPS
HD 88230 $0.63$ $-0.040\pm 0.070$ No No CLS
HD 8828 $0.93$ $-0.160\pm 0.040$ No No HARPS
HD 8859 $0.97$ $-0.100\pm 0.040$ No No HARPS
HD 88742 $1.10$ $-0.050\pm 0.060$ No No AAT
HD 88986 $1.10$ $0.078\pm 0.060$ No No CLS
HD 89269 $0.87$ $-0.138\pm 0.060$ No No CLS
HD 89454 $1.02$ $0.110\pm 0.050$ No No HARPS
HD 89839 $1.24$ $0.010\pm 0.040$ Yes Yes HARPS
HD 89920 $0.78$ $-0.060\pm 0.020$ No No HARPS
HD 89965 $0.84$ $-0.110\pm 0.020$ No No HARPS
HD 90156 $1.00$ $-0.240\pm 0.050$ No No HARPS+CLS*
HD 90422 $1.13$ $-0.650\pm 0.060$ No No HARPS
HD 90520 $1.12$ $0.200\pm 0.100$ No No HARPS
HD 90711 $1.00$ $0.311\pm 0.060$ No No CLS
HD 90812 $0.87$ $-0.380\pm 0.040$ No No HARPS
HD 90875 $0.77$ $0.309\pm 0.070$ No No CLS
HD 90905 $1.14$ $0.070\pm 0.040$ No No HARPS
HD 91204 $1.16$ $0.211\pm 0.060$ No No CLS
HD 91379 $1.17$ $-0.320\pm 0.020$ No No HARPS
HD 92719 $1.04$ $-0.130\pm 0.080$ No No HARPS*+CLS
HD 92788 $1.08$ $0.277\pm 0.060$ Yes Yes CLS
HD 9280 $1.01$ $0.32\pm 0.10$ No No AAT
HD 92987 $1.04$ $-0.010\pm 0.070$ No No AAT
HD 93351 $0.94$ $-0.260\pm 0.020$ No No HARPS
HD 93372 $1.28$ $0.090\pm 0.040$ No No HARPS
HD 93385 $1.10$ $-0.010\pm 0.080$ No No HARPS*+AAT
HD 93745 $1.22$ $0.113\pm 0.060$ No No CLS
HD 9407 $0.93$ $0.029\pm 0.060$ No No CLS
HD 94151 $0.99$ $0.040\pm 0.060$ No No HARPS
HD 94771 $1.00$ $0.190\pm 0.020$ Yes Yes HARPS
HD 95128 $1.01$ $0.026\pm 0.060$ Yes Yes CLS
HD 95456 $1.23$ $0.130\pm 0.060$ No No HARPS
HD 95521 $1.03$ $-0.180\pm 0.060$ No No HARPS
HD 95542 $1.14$ $-0.060\pm 0.020$ No No HARPS
HD 9562 $1.33$ $0.222\pm 0.060$ No No CLS
HD 9578 $1.13$ $0.080\pm 0.040$ No No HARPS
HD 96116 $1.05$ $-0.010\pm 0.050$ No No HARPS
HD 96423 $1.02$ $0.100\pm 0.070$ No No HARPS*+AAT
HD 967 $0.99$ $-0.680\pm 0.040$ No No HARPS
HD 96700 $1.06$ $-0.210\pm 0.060$ No No HARPS
HD 97037 $1.06$ $-0.100\pm 0.080$ No No HARPS
HD 97101 $0.70$ $0.243\pm 0.070$ No No CLS
HD 97320 $1.12$ $-1.17\pm 0.20$ No No HARPS
HD 97343 $0.88$ $0.020\pm 0.060$ No No CLS
HD 97658 $0.77$ $-0.228\pm 0.060$ No No CLS
HD 9782 $1.11$ $0.060\pm 0.030$ No No HARPS
HD 9796 $0.87$ $-0.250\pm 0.040$ No No HARPS
HD 97998 $1.02$ $-0.450\pm 0.060$ No No HARPS
HD 9826 $1.29$ $0.122\pm 0.060$ Yes Yes CLS
HD 98284 $1.12$ $-0.870\pm 0.020$ No No HARPS
HD 984 $1.30$ $0.090\pm 0.080$ No No HARPS
HD 99109 $0.97$ $0.384\pm 0.060$ Yes No CLS
HD 99491 $1.02$ $0.331\pm 0.060$ No No CLS
HD 99492 $0.85$ $0.32\pm 0.10$ No No HARPS+CLS*
HD 9986 $1.04$ $0.083\pm 0.100$ No No HARPS+CLS*
HIP 100233 $1.10$ $-0.080\pm 0.080$ No No HARPS
HIP 100396 $1.18$ $0.100\pm 0.100$ No No HARPS
HIP 105388 $0.98$ $0.02\pm 0.20$ No No HARPS
HIP 106931 $0.98$ $-0.200\pm 0.060$ No No HARPS
HIP 108065 $0.97$ $0.050\pm 0.040$ No No HARPS
HIP 108095 $1.11$ $-0.410\pm 0.060$ No No HARPS
HIP 108950 $1.04$ $0.380\pm 0.020$ No No HARPS
HIP 109149 $0.82$ $-0.130\pm 0.020$ No No HARPS
HIP 110156 $0.77$ $0.050\pm 0.020$ No No HARPS
HIP 11048 $0.60$ $0.220\pm 0.080$ No No CLS
HIP 111978 $0.95$ $0.170\pm 0.040$ No No HARPS
HIP 112414 $1.04$ $-0.12\pm 0.20$ No No HARPS
HIP 115381 $0.87$ $-0.130\pm 0.040$ No No HARPS
HIP 115803 $1.20$ $-0.010\pm 0.050$ No No HARPS
HIP 18918 $0.72$ $-0.230\pm 0.020$ No No HARPS
HIP 19165 $0.66$ $-0.071\pm 0.070$ No No CLS
HIP 2247 $0.74$ $-0.020\pm 0.020$ Yes Yes HARPS
HIP 29788 $1.12$ $0.090\pm 0.080$ No No HARPS
HIP 30112 $0.70$ $0.139\pm 0.070$ No No CLS
HIP 32127 $0.91$ $-0.670\pm 0.020$ No No HARPS
HIP 32812 $0.82$ $-0.010\pm 0.020$ No No HARPS
HIP 39470 $0.72$ $-0.370\pm 0.050$ No No HARPS
HIP 40774 $0.79$ $-0.220\pm 0.060$ No No HARPS
HIP 41659 $0.88$ $-0.550\pm 0.030$ No No HARPS
HIP 42220 $0.64$ $0.136\pm 0.070$ No No CLS
HIP 45301 $0.77$ $-0.140\pm 0.020$ No No HARPS
HIP 49067 $0.78$ $-0.040\pm 0.020$ No No HARPS
HIP 5114 $0.79$ $-0.670\pm 0.080$ No No HARPS
HIP 52776 $0.73$ $-0.150\pm 0.020$ No No HARPS
HIP 54597 $0.78$ $-0.250\pm 0.020$ Yes Yes HARPS
HIP 58289 $0.78$ $-0.130\pm 0.020$ No No HARPS
HIP 60051 $0.98$ $-0.510\pm 0.020$ No No HARPS
HIP 60633 $0.76$ $0.288\pm 0.070$ No No CLS
HIP 6276 $0.92$ $-0.020\pm 0.040$ No No HARPS
HIP 63719 $1.42$ $-0.240\pm 0.020$ No No HARPS
HIP 69570 $0.99$ $-0.300\pm 0.040$ No No HARPS
HIP 70681 $0.98$ $-1.17\pm 0.20$ No No HARPS
HIP 72779 $1.01$ $0.060\pm 0.040$ No No HARPS
HIP 78408 $0.98$ $0.330\pm 0.080$ No No HARPS
HIP 79578 $1.04$ $0.048\pm 0.060$ No No HARPS
HIP 80683 $0.80$ $-0.060\pm 0.040$ No No HARPS
HIP 8361 $0.77$ $0.040\pm 0.040$ No No HARPS
HIP 90194 $0.92$ $-0.320\pm 0.020$ No No HARPS
HIP 93745 $1.16$ $-0.400\pm 0.020$ No No HARPS
HIP 98106 $0.78$ $-0.010\pm 0.070$ No No HARPS
HIP 99695 $0.85$ $-0.090\pm 0.020$ No No HARPS
TYC6296-1176-1 $1.11$ $0.000\pm 0.040$ No No HARPS
TYC6296-478-1 $1.02$ $0.060\pm 0.040$ No No HARPS
TYC9111-1423-1 $1.15$ $-0.68\pm 0.20$ No No HARPS
 [^15] From left to right, the columns report the name of the star, its mass and metallicity as taken from the Hypatia catalog (Hinkel et al. 2014), the presence or absence of confirmed CJs in the $0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ and $0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ mass ranges (“Yes” appears in both columns if at least one CJ in the overlapping range $0.5-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ is present), and the name of the radial velocity survey (AAT and/or CLS and/or HARPS). In case the star was observed by more than one survey, the asterisk indicates the survey with the best RV dataset in terms of baseline and/or RV precision that was used in this work.
Appendix C Comparison of the stellar properties of the small planet and merged RV samples
Figure 4: Stellar metallicity distributions in the small planet sample (red solid line) and merged (AAT+CLS+HARPS) radial velocity comparison sample (black dashed line). Note that the relative stellar counts of the SP and RV merged samples cannot be directly compared as they were normalized from different distributions.
Figure 5: Stellar mass distributions in the small planet sample (red solid line) and merged (AAT+CLS+HARPS) radial velocity comparison sample (black dashed line). Note that the relative stellar counts of the SP and RV merged samples cannot be directly compared as they were normalized from different distributions.
Figure 6: Upper panels: same as Fig. 4 for the three different bins of stellar mass $0.6-0.8\,\hbox{$\mathrm{M}_{\odot}$}$ (left), $0.8-1.0\,\hbox{$\mathrm{M}_{\odot}$}$ (center), and $1.0-1.2\,\hbox{$\mathrm{M}_{\odot}$}$ (right). Note that the relative stellar counts of the SP and RV merged samples cannot be directly compared as they were normalized from different distributions. Lower panels: empirical cumulative distribution functions for the same bins of stellar mass as above. The p-values of the Kolmogorov-Smirnov test are also reported.
Appendix D Completenesses of the considered samples
(a) SP RV sample
(b) SP transit sample
(c) SP (transit+RV) sample
(d) SP and mixed RV sample
(e) SP and mixed transit sample
(f) SP and mixed (transit+RV) sample
(g) AAT RV survey
(h) CLS RV survey
(i) HARPS RV survey
(j) Merged (AAT+CLS+HARPS) RV sample
Figure 7: Average completenesses of all the samples considered in this work. As indicated in the color bars, the detection rate/sensitivity increases from dark blue ($0\%$) to green ($\sim 50-80\%$) and yellow ($100\%$).
Appendix E Cold Jupiters and their occurrence rates in the small planet and radial velocity comparison samples
Table 4: Cold Jupiters in the small planet sample.
Target $M_{\rm p}\sin(i)$ or $M_{\rm p}$ $a$ M [Fe/H] Ref.
[$\mathrm{M}_{\rm Jup}$] [AU] [M] [dex]
BD-11 4672 b $0.65_{-0.06}^{+0.05}$ $2.360\pm 0.040$ $0.650\pm 0.030$ $-0.35\pm 0.20$ Barbato et al. (2020)
GJ 676A b $5.790_{-0.477}^{+0.500}$ $1.735_{-0.060}^{+0.056}$ $0.73\pm 0.20$ $0.080$ Feng et al. (2022)
GJ 676A c $13.500_{-1.127}^{+1.000}$ $9.726_{-0.793}^{+0.629}$ $0.73\pm 0.20$ $0.080$ Feng et al. (2022)
HD 125612 b $2.974_{-0.258}^{+0.245}$ $1.349_{-0.059}^{+0.054}$ $1.11\pm 0.20$ $0.250\pm 0.040$ Feng et al. (2022)
HD 125612 d $7.180_{-0.445}^{+0.900}$ $3.982_{-0.174}^{+0.159}$ $1.11\pm 0.20$ $0.250\pm 0.040$ Feng et al. (2022)
HD 156098 c $5.228_{-1.287}^{+1.939}$ $8.41_{-0.629}^{+0.396}$ $1.29\pm 0.20$ $0.190\pm 0.100$ Feng et al. (2022)
HD 160691 b $1.67\pm 0.11$ $1.50\pm 0.02$ $1.130\pm 0.020$ $0.280\pm 0.030$ McCarthy et al. (2004)
HD 160691 c $3.10\pm 0.71$ $4.17\pm 0.07$ $1.130\pm 0.020$ $0.280\pm 0.030$ McCarthy et al. (2004)
HD 164922 b $0.344\pm 0.013$ $2.149\pm 0.025$ $0.920\pm 0.030$ $0.200\pm 0.060$ Rosenthal et al. (2021)
HD 181433 c $0.674\pm 0.003$ $1.819\pm 0.001$ $0.780$ $0.33\pm 0.10$ Horner et al. (2019)
HD 181433 d $0.612\pm 0.004$ $6.60\pm 0.22$ $0.780$ $0.33\pm 0.10$ Horner et al. (2019)
HD 190360 b $1.492\pm 0.043$ $3.955_{-0.053}^{+0.051}$ $0.990\pm 0.040$ $0.230\pm 0.060$ Rosenthal et al. (2021)
HD 204313 b $4.96_{-0.45}^{+1.70}$ $3.208_{-0.035}^{+0.034}$ $1.045\pm 0.030$ $0.180\pm 0.020$ Xiao et al. (2023)
HD 204313 e $15.300_{-5.183}^{+4.900}$ $7.457_{-0.427}^{+0.399}$ $1.045\pm 0.030$ $0.180\pm 0.020$ Feng et al. (2022)
HD 215497 c $0.35\pm 0.02$ $1.310\pm 0.020$ $0.920\pm 0.050$ $0.230\pm 0.070$ Philipot et al. (2023)
HD 219134 h $0.308\pm 0.014$ $2.968\pm 0.037$ $0.790\pm 0.030$ $0.080\pm 0.060$ Rosenthal et al. (2021)
HD 219828 c $16.10_{-1.27}^{+3.50}$ $5.657_{-0.242}^{+0.228}$ $1.18\pm 0.20$ $0.180$ Feng et al. (2022)
HD 34445 b $0.658_{-0.040}^{+0.041}$ $2.105_{-0.039}^{+0.038}$ $1.110\pm 0.060$ $0.140\pm 0.060$ Rosenthal et al. (2021)
HD 34445 g $0.38\pm 0.13$ $6.36\pm 1.02$ $1.110\pm 0.060$ $0.140\pm 0.060$ Vogt et al. (2017)
HD 39091 b/$\pi$ Men b $12.300_{-1.384}^{+1.200}$ $3.311_{-0.148}^{+0.134}$ $1.070\pm 0.040$ $0.090\pm 0.040$ Feng et al. (2022)
HD 74698 c $0.40\pm 0.06$ $4.5\pm 0.2$ $1.040\pm 0.025$ $0.07\pm 0.02$ Frensch et al. (2023)
HD 86226 b $0.45^{+0.04}_{-0.05}$ $2.730\pm 0.060$ $1.020\pm 0.060$ $0.020\pm 0.050$ Teske et al. (2020)
Kepler-48 e $2.163\pm 0.062$ $1.898$ $0.880\pm 0.060$ $0.170\pm 0.070$ Weiss et al. (2024)
Kepler-48 f $0.939\pm 0.292$ $5.719$ $0.880\pm 0.060$ $0.170\pm 0.070$ Weiss et al. (2024)
Kepler-68 d $0.749\pm 0.017$ $1.469\pm 0.010$ $1.080\pm 0.050$ $0.120\pm 0.070$ Bonomo et al. (2023)
Kepler-94 c $8.885\pm 0.117$ $1.597$ $0.810\pm 0.060$ $0.340\pm 0.070$ Weiss et al. (2024)
Kepler-129 d $8.30_{-0.70}^{+1.10}$ $4.00\pm 0.10$ $1.178\pm 0.030$ $0.290\pm 0.100$ Zhang et al. (2021)
Kepler-139 e $1.341\pm 0.134$ $3.236$ $1.078\pm 0.030$ $0.350\pm 0.040$ Weiss et al. (2024)
Kepler-407 c $11.089\pm 0.049$ $3.27$ $1.000\pm 0.060$ $0.330\pm 0.070$ Weiss et al. (2024)
Kepler-454 c $4.51\pm 0.12$ $1.287\pm 0.017$ $1.030\pm 0.040$ $0.320\pm 0.080$ Bonomo et al. (2023)
Kepler-454 d $2.31_{-0.16}^{+0.27}$ $5.10_{-0.19}^{+0.34}$ $1.030\pm 0.040$ $0.320\pm 0.080$ Bonomo et al. (2023)
K2-312 c/HD 80653 c $5.41_{-0.44}^{+0.52}$ $1.961\pm 0.027$ $1.180\pm 0.040$ $0.255\pm 0.070$ Bonomo et al. (2023)
K2-364 c/HD 137496 c $7.66\pm 0.11$ $1.2163\pm 0.0087$ $1.040\pm 0.020$ $-0.030\pm 0.040$ Azevedo Silva et al. (2022)
TOI-1669 c $0.601_{-0.076}^{+0.079}$ $1.27\pm 0.03$ $1.00\pm 0.05$ $0.26\pm 0.06$ Polanski et al. (2024)
TOI-1736 c $8.70_{-0.60}^{+1.50}$ $1.381\pm 0.017$ $1.080\pm 0.040$ $0.138\pm 0.010$ Martioli et al. (2023)
TOI-1339 f/HD 191939 f $6.50\pm 4.515$ $4.80\pm 2.20$ $0.810\pm 0.040$ $-0.150\pm 0.060$ Lubin et al. (2022)
 [^16] From left to right, the columns report the name of the cold Jupiter, its mass and semi-major axis, the mass and metallicity of the host star, and the reference for the cold-Jupiter mass determination. Asterisks denote true masses instead of minimum masses thanks to the combination of radial velocities and measurements of Hipparcos-Gaia proper motion anomalies (e.g., Kervella et al. 2022).
Table 5: Cold Jupiters in the mixed system sample.
Target $M_{\rm p}\sin(i)$ or $M_{\rm p}$ $a$ M [Fe/H] Ref.
[$\mathrm{M}_{\rm Jup}$] [AU] [M] [dex]
55 Cnc d $3.12\pm 0.10$ $5.96\pm 0.07$ $0.975\pm 0.045$ $0.383\pm 0.055$ Bourrier et al. (2018)
HIP 57274 d $0.53\pm 0.03$ $1.01$ $0.73\pm 0.05$ $0.09\pm 0.05$ Fischer et al. (2012)
Kepler-56 d $5.61\pm 0.38$ $2.16\pm 0.08$ $1.67\pm 0.15$ $0.40\pm 0.04$ Otor et al. (2016)
Kepler-88 d/KOI-142 d $3.15\pm 0.17$ $2.47$ $0.985\pm 0.021$ $0.27\pm 0.04$ Weiss et al. (2024)
TOI-4010 e $2.18_{-0.20}^{+0.21}$ $1.57_{-0.13}^{+0.12}$ $0.89\pm 0.03$ $0.37\pm 0.07$ Kunimoto et al. (2023)
WASP-47 c $1.41_{-0.07}^{+0.06}$ $1.393\pm 0.014$ $1.04\pm 0.03$ $0.38\pm 0.05$ Nascimbeni et al. (2023)
 [^17] Same as Table LABEL:table_CJs_SPsample but for cold Jupiters within mixed systems.
Figure 8: Similarly to Fig. 1, occurrence rates of cold Jupiters with $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ in the AAT (orange), CLS (violet), and HARPS (green) radial velocity surveys, at sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity. Almost identical occurrence rates are obtained for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$. See also Table LABEL:table_RVsurveys.
Figure 9: Same as Fig. 1 for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$. See also Table 19.
Figure 10: Same as Fig. 2 for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$. See also Table 20.
Figure 11: The architectures of the systems with both small planets and cold Jupiters considered in this work and ordered according to increasing semi-major axis of the innermost planet (from bottom to top). The cold Jupiters in the left (27 systems) and right (23 systems) panels have masses $0.3\leq{\rm M_{p}}\leq 13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ and $0.5\leq{\rm M_{p}}\leq 20\,\hbox{$\mathrm{M}_{\rm Jup}$}$, respectively, with a large overlap for cold Jupiters with $0.5\leq{\rm M_{p}}\leq 13\,\hbox{$\mathrm{M}_{\rm Jup}$}$. The planet circles have sizes proportional to their mass (see legend in the top right corner), and colors depending on the orbital eccentricity (more eccentric planets are shown in yellow). The host stars are color coded as a function of stellar metallicity (metallicity increases from blue to red).
Table 6: Cold Jupiters in the AAT, CLS, and HARPS radial velocity surveys.
Target CJ name $M_{\rm p}\sin(i)$ or $M_{\rm p}$ $a$ M [Fe/H] Ref.
[$\mathrm{M}_{\rm Jup}$] [AU] [M] [dex]
HD 19994 HD 19994 b $1.37\pm 0.12$ $1.305\pm 0.016$ $1.14$ $0.21\pm 0.20$ Wittenmyer et al. (2009)
HD 60532 HD 60532 c $2.51\pm 0.16$ $1.600\pm 0.040$ $1.20$ $-0.17\pm 0.20$ Borgniet et al. (2017)
HD 160691 HD 160691 b $1.67\pm 0.11$ $1.50\pm 0.02$ $1.130\pm 0.020$ $0.280\pm 0.030$ McCarthy et al. (2004)
HD 160691 c $3.10\pm 0.71$ $4.17\pm 0.07$ $1.130\pm 0.020$ $0.280\pm 0.030$ McCarthy et al. (2004)
HD 100777 HD 100777 b $1.230_{-0.050}^{+0.100}$ $1.060_{-0.010}^{+0.020}$ $0.97$ $0.250\pm 0.060$ Philipot et al. (2023)
HD 103891 HD 103891 b $2.89_{-0.84}^{+0.90}$ $3.255_{-0.026}^{+0.024}$ $1.10$ $-0.220\pm 0.080$ Xiao et al. (2023)
HD 105779 HD 105779 b $0.640\pm 0.060$ $3.380\pm 0.050$ $1.04$ $-0.250\pm 0.020$ Sreenivas et al. (2022)
HD 10647 HD 10647 b $0.940\pm 0.080$ $2.015\pm 0.011$ $1.17$ $-0.03\pm 0.10$ Marmier et al. (2013)
HD 10697 HD 10697 b $6.77_{-0.56}^{+0.90}$ $2.143_{-0.063}^{+0.062}$ $1.16$ $0.160\pm 0.060$ Xiao et al. (2023)
HD 108341 HD 108341 b $2.25_{-0.25}^{+0.70}$ $2.029_{-0.022}^{+0.018}$ $0.89$ $0.020\pm 0.020$ Xiao et al. (2023)
HD 108874 HD 108874 b $1.320\pm 0.047$ $1.052\pm 0.016$ $1.00$ $0.221\pm 0.060$ Rosenthal et al. (2021)
HD 111232 HD 111232 b $7.47_{-0.26}^{+0.60}$ $2.051_{-0.024}^{+0.023}$ $0.96$ $-0.450\pm 0.090$ Xiao et al. (2023)
HD 114729 HD 114729 b $0.892\pm 0.053$ $2.094\pm 0.022$ $1.05$ $-0.28\pm 0.20$ Rosenthal et al. (2021)
HD 114783 HD 114783 b $1.033\pm 0.034$ $1.164\pm 0.016$ $0.87$ $0.151\pm 0.060$ Rosenthal et al. (2021)
HD 114783 c $1.90_{-0.40}^{+0.50}$ $5.000\pm 0.100$ $0.87$ $0.151\pm 0.060$ Philipot et al. (2023)
HD 117207 HD 117207 b $2.110_{-0.089}^{+0.200}$ $3.773_{-0.035}^{+0.036}$ $0.99$ $0.220\pm 0.080$ Xiao et al. (2023)
HD 11964 HD 11964 b $0.631\pm 0.027$ $3.185_{-0.032}^{+0.031}$ $0.91$ $0.080\pm 0.100$ Rosenthal et al. (2021)
HD 12661 HD 12661 c $1.855\pm 0.054$ $2.861_{-0.040}^{+0.038}$ $1.09$ $0.338\pm 0.060$ Rosenthal et al. (2021)
HD 126614 HD 126614 b $0.339_{-0.019}^{+0.020}$ $2.249_{-0.096}^{+0.088}$ $1.02$ $0.408\pm 0.050$ Feng et al. (2022)
HD 128311 HD 128311 b $2.00\pm 0.16$ $1.088_{-0.013}^{+0.012}$ $0.84$ $0.134\pm 0.060$ Rosenthal et al. (2021)
HD 128311 c $3.25\pm 0.12$ $1.742\pm 0.020$ $0.84$ $0.134\pm 0.060$ Rosenthal et al. (2021)
HD 134987 HD 134987 c $0.935_{-0.060}^{+0.063}$ $6.62_{-0.15}^{+0.16}$ $1.02$ $0.26\pm 0.20$ Rosenthal et al. (2021)
HD 135625 HD 135625 b $2.30_{-0.30}^{+0.80}$ 5.400±+0.1005.400$\pm$+$0$ + 0.100 $1.12$ $0.090\pm 0.020$ Philipot et al. (2023)
HD 136925 HD 136925 b $0.840_{-0.074}^{+0.078}$ $5.13_{-0.11}^{+0.12}$ $0.88$ $-0.217\pm 0.060$ Rosenthal et al. (2021)
HD 13931 HD 13931 b $2.80_{-0.64}^{+0.80}$ $5.338_{-0.08}^{+0.082}$ $1.02$ $0.041\pm 0.060$ Xiao et al. (2023)
HD 141399 HD 141399 d $1.263\pm 0.047$ $2.114_{-0.036}^{+0.036}$ $1.09$ $0.366\pm 0.060$ Rosenthal et al. (2021)
HD 141399 e $0.644_{-0.040}^{+0.042}$ $4.50\pm 0.11$ $1.09$ $0.366\pm 0.060$ Rosenthal et al. (2021)
HD 142A HD 142A b $1.249_{-0.130}^{+0.120}$ $1.039_{-0.051}^{+0.046}$ $1.21$ $0.07\pm 0.20$ Feng et al. (2022)
HD 142A c $10.901_{-0.937}^{+1.278}$ $9.811_{-0.520}^{+0.515}$ $1.21$ $0.07\pm 0.20$ Feng et al. (2022)
HD 143361 HD 143361 b $4.35_{-0.66}^{+1.20}$ $1.994\pm 0.018$ $0.97$ $0.250\pm 0.080$ Xiao et al. (2023)
HD 145675 HD 145675 b $8.053_{-0.883}^{+1.632}$ $2.771_{-0.120}^{+0.109}$ $0.97$ $0.405\pm 0.050$ Feng et al. (2022)
HD 147873 HD 147873 c $2.30\pm 0.18$ $1.360\pm 0.050$ $1.18$ $0.08\pm 0.30$ Jenkins et al. (2017)
HD 148156 HD 148156 b $0.85_{-0.05}^{+0.67}$ $2.450_{-0.050}^{+0.040}$ $1.18$ $0.220\pm 0.030$ Naef et al. (2010)
HD 153950 HD 153950 b $2.730\pm 0.050$ $1.280\pm 0.010$ $1.11$ $-0.040\pm 0.030$ Moutou et al. (2009)
HD 154345 HD 154345 b $0.905_{-0.089}^{+0.071}$ $4.272_{-0.079}^{+0.086}$ $0.89$ $-0.070\pm 0.060$ Rosenthal et al. (2021)
HD 154857 HD 154857 b $1.80\pm 0.40$ $1.20\pm 0.20$ $0.99$ $-0.220\pm 0.040$ Stassun et al. (2017)
HD 154857 c $2.58\pm 0.16$ $5.360\pm 0.090$ $0.99$ $-0.220\pm 0.040$ Wittenmyer et al. (2014)
HD 156098 HD 156098 c $5.228_{-1.287}^{+1.938}$ $8.415_{-0.629}^{+0.396}$ $1.29$ $0.150\pm 0.100$ Feng et al. (2022)
HD 156279 HD 156279 c $9.750_{-0.605}^{+1.300}$ $5.486_{-0.240}^{+0.219}$ $0.96$ $0.192\pm 0.060$ Feng et al. (2022)
HD 156411 HD 156411 b $0.740_{-0.040}^{+0.050}$ $1.880_{-0.040}^{+0.030}$ $1.07$ $-0.110\pm 0.100$ Naef et al. (2010)
HD 159868 HD 159868 b $2.218\pm 0.059$ $2.32\pm 0.14$ $0.97$ $-0.07\pm 0.10$ Luhn et al. (2019)
HD 159868 c $0.768\pm 0.044$ $1.032\pm 0.064$ $0.97$ $-0.07\pm 0.10$ Luhn et al. (2019)
HD 164922 HD 164922b $0.344\pm 0.013$ $2.149\pm 0.025$ $0.92$ $0.197\pm 0.060$ Rosenthal et al. (2021)
HD 165131 HD 165131 b $18.70_{-1.00}^{+1.40}$ $3.540\pm 0.054$ $1.06$ $0.030\pm 0.020$ Xiao et al. (2023)
HD 165155 HD 165155 b $2.89\pm 0.23$ $1.130\pm 0.040$ $0.94$ $0.210\pm 0.080$ Jenkins et al. (2017)
HD 166724 HD 166724 b $3.80_{-0.29}^{+0.70}$ $5.17_{-0.49}^{+0.38}$ $0.86$ $-0.090\pm 0.050$ Xiao et al. (2023)
HD 167677 HD 167677 b $2.85\pm 1.00$ $2.877\pm 0.025$ $0.96$ $-0.320\pm 0.100$ Xiao et al. (2023)
HD 168443 HD 168443 c $17.300_{-0.906}^{+2.600}$ $2.842_{-0.120}^{+0.110}$ $1.04$ $0.076\pm 0.060$ Feng et al. (2022)
HD 16905 HD 16905 b $11.30_{-0.70}^{+0.60}$ $8.80_{-0.30}^{+0.40}$ $0.79$ $0.170\pm 0.080$ Philipot et al. (2023)
HD 169830 HD 169830 c $8.90_{-2.60}^{+2.80}$ $3.290_{-0.050}^{+0.040}$ $1.17$ $0.17\pm 0.20$ Philipot et al. (2023)
HD 170469 HD 170469 b $0.555_{-0.072}^{+0.075}$ $2.213_{-0.044}^{+0.040}$ $1.14$ $0.311\pm 0.060$ Rosenthal et al. (2021)
HD 171028 HD 171028 b $2.62\pm 0.16$ $1.320\pm 0.030$ $1.01$ $-0.480\pm 0.100$ Stassun et al. (2017)
HD 181433 HD 181433 c $0.674\pm 0.003$ $1.819\pm 0.001$ $0.81$ $0.38\pm 0.10$ Horner et al. (2019)
HD 181433 d $0.612\pm 0.004$ $6.60\pm 0.22$ $0.81$ $0.38\pm 0.10$ Horner et al. (2019)
HD 181720 HD 181720 b $0.40\pm 0.06$ $1.85\pm 0.040$ $1.04$ $-0.560\pm 0.030$ Stassun et al. (2017)
HD 183263 HD 183263 b $3.471_{-0.304}^{+0.289}$ $1.460_{-0.065}^{+0.060}$ $1.17$ $0.278\pm 0.060$ Feng et al. (2022)
HD 183263 c $9.310_{-1.822}^{+1.500}$ $5.585_{-0.251}^{+0.228}$ $1.17$ $0.278\pm 0.060$ Feng et al. (2022)
HD 185283 HD 185283 b $1.30_{-0.10}^{+0.30}$ $4.60_{-0.10}^{+0.20}$ $0.79$ $-0.070\pm 0.060$ Philipot et al. (2023)
HD 186427 16 Cyg B b $1.752_{-0.053}^{+0.054}$ $1.676\pm 0.025$ $0.98$ $0.070\pm 0.060$ Rosenthal et al. (2021)
HD 187085 HD 187085 b $0.836\pm 0.011$ $2.100\pm 0.032$ $1.15$ $0.090\pm 0.100$ Barbato et al. (2018)
HD 187123 HD 187123 c $1.713_{-0.057}^{+0.058}$ $4.432_{-0.072}^{+0.070}$ $1.02$ $0.106\pm 0.060$ Rosenthal et al. (2021)
HD 188015 HD 188015 b $1.454_{-0.060}^{+0.061}$ $1.187_{-0.019}^{+0.018}$ $1.04$ $0.253\pm 0.060$ Rosenthal et al. (2021)
HD 190360 HD 190360 b $1.492\pm 0.043$ $3.955_{-0.053}^{+0.051}$ $0.99$ $0.230\pm 0.060$ Rosenthal et al. (2021)
HD 190647 HD 190647 b $1.985\pm 0.033$ $2.231\pm 0.019$ $1.00$ $0.230\pm 0.100$ Barbato et al. (2018)
HD 190984 HD 190984 b $3.58_{-0.45}^{+1.20}$ $8.80_{-1.40}^{+2.50}$ $1.10$ $-0.520\pm 0.020$ Xiao et al. (2023)
HD 196050 HD 196050 b $4.55_{-0.72}^{+0.70}$ $2.585_{-0.035}^{+0.032}$ $1.07$ $0.22\pm 0.20$ Xiao et al. (2023)
HD 202206 HD 202206 c $17.90_{-1.80}^{+2.90}$ $2.410$ $1.03$ $0.29\pm 0.16$ Benedict & Harrison (2017)
HD 2039 HD 2039 b $6.29\pm 1.16$ $2.200\pm 0.060$ $1.09$ $0.31\pm 0.20$ Stassun et al. (2017)
HD 204313 HD 204313 b $4.96_{-0.45}^{+1.70}$ $3.208_{-0.035}^{+0.034}$ $1.03$ $0.160\pm 0.040$ Xiao et al. (2023)
HD 204313 e $15.300_{-5.183}^{+4.900}$ $7.457_{-0.427}^{+0.399}$ $1.03$ $0.160\pm 0.040$ Feng et al. (2022)
HD 20782 HD 20782 b $0.730\pm 0.050$ $1.359\pm 0.005$ $1.04$ $-0.060\pm 0.080$ Stassun et al. (2017)
HD 207832 HD 207832 c $0.73_{-0.05}^{+0.18}$ $2.112_{-0.045}^{+0.087}$ $1.02$ $0.130\pm 0.040$ Haghighipour et al. (2012)
HD 210277 HD 210277 b $1.236\pm 0.032$ $1.123\pm 0.014$ $0.96$ $0.20\pm 0.10$ Rosenthal et al. (2021)
HD 213240 HD 213240 b $5.21_{-0.49}^{+1.50}$ $1.920\pm 0.026$ $1.10$ $0.14\pm 0.20$ Xiao et al. (2023)
HD 215497 HD 215497 c $0.35\pm 0.02$ $1.310\pm 0.020$ $0.86$ $0.170\pm 0.020$ Philipot et al. (2023)
HD 216435 HD 216435 b $1.26\pm 0.13$ $2.56\pm 0.17$ $1.10$ $0.21\pm 0.20$ Butler et al. (2006)
HD 216437 HD 216437 b $3.88_{-0.73}^{+0.70}$ $2.501_{-0.037}^{+0.036}$ $1.06$ $0.22\pm 0.20$ Xiao et al. (2023)
HD 217107 HD 217107 c $4.31\pm 0.13$ $5.942_{-0.086}^{+0.083}$ $1.06$ $0.340\pm 0.060$ Rosenthal et al. (2021)
HD 217786 HD 217786 b $13.900_{-1.311}^{+1.300}$ $2.446_{-0.119}^{+0.109}$ $1.14$ $-0.170\pm 0.030$ Feng et al. (2022)
HD 219077 HD 219077 b $10.60_{-0.17}^{+0.20}$ $6.13_{-0.10}^{+0.11}$ $0.92$ $-0.16\pm 0.10$ Xiao et al. (2023)
HD 219134 HD 219134 h $0.308\pm 0.014$ $2.968\pm 0.037$ $0.79$ $0.083\pm 0.060$ Rosenthal et al. (2021)
HD 219828 HD 219828 c $16.10_{-1.27}^{+3.50}$ $5.657_{-0.242}^{+0.228}$ $1.04$ $0.15\pm 0.10$ Feng et al. (2022)
HD 22049 eps Eri b $0.651_{-0.039}^{+0.038}$ $3.500\pm 0.039$ $0.85$ $-0.09\pm 0.50$ Rosenthal et al. (2021)
HD 221287 HD 221287 b $3.09\pm 0.79$ $1.250\pm 0.040$ $1.27$ $0.01\pm 0.10$ Naef et al. (2007)
HD 222582 HD 222582 b $7.88\pm 0.24$ $1.335\pm 0.020$ $1.04$ $-0.01\pm 0.20$ Rosenthal et al. (2021)
HD 224538 HD 224538 b $6.530_{-0.349}^{+1.800}$ $2.375_{-0.123}^{+0.111}$ $1.23$ $0.28\pm 0.10$ Feng et al. (2022)
HD 23079 HD 23079 b $2.61\pm 0.11$ $1.600\pm 0.030$ $1.10$ $-0.14\pm 0.10$ Stassun et al. (2017)
HD 23127 HD 23127 b $1.527_{-0.038}^{+0.037}$ $2.370\pm 0.032$ $1.05$ $0.34\pm 0.20$ Barbato et al. (2018)
HD 24040 HD 24040 b $4.05\pm 0.15$ $4.708_{-0.077}^{+0.074}$ $1.10$ $0.197\pm 0.060$ Rosenthal et al. (2021)
HD 25171 HD 25171 b $0.915_{-0.012}^{+0.011}$ $2.971_{-0.032}^{+0.033}$ $1.20$ $-0.140\pm 0.040$ Barbato et al. (2018)
HD 27442 A HD 27442 A b $1.56\pm 0.14$ $1.271\pm 0.073$ $1.21$ $0.39\pm 0.20$ Butler et al. (2006)
HD 27894 HD 27894 d $6.490_{-0.353}^{+1.000}$ $5.36_{-0.223}^{+0.206}$ $0.81$ $0.26\pm 0.20$ Feng et al. (2022)
HD 28185 HD 28185 b $5.837_{-0.510}^{+0.486}$ $1.035_{-0.046}^{+0.042}$ $1.02$ $0.227\pm 0.060$ Feng et al. (2022)
HD 28254 HD 28254 b $3.80_{-2.20}^{+3.00}$ $2.450_{-0.040}^{+0.030}$ $1.00$ $0.330\pm 0.020$ Philipot et al. (2023)
HD 290327 HD 290327 b $2.43\pm 0.42$ $3.43\pm 0.16$ $0.97$ $-0.170\pm 0.020$ Stassun et al. (2017)
HD 30177 HD 30177 b $8.400_{-0.489}^{+1.200}$ $3.60_{-0.147}^{+0.135}$ $0.98$ $0.37\pm 0.20$ Feng et al. (2022)
HD 30177 c $7.60\pm 3.10$ $9.89\pm 1.04$ $0.98$ $0.37\pm 0.20$ Wittenmyer et al. (2017)
HD 30669 HD 30669 b $0.470\pm 0.060$ $2.690\pm 0.080$ $0.94$ $0.130\pm 0.040$ Moutou et al. (2015)
HD 31253 HD 31253 b $0.446\pm 0.063$ $1.295_{-0.025}^{+0.024}$ $1.33$ $0.140\pm 0.060$ Rosenthal et al. (2021)
HD 32963 HD 32963 b $2.07_{-0.64}^{+0.80}$ $3.409_{-0.064}^{+0.063}$ $0.98$ $0.075\pm 0.060$ Xiao et al. (2023)
HD 34445 HD 34445 b $0.658_{-0.040}^{+0.041}$ $2.105_{-0.039}^{+0.038}$ $1.11$ $0.138\pm 0.060$ Rosenthal et al. (2021)
HD 34445 g $0.38\pm 0.13$ $6.36\pm 1.02$ $1.11$ $0.138\pm 0.060$ Vogt et al. (2017)
HD 37124 HD 37124 c $0.647_{-0.090}^{+0.080}$ $1.687\pm 0.011$ $0.81$ $-0.326\pm 0.060$ Rosenthal et al. (2021)
HD 38283 HD 38283 b $0.400\pm 0.040$ $1.020\pm 0.020$ $1.10$ $-0.250\pm 0.100$ Stassun et al. (2017)
HD 39091 $\pi$ Men b $12.300_{-1.384}^{+1.200}$ $3.311_{-0.148}^{+0.134}$ $1.10$ $0.07\pm 0.10$ Feng et al. (2022)
HD 4203 HD 4203 b $1.822\pm 0.078$ $1.177_{-0.022}^{+0.021}$ $1.14$ $0.380\pm 0.060$ Rosenthal et al. (2021)
HD 4203 c $2.68_{-0.24}^{+0.99}$ $7.80_{-0.78}^{+5.40}$ $1.14$ $0.380\pm 0.060$ Rosenthal et al. (2021)
HD 4208 HD 4208 b $0.771\pm 0.036$ $1.634\pm 0.022$ $0.85$ $-0.220\pm 0.060$ Rosenthal et al. (2021)
HD 43197 HD 43197 c $7.870_{-1.599}^{+1.800}$ $8.540_{-1.584}^{+2.334}$ $0.96$ $0.360\pm 0.040$ Feng et al. (2022)
HD 44219 HD 44219 b $0.580_{-0.080}^{+0.100}$ $1.260_{-0.070}^{+0.020}$ $1.03$ $0.020\pm 0.020$ Philipot et al. (2023)
HD 45350 HD 45350 b $1.821_{-0.071}^{+0.075}$ $1.958_{-0.030}^{+0.029}$ $1.05$ $0.292\pm 0.060$ Rosenthal et al. (2021)
HD 48265 HD 48265 b $1.525_{-0.050}^{+0.049}$ $1.814_{-0.031}^{+0.030}$ $1.04$ $0.35\pm 0.20$ Barbato et al. (2018)
HD 50499 HD 50499 b $1.346_{-0.087}^{+0.083}$ $3.846\pm 0.039$ $1.25$ $0.305\pm 0.060$ Rosenthal et al. (2021)
HD 50499 c $2.93_{-0.18}^{+0.73}$ $9.02_{-0.33}^{+1.73}$ $1.25$ $0.305\pm 0.060$ Rickman et al. (2019)
HD 50554 HD 50554 b $4.35_{-0.18}^{+0.19}$ $2.265_{-0.036}^{+0.035}$ $1.04$ $-0.021\pm 0.060$ Rosenthal et al. (2021)
HD 564 HD 564 b $0.330\pm 0.030$ $0.0201.200_{\pm}$ 0.020 $1.11$ $-0.230\pm 0.020$ Moutou et al. (2015)
HD 62364 HD 62364 b $17.50_{-0.59}^{+0.60}$ $6.248_{-0.072}^{+0.070}$ $1.19$ $-0.140\pm 0.030$ Xiao et al. (2023)
HD 66428 HD 66428 b $10.900_{-3.845}^{+2.400}$ $3.365_{-0.157}^{+0.141}$ $1.01$ $0.250\pm 0.090$ Feng et al. (2022)
HD 66428 c $1.760_{-0.041}^{+3.400}$ $9.408_{-1.267}^{+1.945}$ $1.01$ $0.250\pm 0.090$ Feng et al. (2022)
HD 6718 HD 6718 b $1.68\pm 0.24$ $3.55\pm 0.18$ $1.02$ $-0.07\pm 0.10$ Stassun et al. (2017)
HD 70642 HD 70642 b $3.90_{-0.27}^{+0.30}$ $3.295\pm 0.021$ $1.01$ $0.170\pm 0.080$ Xiao et al. (2023)
HD 72659 HD 72659 b $2.990_{-0.098}^{+2.600}$ $4.691_{-0.202}^{+0.185}$ $1.09$ $-0.020\pm 0.080$ Feng et al. (2022)
HD 73256 HD 73256 c $16.00\pm 1.00$ $3.800\pm 0.100$ $0.97$ $0.230\pm 0.060$ Philipot et al. (2023)
HD 73267 HD 73267 b $4.60_{-1.10}^{+1.20}$ $2.195_{-0.025}^{+0.024}$ $0.93$ $0.050\pm 0.030$ Xiao et al. (2023)
HD 73526 HD 73526 c $2.25\pm 0.13$ $0.0201.030_{\pm}$ 0.020 $0.98$ $0.25\pm 0.20$ Wittenmyer et al. (2014)
HD 74156 HD 74156 c $9.84_{-0.86}^{+0.90}$ $3.823\pm 0.043$ $1.15$ $0.079\pm 0.060$ Xiao et al. (2023)
HD 7449 HD 7449 b $8.170_{-2.699}^{+3.100}$ $2.438_{-0.063}^{+0.062}$ $1.12$ $-0.140\pm 0.040$ Feng et al. (2022)
HD 74698 HD 74698 c $0.400\pm 0.060$ $4.50\pm 0.20$ $1.05$ $0.060\pm 0.040$ Frensch et al. (2023)
HD 75302 HD 75302 b $5.40_{-0.40}^{+0.50}$ $5.30_{-0.10}^{+0.20}$ $1.01$ $0.08\pm 0.10$ Philipot et al. (2023)
HD 75732 55 Cnc d $3.86_{-0.15}^{+0.60}$ $5.446\pm 0.020$ $0.98$ $0.383\pm 0.060$ Stassun et al. (2017)
HD 82943 HD 82943 b $1.539_{-0.052}^{+0.053}$ $1.189_{-0.018}^{+0.017}$ $1.08$ $0.25\pm 0.20$ Rosenthal et al. (2021)
HD 83443 HD 83443 c $1.50_{-0.20}^{+0.50}$ $8.00\pm 0.80$ $0.95$ $0.34\pm 0.20$ Errico et al. (2022)
HD 8535 HD 8535 b $0.680_{-0.040}^{+0.070}$ $2.450_{-0.060}^{+0.040}$ $1.18$ $0.010\pm 0.020$ Naef et al. (2010)
HD 87883 HD 87883 b $5.410_{-0.635}^{+0.500}$ $4.055_{-0.161}^{+0.147}$ $0.83$ $0.104\pm 0.060$ Feng et al. (2022)
HD 89839 HD 89839 b $5.01_{-0.76}^{+0.80}$ $4.761\pm 0.042$ $1.24$ $0.010\pm 0.040$ Xiao et al. (2023)
HD 92788 HD 92788 c $2.81_{-0.17}^{+0.18}$ $8.26_{-0.28}^{+0.36}$ $1.08$ $0.277\pm 0.060$ Rosenthal et al. (2021)
HD 94771 HD 94771 b $0.530\pm 0.030$ $3.480\pm 0.030$ $1.00$ $0.190\pm 0.020$ Frensch et al. (2023)
HD 95128 47 UMa b $2.438\pm 0.085$ $2.059_{-0.032}^{+0.031}$ $1.01$ $0.026\pm 0.060$ Rosenthal et al. (2021)
47 UMa c $0.497_{-0.030}^{+0.032}$ $3.404_{-0.056}^{+0.054}$ $1.01$ $0.026\pm 0.060$ Rosenthal et al. (2021)
HD 9826 ups And d $4.100\pm 0.100$ $2.517_{-0.027}^{+0.026}$ $1.29$ $0.122\pm 0.060$ Rosenthal et al. (2021)
HD 99109 HD 99109 b $0.474\pm 0.035$ $1.122\pm 0.017$ $0.97$ $0.384\pm 0.060$ Rosenthal et al. (2021)
HIP 2247 BD-17 63 b $2.85\pm 1.16$ $1.340\pm 0.020$ $0.74$ $-0.020\pm 0.020$ Stassun et al. (2017)
HIP 54597 HIP 54597 b $2.010\pm 0.030$ $3.810\pm 0.020$ $0.78$ $-0.250\pm 0.020$ Frensch et al. (2023)
 [^18] From left to right, the columns report the HD or HIP identifier of the host star, the name of the cold Jupiter, its mass and semi-major axis, the stellar mass and metallicity, and the reference for the cold-Jupiter mass determination. Asterisks denote true masses instead of minimum masses thanks to the combination of radial velocities and measurements of Hipparcos-Gaia proper motion anomalies (e.g., Kervella et al. 2022).
Appendix F Occurrence rates of cold Jupiters in the small planet and in the large RV comparison sample as a function of stellar metallicity and/or mass.
Table 7: Occurrence rates of cold Jupiters in the small planet sample and in the large RV comparison one as a function of stellar metallicity.
$M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$
[Fe/H] $N_{\star}\,|\,d$ $C$ $f_{\rm CJ}$ [%] $N_{\star}\,|\,d$ $C$ $f_{\rm CJ}$ [%]
Small planet (transit+radial velocity) sample
$\overline{\rm[Fe/H]}=-0.011\pm 0.005$ 217| 27conditional21727217\,|\,27217 | 27 0.928 $13.4^{+2.8}_{-2.0}$ 217| 23conditional21723217\,|\,23217 | 23 0.958 $11.1^{+2.5}_{-1.8}$
$\rm[Fe/H]<-0.1$ 64| 2conditional64264\,|\,264 | 2 0.943 $3.3^{+4.1}_{-1.1}$ 64| 2conditional64264\,|\,264 | 2 0.969 $3.2^{+4.0}_{-1.0}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 86| 6conditional86686\,|\,686 | 6 0.918 $7.6^{+4.1}_{-2.0}$ 86| 3conditional86386\,|\,386 | 3 0.952 $3.7^{+3.3}_{-1.1}$
$\rm[Fe/H]>+0.1$ 67| 19conditional671967\,|\,1967 | 19 0.923 $30.7^{+6.4}_{-5.2}$ 67| 18conditional671867\,|\,1867 | 18 0.954 $28.2^{+6.2}_{-4.9}$
$\rm[Fe/H]\leq 0$ 102| 3conditional1023102\,|\,3102 | 3 0.932 $3.2^{+2.9}_{-1.0}$ 102| 3conditional1023102\,|\,3102 | 3 0.961 $3.1^{+2.8}_{-0.9}$
$\rm[Fe/H]>0$ 115| 24conditional11524115\,|\,24115 | 24 0.923 $22.6^{+4.5}_{-3.5}$ 115| 20conditional11520115\,|\,20115 | 20 0.954 $18.2^{+4.2}_{-3.1}$
Merged (AAT+CLS+HARPS) radial velocity comparison sample
$\overline{\rm[Fe/H]}=-0.072\pm 0.009$ 1167| 118conditional11671181167\,|\,1181167 | 118 0.974 $10.4^{+1.0}_{-0.8}$ 1167| 113conditional11671131167\,|\,1131167 | 113 0.988 $9.8^{+0.9}_{-0.8}$
$\rm[Fe/H]<-0.1$ 430| 22conditional43022430\,|\,22430 | 22 0.974 $5.2^{+1.3}_{-0.9}$ 430| 21conditional43021430\,|\,21430 | 21 0.988 $4.9^{+1.3}_{-0.8}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 413| 31conditional41331413\,|\,31413 | 31 0.971 $7.7^{+1.5}_{-1.1}$ 413| 31conditional41331413\,|\,31413 | 31 0.986 $7.6^{+1.5}_{-1.1}$
$\rm[Fe/H]>+0.1$ 324| 65conditional32465324\,|\,65324 | 65 0.977 $20.5^{+2.4}_{-2.1}$ 324| 61conditional32461324\,|\,61324 | 61 0.990 $19.0^{+2.4}_{-2.0}$
$\rm[Fe/H]\leq 0$ 624| 33conditional62433624\,|\,33624 | 33 0.973 $5.4^{+1.1}_{-0.8}$ 624| 32conditional62432624\,|\,32624 | 32 0.988 $5.2^{+1.0}_{-0.7}$
$\rm[Fe/H]>0$ 543| 85conditional54385543\,|\,85543 | 85 0.974 $16.1^{+1.7}_{-1.5}$ 543| 81conditional54381543\,|\,81543 | 81 0.988 $15.1^{+1.7}_{-1.4}$
 [^19] From left to right, the columns report the stellar metallicity ([Fe/H]), the total number of stars ($N_{\star}$) and the number of stars with confirmed CJs ($d$), the average completeness ($C$), and the occurrence rate of CJs ($f_{\rm CJ}$), for the small planet (top) and RV comparison sample (bottom) and their sub-samples at sub-solar, solar, and super-solar metallicity, by considering two ranges for the CJ mass: $0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ (columns 2-4) and $0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ (columns 5-7). We note that $f_{\rm CJ}$ for the small planet sample is equivalent to $f_{\rm CJ|SP}$.
Table 8: Occurrence rates of cold Jupiters in the small planet and in the large RV comparison sample as a function of stellar metallicity in three intervals of stellar mass.
$M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$
[Fe/H] $N_{\star}\,|\,d$ $C$ $f_{\rm CJ}$ [%] $N_{\star}\,|\,d$ $C$ $f_{\rm CJ}$ [%]
$0.6\leq M_{\star}<0.8\,\rm M_{\odot}$
Small planet (transit+radial velocity) sample
$\overline{\rm[Fe/H]}=-0.114\pm 0.011$ 58| 4conditional58458\,|\,458 | 4 0.947 $7.3^{+5.1}_{-2.1}$ 58| 3conditional58358\,|\,358 | 3 0.971 $5.3^{+4.7}_{-1.6}$
$\rm[Fe/H]<-0.1$ 26| 1conditional26126\,|\,126 | 1 0.959 $4.0^{+8.1}_{-1.2}$ 26| 1conditional26126\,|\,126 | 1 0.979 $3.9^{+8.0}_{-1.2}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 23| 2conditional23223\,|\,223 | 2 0.924 $9.4^{+10.2}_{-3.1}$ 23| 1conditional23123\,|\,123 | 1 0.955 $4.5^{+9.1}_{-1.4}$
$\rm[Fe/H]>+0.1$ 9| 1conditional919\,|\,19 | 1 0.972 $11.4^{+18.6}_{-4.0}$ 9| 1conditional919\,|\,19 | 1 0.986 $11.3^{+18.4}_{-3.9}$
Merged (AAT+CLS+HARPS) radial velocity comparison sample
$\overline{\rm[Fe/H]}=-0.121\pm 0.014$ 142| 4conditional1424142\,|\,4142 | 4 0.983 $2.9^{+2.2}_{-0.8}$ 142| 4conditional1424142\,|\,4142 | 4 0.993 $2.8^{+2.1}_{-0.9}$
$\rm[Fe/H]<-0.1$ 68| 1conditional68168\,|\,168 | 1 0.980 $1.5^{+3.3}_{-0.4}$ 68| 1conditional68168\,|\,168 | 1 0.991 $1.5^{+3.2}_{-0.4}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 48| 2conditional48248\,|\,248 | 2 0.984 $4.2^{+5.1}_{-1.4}$ 48| 2conditional48248\,|\,248 | 2 0.993 $4.2^{+5.1}_{-1.4}$
$\rm[Fe/H]>+0.1$ 26| 1conditional26126\,|\,126 | 1 0.991 $3.9^{+7.9}_{-1.2}$ 26| 1conditional26126\,|\,126 | 1 0.997 $3.9^{+7.8}_{-1.2}$
$0.8\leq M_{\star}<1.0\,\rm M_{\odot}$
Small planet (transit+radial velocity) sample
$\overline{\rm[Fe/H]}=-0.030\pm 0.007$ 82| 6conditional82682\,|\,682 | 6 0.923 $7.9^{+4.3}_{-2.1}$ 82| 4conditional82482\,|\,482 | 4 0.954 $5.1^{+3.7}_{-1.5}$
$\rm[Fe/H]<-0.1$ 27| 1conditional27127\,|\,127 | 1 0.948 $3.9^{+7.9}_{-1.2}$ 27| 1conditional27127\,|\,127 | 1 0.971 $3.8^{+7.8}_{-1.2}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 37| 0conditional37 037\,|\,037 | 0 0.922 $<3.2$ 37| 0conditional37 037\,|\,037 | 0 0.955 $<3.1$
$\rm[Fe/H]>+0.1$ 18| 5conditional18518\,|\,518 | 5 0.886 $31.3^{+13.1}_{-8.9}$ 18| 3conditional18318\,|\,318 | 3 0.926 $18.0^{+12.7}_{-5.8}$
Merged (AAT+CLS+HARPS) radial velocity comparison sample
$\overline{\rm[Fe/H]}=-0.073\pm 0.016$ 387| 39conditional38739387\,|\,39387 | 39 0.981 $10.3^{+1.8}_{-1.3}$ 387| 36conditional38736387\,|\,36387 | 36 0.992 $9.4^{+1.7}_{-1.3}$
$\rm[Fe/H]<-0.1$ 147| 8conditional1478147\,|\,8147 | 8 0.981 $5.5^{+2.6}_{-1.3}$ 147| 8conditional1478147\,|\,8147 | 8 0.992 $5.5^{+2.5}_{-1.3}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 141| 9conditional1419141\,|\,9141 | 9 0.981 $6.5^{+2.7}_{-1.5}$ 141| 9conditional1419141\,|\,9141 | 9 0.991 $6.4^{+2.7}_{-1.5}$
$\rm[Fe/H]>+0.1$ 99| 22conditional992299\,|\,2299 | 22 0.981 $22.6^{+4.8}_{-3.7}$ 99| 19conditional991999\,|\,1999 | 19 0.992 $19.3^{+4.6}_{-1.3}$
$1.0\leq M_{\star}<1.2\,\rm M_{\odot}$
Small planet (transit+radial velocity) sample
$\overline{\rm[Fe/H]}=0.077\pm 0.008$ 65| 16conditional651665\,|\,1665 | 16 0.915 $26.9^{+6.4}_{-4.9}$ 65| 15conditional651565\,|\,1565 | 15 0.951 $24.3^{+6.2}_{-4.6}$
$\rm[Fe/H]<-0.1$ 10| 0conditional10 010\,|\,010 | 0 0.903 $<10.8$ 10| 0conditional10 010\,|\,010 | 0 0.950 $<10.3$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 24| 4conditional24424\,|\,424 | 4 0.899 $18.5^{+10.9}_{-5.6}$ 24| 2conditional24224\,|\,224 | 2 0.941 $8.9^{+9.7}_{-3.0}$
$\rm[Fe/H]>+0.1$ 31| 12conditional311231\,|\,1231 | 12 0.936 $41.4^{+9.4}_{-8.2}$ 31| 13conditional311331\,|\,1331 | 13 0.959 $43.7^{+9.2}_{-8.4}$
Merged (AAT+CLS+HARPS) radial velocity comparison sample
$\overline{\rm[Fe/H]}=-0.056\pm 0.014$ 549| 63conditional54963549\,|\,63549 | 63 0.972 $11.8^{+1.5}_{-1.3}$ 549| 62conditional54962549\,|\,62549 | 62 0.988 $11.4^{+1.5}_{-1.2}$
$\rm[Fe/H]<-0.1$ 181| 11conditional18111181\,|\,11181 | 11 0.973 $6.2^{+2.3}_{-1.4}$ 181| 10conditional18110181\,|\,10181 | 10 0.988 $5.6^{+2.2}_{-1.3}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 200| 17conditional20017200\,|\,17200 | 17 0.966 $8.8^{+2.5}_{-1.6}$ 200| 17conditional20017200\,|\,17200 | 17 0.984 $8.6^{+2.4}_{-1.6}$
$\rm[Fe/H]>+0.1$ 168| 35conditional16835168\,|\,35168 | 35 0.978 $21.3^{+3.5}_{-2.8}$ 168| 35conditional16835168\,|\,35168 | 35 0.991 $21.0^{+3.5}_{-2.8}$
 [^20] Same columns as Table 19.
Appendix G Occurrence rates of cold Jupiters in the small planet transit, RV, and transit+RV samples, without and with mixed systems, as a function of stellar metallicity.
Table 9: Occurrence rates of cold Jupiters in the small planet transit, RV, and transit+RV samples, without and with the inclusion of mixed systems, as a function of stellar metallicity.
$M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$
[Fe/H] $N_{\star}\,|\,d$ $C$ $f_{\rm CJ}$ [%] $N_{\star}\,|\,d$ $C$ $f_{\rm CJ}$ [%]
Small planet transit sample
$\overline{\rm[Fe/H]}=+0.002\pm 0.006$ 134| 12conditional13412134\,|\,12134 | 12 0.894 $10.0^{+3.4}_{-2.1}$ 134| 12conditional13412134\,|\,12134 | 12 0.938 $9.5^{+3.3}_{-2.0}$
$\rm[Fe/H]<-0.1$ 36| 1conditional36136\,|\,136 | 1 0.910 $3.0^{+6.4}_{-0.9}$ 36| 1conditional36136\,|\,136 | 1 0.949 $2.9^{+6.1}_{-0.9}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 56| 1conditional56156\,|\,156 | 1 0.884 $2.0^{+4.3}_{-0.6}$ 56| 1conditional56156\,|\,156 | 1 0.931 $1.9^{+4.1}_{-0.6}$
$\rm[Fe/H]>+0.1$ 42| 10conditional421042\,|\,1042 | 10 0.889 $26.8^{+8.3}_{-5.9}$ 42| 10conditional421042\,|\,1042 | 10 0.932 $25.5^{+8.0}_{-5.6}$
$\rm[Fe/H]\leq 0$ 62| 2conditional62262\,|\,262 | 2 0.900 $3.6^{+4.4}_{-1.1}$ 62| 2conditional62262\,|\,262 | 2 0.942 $3.4^{+4.2}_{-1.1}$
$\rm[Fe/H]>0$ 72| 10conditional721072\,|\,1072 | 10 0.886 $15.7^{+5.6}_{-3.5}$ 72| 10conditional721072\,|\,1072 | 10 0.931 $14.9^{+5.4}_{-3.3}$
Small planet radial velocity sample
$\overline{\rm[Fe/H]}=-0.033\pm 0.008$ 83| 15conditional831583\,|\,1583 | 15 0.982 $18.4^{+5.0}_{-3.5}$ 83| 11conditional831183\,|\,1183 | 11 0.992 $13.4^{+4.6}_{-2.9}$
$\rm[Fe/H]<-0.1$ 28| 1conditional28128\,|\,128 | 1 0.985 $3.6^{+7.4}_{-1.1}$ 28| 1conditional28128\,|\,128 | 1 0.994 $3.6^{+7.4}_{-1.1}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 30| 5conditional30530\,|\,530 | 5 0.979 $17.0^{+9.0}_{-4.8}$ 30| 2conditional30230\,|\,230 | 2 0.990 $6.7^{+7.7}_{-2.2}$
$\rm[Fe/H]>+0.1$ 25| 9conditional25925\,|\,925 | 9 0.982 $36.7^{+10.3}_{-8.3}$ 25| 8conditional25825\,|\,825 | 8 0.992 $32.3^{+10.4}_{-7.7}$
$\rm[Fe/H]\leq 0$ 40| 1conditional40140\,|\,140 | 1 0.981 $2.5^{+5.4}_{-0.8}$ 40| 1conditional40140\,|\,140 | 1 0.991 $2.5^{+5.3}_{-0.8}$
$\rm[Fe/H]>0$ 43| 14conditional431443\,|\,1443 | 14 0.983 $33.1^{+7.8}_{-6.3}$ 43| 10conditional431043\,|\,1043 | 10 0.992 $23.4^{+7.6}_{-5.2}$
Small planet (transit+radial velocity) sample
$\overline{\rm[Fe/H]}=-0.011\pm 0.005$ 217| 27conditional21727217\,|\,27217 | 27 0.928 $13.4^{+2.8}_{-2.0}$ 217| 23conditional21723217\,|\,23217 | 23 0.958 $11.1^{+2.5}_{-1.8}$
$\rm[Fe/H]<-0.1$ 64| 2conditional64264\,|\,264 | 2 0.943 $3.3^{+4.1}_{-1.1}$ 64| 2conditional64264\,|\,264 | 2 0.969 $3.2^{+4.0}_{-1.0}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 86| 6conditional86686\,|\,686 | 6 0.918 $7.6^{+4.1}_{-2.0}$ 86| 3conditional86386\,|\,386 | 3 0.952 $3.7^{+3.3}_{-1.1}$
$\rm[Fe/H]>+0.1$ 67| 19conditional671967\,|\,1967 | 19 0.923 $30.7^{+6.4}_{-5.2}$ 67| 18conditional671867\,|\,1867 | 18 0.954 $28.2^{+6.2}_{-4.9}$
$\rm[Fe/H]\leq 0$ 102| 3conditional1023102\,|\,3102 | 3 0.932 $3.2^{+2.9}_{-1.0}$ 102| 3conditional1023102\,|\,3102 | 3 0.961 $3.1^{+2.8}_{-0.9}$
$\rm[Fe/H]>0$ 115| 24conditional11524115\,|\,24115 | 24 0.923 $22.6^{+4.5}_{-3.5}$ 115| 20conditional11520115\,|\,20115 | 20 0.954 $18.2^{+4.2}_{-3.1}$
Small planet and mixed transit sample
$\overline{\rm[Fe/H]}=+0.022\pm 0.006$ 151| 16conditional15116151\,|\,16151 | 16 0.891 $11.9^{+3.4}_{-2.2}$ 151| 16conditional15116151\,|\,16151 | 16 0.935 $11.3^{+3.2}_{-2.1}$
$\rm[Fe/H]<-0.1$ 36| 1conditional36136\,|\,136 | 1 0.910 $3.0^{+6.4}_{-0.9}$ 36| 1conditional36136\,|\,136 | 1 0.949 $2.9^{+6.1}_{-0.9}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 63| 1conditional63163\,|\,163 | 1 0.881 $1.8^{+3.9}_{-0.5}$ 63| 1conditional63163\,|\,163 | 1 0.929 $1.7^{+3.7}_{-0.5}$
$\rm[Fe/H]>+0.1$ 52| 14conditional521452\,|\,1452 | 14 0.883 $30.5^{+7.5}_{-5.8}$ 52| 14conditional521452\,|\,1452 | 14 0.928 $29.0^{+7.2}_{-5.6}$
$\rm[Fe/H]\leq 0$ 64| 2conditional64264\,|\,264 | 2 0.899 $3.5^{+4.2}_{-1.1}$ 64| 2conditional64264\,|\,264 | 2 0.942 $3.3^{+4.1}_{-1.1}$
$\rm[Fe/H]>0$ 87| 14conditional871487\,|\,1487 | 14 0.881 $18.3^{+5.2}_{-3.6}$ 87| 14conditional871487\,|\,1487 | 14 0.927 $17.4^{+5.0}_{-3.4}$
Small planet and mixed radial velocity sample
$\overline{\rm[Fe/H]}=-0.018\pm 0.007$ 91| 17conditional911791\,|\,1791 | 17 0.983 $19.0^{+4.8}_{-3.4}$ 91| 13conditional911391\,|\,1391 | 13 0.992 $14.4^{+4.5}_{-2.9}$
$\rm[Fe/H]<-0.1$ 29| 1conditional29129\,|\,129 | 1 0.985 $3.5^{+7.2}_{-1.1}$ 29| 1conditional29129\,|\,129 | 1 0.993 $3.5^{+7.2}_{-1.1}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 34| 6conditional34634\,|\,634 | 6 0.979 $18.0^{+8.4}_{-4.8}$ 34| 3conditional34334\,|\,334 | 3 0.990 $8.9^{+7.4}_{-2.8}$
$\rm[Fe/H]>+0.1$ 28| 10conditional281028\,|\,1028 | 10 0.983 $36.3^{+9.8}_{-7.9}$ 28| 9conditional28928\,|\,928 | 9 0.993 $32.4^{+9.8}_{-7.4}$
$\rm[Fe/H]\leq 0$ 41| 1conditional41141\,|\,141 | 1 0.981 $2.5^{+5.3}_{-0.8}$ 41| 1conditional41141\,|\,141 | 1 0.991 $2.5^{+5.2}_{-0.7}$
$\rm[Fe/H]>0$ 50| 16conditional501650\,|\,1650 | 16 0.983 $32.5^{+7.2}_{-5.9}$ 50| 12conditional501250\,|\,1250 | 12 0.993 $24.2^{+7.0}_{-5.0}$
Small planet and mixed (transit+radial velocity) sample
$\overline{\rm[Fe/H]}=0.007\pm 0.004$ 242| 33conditional24233242\,|\,33242 | 33 0.926 $14.7^{+2.7}_{-2.1}$ 242| 29conditional24229242\,|\,29242 | 29 0.957 $12.5^{+2.5}_{-1.8}$
$\rm[Fe/H]<-0.1$ 65| 2conditional65265\,|\,265 | 2 0.944 $3.3^{+4.0}_{-1.0}$ 65| 2conditional65265\,|\,265 | 2 0.969 $3.2^{+3.9}_{-1.0}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 97| 7conditional97797\,|\,797 | 7 0.915 $7.9^{+3.8}_{-2.0}$ 97| 4conditional97497\,|\,497 | 4 0.950 $4.3^{+3.2}_{-1.3}$
$\rm[Fe/H]>+0.1$ 80| 24conditional802480\,|\,2480 | 24 0.918 $32.7^{+5.8}_{-4.9}$ 80| 23conditional802380\,|\,2380 | 23 0.951 $30.2^{+5.7}_{-4.7}$
$\rm[Fe/H]\leq 0$ 105| 3conditional1053105\,|\,3105 | 3 0.931 $3.1^{+2.8}_{-0.9}$ 105| 3conditional1053105\,|\,3105 | 3 0.961 $3.0^{+2.7}_{-0.9}$
$\rm[Fe/H]>0$ 137| 30conditional13730137\,|\,30137 | 30 0.918 $23.8^{+4.2}_{-3.4}$ 137| 26conditional13726137\,|\,26137 | 26 0.951 $20.0^{+3.9}_{-3.0}$
 [^21] Same columns as Table 19.
Appendix H Occurrence rates of cold Jupiters in the AAT, CLS, and HARPS RV surveys
Table 10: Occurrence rates of cold Jupiters in the CLS survey by using the selection criteria in BL24 (top) and this work (bottom).
$M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$
[Fe/H] $N_{\star}\,|\,d$ $C$ $f_{\rm CJ}$ [%] $N_{\star}\,|\,d$ $C$ $f_{\rm CJ}$ [%]
CLS radial velocity sample (BL24)
$\overline{\rm[Fe/H]}=+0.008\pm 0.004$ 562| 51conditional56251562\,|\,51562 | 51 0.981 $9.2^{+1.9}_{-1.1}$ 562| 46conditional56246562\,|\,46562 | 46 0.993 $8.2^{+1.3}_{-1.0}$
$\rm[Fe/H]<-0.1$ 135| 4conditional1354135\,|\,4135 | 4 0.980 $3.0^{+2.3}_{-0.9}$ 135| 4conditional1354135\,|\,4135 | 4 0.993 $3.0^{+2.3}_{-0.9}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 201| 12conditional20112201\,|\,12201 | 12 0.985 $6.1^{+2.2}_{-1.3}$ 201| 11conditional20111201\,|\,11201 | 11 0.994 $5.5^{+2.1}_{-1.2}$
$\rm[Fe/H]>+0.1$ 226| 35conditional22635226\,|\,35226 | 35 0.979 $15.8^{+2.8}_{-2.1}$ 226| 31conditional22631226\,|\,31226 | 31 0.992 $13.8^{+2.6}_{-2.0}$
CLS radial velocity sample (this work)
$\overline{\rm[Fe/H]}=+0.025\pm 0.004$ 402| 50conditional40250402\,|\,50402 | 50 0.987 $12.6^{+1.8}_{-1.5}$ 402| 45conditional40245402\,|\,45402 | 45 0.995 $11.2^{+1.8}_{-1.4}$
$\rm[Fe/H]<-0.1$ 95| 4conditional95495\,|\,495 | 4 0.990 $4.2^{+3.1}_{-1.2}$ 95| 4conditional95495\,|\,495 | 4 0.996 $4.2^{+3.1}_{-1.2}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 149| 12conditional14912149\,|\,12149 | 12 0.989 $8.1^{+2.8}_{-1.7}$ 149| 11conditional14911149\,|\,11149 | 11 0.996 $7.4^{+2.7}_{-1.6}$
$\rm[Fe/H]>+0.1$ 158| 34conditional15834158\,|\,34158 | 34 0.985 $21.8^{+3.6}_{-2.9}$ 158| 30conditional15830158\,|\,30158 | 30 0.994 $19.1^{+3.5}_{-2.7}$
 [^22] Same columns as Table 19.
Table 11: Occurrence rates of cold Jupiters in the AAT (top), CLS (middle), and HARPS (bottom) surveys.
$M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$
[Fe/H] $N_{\star}\,|\,d$ $C$ $f_{\rm CJ}$ [%] $N_{\star}\,|\,d$ $C$ $f_{\rm CJ}$ [%]
AAT radial velocity survey
$\overline{\rm[Fe/H]}=-0.045\pm 0.018$ 196| 20conditional19620196\,|\,20196 | 20 0.971 $10.5^{+2.6}_{-1.8}$ 196| 20conditional19620196\,|\,20196 | 20 0.989 $10.3^{+2.6}_{-1.8}$
$\rm[Fe/H]<-0.1$ 69| 3conditional69369\,|\,369 | 3 0.977 $4.4^{+4.0}_{-1.4}$ 69| 3conditional69369\,|\,369 | 3 0.992 $4.4^{+3.9}_{-1.3}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 66| 5conditional66566\,|\,566 | 5 0.969 $7.8^{+4.7}_{-2.2}$ 66| 5conditional66566\,|\,566 | 5 0.989 $7.7^{+4.6}_{-2.1}$
$\rm[Fe/H]>+0.1$ 61| 12conditional611261\,|\,1261 | 12 0.966 $20.4^{+6.2}_{-4.2}$ 61| 12conditional611261\,|\,1261 | 12 0.988 $19.9^{+6.1}_{-4.1}$
CLS radial velocity survey
$\overline{\rm[Fe/H]}=+0.025\pm 0.004$ 402| 50conditional40250402\,|\,50402 | 50 0.987 $12.6^{+1.8}_{-1.5}$ 402| 45conditional40245402\,|\,45402 | 45 0.995 $11.2^{+1.8}_{-1.4}$
$\rm[Fe/H]<-0.1$ 95| 4conditional95495\,|\,495 | 4 0.990 $4.2^{+3.1}_{-1.2}$ 95| 4conditional95495\,|\,495 | 4 0.996 $4.2^{+3.1}_{-1.2}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 149| 12conditional14912149\,|\,12149 | 12 0.989 $8.1^{+2.8}_{-1.7}$ 149| 11conditional14911149\,|\,11149 | 11 0.996 $7.4^{+2.7}_{-1.6}$
$\rm[Fe/H]>+0.1$ 158| 34conditional15834158\,|\,34158 | 34 0.985 $21.8^{+3.6}_{-2.9}$ 158| 30conditional15830158\,|\,30158 | 30 0.994 $19.1^{+3.5}_{-2.7}$
HARPS radial velocity survey
$\overline{\rm[Fe/H]}=-0.122\pm 0.011$ 782| 75conditional78275782\,|\,75782 | 75 0.968 $9.9^{+1.2}_{-1.0}$ 782| 74conditional78274782\,|\,74782 | 74 0.984 $9.6^{+1.2}_{-1.0}$
$\rm[Fe/H]<-0.1$ 344| 19conditional34419344\,|\,19344 | 19 0.970 $5.7^{+1.5}_{-1.0}$ 344| 18conditional34418344\,|\,18344 | 18 0.986 $5.3^{+1.5}_{-1.0}$
$-0.1\leq\rm[Fe/H]\leq+0.1$ 272| 24conditional27224272\,|\,24272 | 24 0.963 $9.2^{+2.1}_{-1.5}$ 272| 24conditional27224272\,|\,24272 | 24 0.981 $9.0^{+2.1}_{-1.5}$
$\rm[Fe/H]>+0.1$ 166| 32conditional16632166\,|\,32166 | 32 0.972 $19.8^{+3.5}_{-2.8}$ 166| 32conditional16632166\,|\,32166 | 32 0.986 $19.5^{+3.4}_{-2.7}$
 [^23] Same columns as Table 19.